### v2

# KeyDiff: Key Similarity-Based KV Cache Eviction for Long-Context LLM Inference in Resource-Constrained Environments

Abstract
In this work, we demonstrate that distinctive keys during LLM inference tend to have high attention scores. We explore this phenomenon and propose KeyDiff, a training-free KV cache eviction method based on key similarity. This method facilitates the deployment of LLM-based application requiring long input prompts in resource-constrained environments with limited memory and compute budgets. Unlike other KV cache eviction methods, KeyDiff can process arbitrarily long prompts within strict resource constraints and efficiently generate responses. We demonstrate that KeyDiff computes the optimal solution to a KV cache selection problem that maximizes key diversity, providing a theoretical understanding of KeyDiff. Notably, KeyDiff does not rely on attention scores, allowing the use of optimized attention mechanisms like FlashAttention. We demonstrate the effectiveness of KeyDiff across diverse tasks and models, illustrating a performance gap of less than 0.04% with 8K cache budget ($\sim$ 23% KV cache reduction) from the non-evicting baseline on the LongBench benchmark for Llama 3.1-8B and Llama 3.2-3B.
Machine Learning, ICML
1 Introduction
Large language models (LLMs) have revolutionized natural language processing, enabling tasks like document summarization, code generation, question answering (Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023; Dubey et al., 2024). Many exciting applications of LLMs, such as retrieval augmented generation (Lewis et al., 2020) and reasoning tasks (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2024), require very long input prompts, or long context support, to provide a valuable end-user experience. However, deploying these models in production environments has grown in importance and managing long input prompts remains a challenge due to large compute and memory requirements.
Due to cost and privacy concerns, there has been a recent focus on deploying LLMs on the edge, i.e. directly on devices such as mobile phones or personal computers (Alizadeh et al., 2023; Liu et al., 2024c; van Baalen et al., 2024; Xu et al., 2024a). This imposes tight resource constraints, such as limited or shared compute resources, limited available DRAM, and even power limits in the mobile setting. Many techniques, such as system optimization (Alizadeh et al., 2023; Liu et al., 2024c; Xu et al., 2024a), model quantization (van Baalen et al., 2024), and efficient text generation (Goel et al., 2024) help to enable on-device LLM inference, but the main challenge for long context application on-device deployment is effectively managing key and value states.
In a naïve implementation of autoregressive text generation, key and value states (KVs) must be recomputed for each new generated token. In a decoder-only transformer model, one can instead store the previously computed KVs in a KV cache for later use, allowing for a significant decrease in LLM inference latency. This comes at the cost of maintaining the KV cache, whose memory footprint grows linearly with the input sequence length (Yang et al., 2024a) and will violate any imposed memory limit with a sufficiently large input.
One solution to bound this memory overhead is to fix the number of KVs in the cache, or cache budget, and use a cache eviction policy to remove KVs from the cache when the budget has been exceeded, bounding the KV cache memory usage. However, state-of-the-art cache eviction policies retain KVs with high attention weights (Xiao et al., 2024; Oren et al., 2024; Zhang et al., 2024b), which requires explicitly materializing the attention matrix of the entire input prompt. This results in large intermediate memory overhead during cache eviction, often violating memory limits.
This intermediate cache memory overhead can be mitigated by segmenting the input prompt into blocks, sequentially processing each block by the LLM and caching the intermediate KVs (Kwon et al., 2023; Agrawal et al., 2023; Xu et al., 2024b). The cache eviction policy can be applied block-wise to satisfy imposed memory constraints (Figure 1). However, our experiments show that applying existing attention-based eviction methods with block prompt processing results in performance degradation of downstream tasks (Table 1). This is due to the limited scope of the attention computation during block prompt processing: attention scores are only computed for a subsequence of the original prompt and only contain local information. These partial scores may not be representative of a token’s global importance across the entire prompt and can result in incorrect cache eviction. This means that we need an eviction method that is independent from attention and cheap to compute during inference while maintaining model capability.
The “attention sink” phenomenon (Xiao et al., 2024; Sun et al., 2024) describes the observation that, in decoder-only transformer-based LLMs, high attention scores are assigned to the first few tokens in the KV cache regardless of the input. This implies that eviction can be determined by only inspecting KVs and that an attention-free cache eviction criteria exists. In fact, we observe that keys with lower average pairwise cosine similarity also have higher attention outputs, as shown in Figure 2.
Motivated by these observations, we propose an eviction method called KeyDiff, that evicts keys from the cache with the highest cosine similarity to a cache representative called an anchor. We also show that KeyDiff computes the optimal solution of the KV cache selection problem that minimizes pairwise key cosine similarity. Our experiments demonstrate that KeyDiff can achieve an accuracy drop of $\leq$ .04% on LongBench (Bai et al., 2024) with 8K token budget ($\sim$ 23% KV cache reduction), and can match or outperform state-of-the-art eviction methods, especially when using block prompt processing. Our contributions are as follows:
  • We present the empirical observation that low pairwise cosine similarity among keys is correlated with higher attention output.
  • We design an eviction policy based on this observation, called KeyDiff, that enables long context LLM inference with limited memory overhead.
  • We show that KeyDiff computes the optimal solution to a KV cache selection problem that maximizes key diversity measured by pairwise cosine similarity.
  • We demonstrate the effectiveness of KeyDiff on the LongBench task suite, achieving only 1.5% and 0.04% performance gap from the eviction-free baseline, using 6K and 8K cache budgets, respectively.
2 Background
Figure 1: An example of block prompt processing with KV cache eviction. The input prompt having length of 7 is segmented by three blocks, and a transformer layer in LLM processes each block by (1) Computing Key-Value states from inputs, (2) computing attention, (3) computing the eviction score, and (4) performing eviction based on the eviction score to satisfy the memory constraints (e.g., at most 4 tokens can reside in the cache). After each block processing, the KV cache is updated and passed to the next round of block processing, satisfying imposed memory constraints on the KV cache.
2.1 Transformers
The Transformer architecture (Vaswani, 2017) processes input data using a sequence of transformer blocks. A transformer block $f$ takes a sequence $X=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T\times d}$ as input and applies the causal self-attention operator $\operatorname{Attention}$ followed by a feed-forward network $\operatorname{FF}$ with optional gating (Shazeer, 2020) to produce the output $X^{\prime}=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{T})\in\mathbb{R}^{T\times d}$:
$X^{\prime}=f(X)=\operatorname{FF}(\operatorname{Attention}(X)),$ (1)
The causal $\operatorname{Attention}$ operator projects each input token $x_{t}$ with matrices $W_{q},W_{k},W_{v}\in\mathbb{R}^{d\times d}$ into key, query, and value matrices ($K=XW_{k}$, $Q=XW_{Q}$, $V=XW_{V}$, respectively) then applies the following relation to produce the attention output  [^1] The multi-head extension and output projections are omitted for brevity.:
$O^{\operatorname{attn}}=\operatorname{Softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+M\right)V=AV$ (2)
where $O^{\operatorname{attn}}\in\mathbb{R}^{T\times d}$, and the causal attention mask $M$ is an upper triangular matrix with nonzero values of $\infty$- ∞.
2.2 KV Caching
When the $\operatorname{Attention}$ operator processes a new token $x_{T+1}$, it must also recompute the prior KV states for tokens $x_{0},\ldots,x_{T}$. This can be avoided by storing previously computed KVs in a KV cache $\mathcal{C}=(K,V)$ for later reuse and append the new KV corresponding $x_{T+1}$ to the cache. We can apply Equation 2 to an existing KV cache $\mathcal{C}$ as follows:
$o^{\operatorname{attn}}_{T+1}=\operatorname{Softmax}\left(\frac{q_{T+1}[K\|k_{T+1}]^{\top}}{\sqrt{d}}+M\right)[V\|v_{T+1}],$ (3)
where $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ are the key, query, and value states of $x_{T+1}$, and $[X\|x_{T+1}]$ represents the concatenation of $x_{T+1}$ to an existing tensor $X$ along the time dimension, $M$ is the causal attention mask accounting for both the KV cache and $x_{T+1}$.
KV caching dramatically reduces the latency of $\operatorname{Attention}$ by only computing $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ for each token $x_{T+1}$ and reusing the KVs in $\mathcal{C}$. However, the size of the KV cache increases linearly with the number of processed tokens and dominates the memory footprint in long-context applications (Yang et al., 2024a).
2.3 KV Cache Eviction Methods
To limit the memory footprint of the KV cache, we fix a cache budget $N$, which is the maximum number of tokens to be stored in the cache. If a new KV is added to the cache and the updated cache size is greater than $N$, we must evict KVs from the cache until the cache budget is met. The eviction policy $\pi_{N}(\mathcal{C})$ evicts a subset of KVs from $\mathcal{C}$ and returns a new cache $\mathcal{C}^{\prime}$ containing at most $N$ KVs:
$\mathcal{C}$ $\leftarrow([K\|k_{t+1}],[V\|v_{t+1}])$ (4)
$\mathcal{C}^{\prime}$ $\leftarrow\pi_{N}(\mathcal{C})$
Attention-Based Eviction Policies
Attention-based eviction policies $\pi_{N}^{\text{attn}}$ use aggregated attention values to rank each KVs’ relative importance and keep the $N$ highest scoring KVs. For a given attention weight aggregation function $\phi$, the eviction policy $\pi_{N}^{\text{attn}}$ performs the following steps:
$S$ $=\texttt{topk}(\phi(A),N)$ (5)
$K^{\prime}$ $=\texttt{gather}(K,S)$
$V^{\prime}$ $=\texttt{gather}(V,S)$
where $\texttt{topk}(x,N)$ returns the indices of $N$-largest values of $x$ and $\texttt{gather}(X,S)$ gathers columns indexed by $S$.
Attention-based eviction methods prioritize KV pairs with higher attention scores to past tokens. This is problematic when applying block prompt processing: all input tokens are not simultaneously accessible within $\operatorname{Attention}$, only those in the current block and cache. This can result in an incorrect eviction decision. Additionally, attention-based eviction often requires explicitly materializing $A$, which can be resource intensive. We discuss the attention-based eviction policies further in Section 5.
2.4 KV Caching in Resource-Constrained Environments
Existing eviction policies like Zhang et al. (2024b); Oren et al. (2024) focus on processing the entire input prompt at once: KVs are computed for each token in the prompt and stored in a cache $\mathcal{C}$, then the eviction policy $\pi_{N}$ is applied to reduce the number of tokens in $\mathcal{C}^{\prime}$ to $N$, before token generation. However, the intermediate cache $\mathcal{C}$ before eviction will grow to the size of the input prompt. This can often exceed model’s allocated memory limit when deploying long context applications in resource constrained environments.
As demonstrated in efficient LLM inference frameworks (Agrawal et al., 2023; Holmes et al., 2024; Kwon et al., 2023; Xu et al., 2024a), one solution is to apply $\pi_{N}$ more frequently by segmenting $X$ into non-overlapping blocks $X=[X_{0},X_{1},\ldots,X_{m-1}]$, where $X_{i}=[x_{Bi},\ldots,x_{B(i+1)-1}]$, $B$ is the block size, and $m=\lceil T/B\rceil$, and iteratively updating the cache by exploiting causality, applying Equation 4 in a block-wise fashion:
$\mathcal{C}_{i}$ $\leftarrow([K_{i-1}\|k_{Bi:B(i+1)-1}],[V_{i-1}\|v_{Bi:B(i+1)-1}])$ (6)
$\mathcal{C}_{i}^{\prime}$ $\leftarrow\pi_{N}(\mathcal{C}_{i}),\quad\mathcal{C}_{i}\leftarrow\mathcal{C}_{i}^{\prime},\quad\mathcal{C}_{0}=\emptyset,$
where $k_{Bi:B(i+1)-1}$ and $v_{Bi:B(i+1)-1}$ are the keys and values selected from $X_{i}$ respectively, and $\mathcal{C}_{i}=(K_{i},V_{i})$ is the KV cache after processing the first $i$ prompt blocks. As in Equation 4, we concatenate the $B$ new KVs to the current cache, apply the eviction policy $\pi_{N}$ and update the cache in Equation 6.
We refer to this as block prompt processing. Its main advantage is the ability to control of the compute and memory overhead of KV cache management by adjusting the block size $B$ and cache budget $N$. Note that, in a decoder-based architecture, applying block prompt processing to $X$ with $B=T$ yields the same result as processing all of $X$ at once and choosing $B=1$ corresponds to the token generation phase of LLM evaluation.
Attention-Based Token Eviction Challenges
Despite its advantages, block prompt processing introduces a challenge for KV cache eviction: eviction decisions in block $X_{i}$ impact the cache used by $X_{i+1}$, causing eviction errors to compound over time. When the model processes $X_{i}$, attention-based eviction methods retain KVs with high attention weights derived from $X_{0},\ldots,X_{i}$ rather than all of $X$, which may prematurely evict KVs with high weights in upcoming blocks.
(a) Layer 2 and Head 1
(b) Layer 11 and Head 1
Figure 2: Cosine similarity of the keys and attention weights. Measured from Llama 3.2-3B-Instruct and the first sample from the NarrativeQA dataset in LongBench. Truncated to the first 64 tokens for visualization.
3 Method
We demonstrate a negative correlation between attention scores and the similarity among keys (Section 3.1) and leverage this observation to develop KeyDiff (Section 3.2), followed by a theoretical justification of the method and preliminary evidence of its efficacy (Section 3.3).
3.1 Correlation of Attention Scores and Key Dissimilarity
To address the shortcomings of attention-based KV cache eviction in Section 2.4, we develop an alternative attention-free scoring metric that retains significant KVs across blocks while being resource efficient. We recall the “attention sink” phenomenon: LLMs often assign high attention weight to the first few tokens, regardless of the input (Xiao et al., 2024; Sun et al., 2024); these highly weighted tokens are called sink tokens. However, the index of the sink tokens can vary across heads and layers and reside deeper in the sequence than the first few tokens. This observation motivates the following hypothesis: high attention scores can be determined by the intrinsic properties of the keys rather than by any particular combination of keys and queries.
Correlation of Key Similarity and Attention Scores
We evaluate our hypothesis by inspecting the cosine similarities between keys computed inside an attention block. We visualize the pairwise cosine similarities between keys along with the attention weights in two particular heads and layers in Figure 2. We observe that keys with lower cosine similarity with other keys exhibit higher relative attention scores regardless of the choice of query, such as the 4th and 15th keys in Figure 2(a), or the 1st key in Figure 2(b). Pairwise cosine similarity of keys is solely a function of the keys in the cache, which are independent of input queries; the surprising aspect of Figure 2 is the negative correlation with attention weights. These distinctive keys essentially recover the attention sink phenomenon (Xiao et al., 2024).
3.2 KeyDiff
Based on the observation in Section 3.1, we propose KeyDiff, which evicts tokens from the KV cache based on key similarity. If the cache $\mathcal{C}$ has intermediate size $n$ and budget $N$ where $n>N$, $\pi^{\textsc{KeyDiff}{}}_{N}$ is defined as:
$S$ $=\texttt{topk}(-\operatorname{CosSim}(K)\mathbf{1},N)$ (7)
$K^{\prime}$ $=\texttt{gather}(K,S)$
$V^{\prime}$ $=\texttt{gather}(V,S)$
where $K\in\mathbb{R}^{n\times d}$ and $V\in\mathbb{R}^{n\times d}$ are the cached keys and values, $\operatorname{CosSim}(K)\in\mathbb{R}^{n\times n}$ is the pairwise cosine similarity matrix of keys in $K$ with $\operatorname{CosSim}(K)_{ij}=\frac{\langle k_{i},k_{j}\rangle}{\|k_{i}\|\|k_{j}\|}$, and $\mathbf{1}\in\mathbb{R}^{n}$ is a vector of ones.
Figure 3: An overview of KeyDiff. (1) KeyDiff first computes the anchor vector by taking the average of the keys in the KV cache, (2) computes the cosine similarity between the keys and the anchor resulting in eviction scores whose color intensities indicate the score values, and (3) retains the KV pairs with the lowest similarities.
Efficient Variant of KeyDiff
Unlike attention-based eviction policies, KeyDiff does not require access to the attention weights $A$, facilitating optimized attention kernels that do not materialize $A$, such as FlashAttention (Dao et al., 2022). However, computing the pairwise cosine similarities runs in $\mathcal{O}(n^{2})$ time. Fortunately, the operation
$\texttt{topk}(-\operatorname{CosSim}(K)\mathbf{1},N)$
can be computed in $\mathcal{O}(n)$ operations on average by computing the score of each token as
$s_{i}=-\operatorname{CosSim}(\mu(\hat{K}),\hat{k}_{i})$
where $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\hat{k}_{i}$ and $\hat{k}_{i}=\frac{k_{i}}{||k_{i}||}$, and retaining the tokens with the $N$ largest scores with a variation of the quick select algorithm. This simplification is derived in Section 3.3.
The eviction score of the efficient expression of KeyDiff is defined as follows:
$s$ $=-\operatorname{CosSim}(\mu(\hat{K}),K)$ (8)
$S$ $=\texttt{topk}(s,N)$
We refer to $\mu(\hat{K})$ as the anchor vector. Experimentation has shown that the anchor vector $\mu(\hat{K})$ can be replaced with $\mu(K)$ without losing accuracy (see Table 2). We evaluate this efficient version of KeyDiff described in Figure 3 in the following sections.
3.3 Derivation of KeyDiff from an Optimization Perspective
[…]
This is a combinatorial optimization problem, which is difficult to solve efficiently, particularly during inference. However, we can relax Section 3.3 to produce a more tractable approximation to the original problem. First, we rewrite Section 3.3 by normalizing keys $k_{i}$ such that $\hat{k}_{i}=\frac{k_{i}}{\|k_{i}\|}$ resulting in:
$\sum\limits_{i\in S}\sum\limits_{j\in S}\hat{k}_{i}\cdot\hat{k}_{j}$ $=\sum\limits_{i\in S}\hat{k}_{i}\cdot\left(\sum\limits_{j\in S}\hat{k}_{j}\right)=\sum\limits_{i\in S}\langle\hat{k}_{i},N\mu(\hat{K}_{S})\rangle$
[…]
Key Cache Diversity and KeyDiff
To empirically verify KeyDiff’s ability to retain diverse keys, we apply PCA to the keys computed in an attention block of Llama 3.2-3B-Instruct after evaluating a long context prompt. We visualize the distribution of retained and evicted keys by applying sink attention (Xiao et al., 2024), TOVA (Oren et al., 2024), and KeyDiff as eviction policies in Figure 4. Visual observation reveals the tokens retained by sink attention and TOVA tend to tightly cluster together while KeyDiff’s retained tokens are more evenly distributed. This observation generalizes across heads and layers, as shown in Section A.3.
We also visualize the log determinant of the Gram matrix of the key cache $\log(\det(KK^{T}))$ generated using different eviction policies in Figure 5. This quantity corresponds to the volume of space spanned by the keys in $\mathcal{C}$. The distribution of volumes for KeyDiff attains higher values, indicating that the retained keys are more distinctive than TOVA and sink, corroborating the results of Figure 4. Details of how these plots were generated are discussed in Section B.1.
(a) Sink Attention
(b) TOVA
(c) KeyDiff
(d) Retained keys only
Figure 4: (a, b, and c) PCA Visualizations in two dimensions of a key cache managed with Sink, TOVA, and KeyDiff. Retained tokens are blue, while evicted tokens are orange. The retained tokens using Sink and TOVA cluster together, while KeyDiff retains more diverse keys. Keys are taken from layer $5$ and head $3$ of Llama3.2-3B-Instruct, and generated using the NarrativeQA dataset. (d) PCA visualization of the retained keys for each KV cache eviction method. We present distributions of additional heads and layers in Section A.3.
Figure 5: Distribution of $\log\left(\textrm{det}\left(KK^{T}\right)\right)$ from the Qasper dataset in LongBench. Larger values mean more of the key space is spanned by the key cache. KeyDiff retains keys that span a greater volume of the ambient space than TOVA or sink attention.
4 Experiment
In this section, we empirically demonstrate the effectiveness of KeyDiff. We begin with a description of competing, state-of-the-art eviction methods, followed by a detailed description of the evaluation setup. We then test each eviction policy on the LongBench task suite (Section 4.2) and report results in Table 1 followed by an ablation study (Section 4.3) of important properties of KeyDiff. Our findings can be summarized as follows:
  • KeyDiff outperforms competing eviction policies on the Needle-In-A-Haystack benchmark (Kamradt, 2023) (Section 4.1).
  • KeyDiff outperforms competing eviction policies with block size $B=128$ on LongBench, achieving an 1.5% accuracy drop with a 6K cache budget ($\sim$33compression rate) and $\leq$ .04% with a 8k cache budget ($\sim$23compression rate) with Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Section 4.2).
  • KeyDiff, utilizing negative cosine similarity as the eviction criteria and the mean of cached keys as the anchor vector, generally outperforms other variants of KeyDiff that use different similarity metrics or anchor vectors. (Section 4.3).
Experimental Setup
We apply several cache eviction methods to several decoder-only transformer-based language models, including Llama $3.1$-$8$B-Instruct (Dubey et al., 2024), Llama $3.2$-$3$B-Instruct, and Qwen $2.5$-$3$B/7B-Instruct. We evaluate these models using H2O (Zhang et al., 2024b), TOVA (Oren et al., 2024), SnapKV (Li et al., 2024), and StreamingLLM (Xiao et al., 2024), (or “sink attention”) cache eviction policies, along with the eviction-free model performance as a baseline. We simulate a resource constrained environment by processing prompts and generating responses using Equation 6, with a block size of $B=128$ for prompt processing and $B=1$ for token generation using greedy decoding for all experiments. We denote the cache budgets of 2048, 4096, 6144 and 8192 as 2K, 4K, 6K and 8K, respectively.
4.1 Needle In a Haystack
To compare the impact of various cache eviction policies on fact retrieval, we conduct the “Needle In a Haystack” test (Liu et al., 2024a; Kamradt, 2023). This test embeds specific information (“needle”) at different points within a body of unrelated text (“haystack”); finding and retaining the needle is challenging for eviction policies, which can’t know during eviction what information must be retained. The results are shown in Figure 6, where we show the recall accuracy of Llama3.23.23.23.2-$3$B-Instruct across different document lengths (x-axis) and needle depths (y-axis) with a cache size of $6$K. KeyDiff performs similarly to TOVA and sink attention for shorter documents and outperforms both methods as the document length increases.
(a) Sink
(b) TOVA
(c) KeyDiff
Figure 6: Accuracy across document length and needle depth for needle in a haystack test. Cache size is $6$K with $B=128$.
Table 1: Llama-3.1-8B/3.2-3B-Instruct LongBench results with $B=128$ (Higher is better). We highlight the best and second best methods within a given budget with bold and underline. We omit Chinese dataset results and other model results due to space limit. The full evaluation results are in Table 11. †: A subset of samples (183/200) were evaluated due to OOM errors.
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code
Narrative QA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PR-en Lcc RB-P Avg.
Llama3.1-8B 30.05 47.00 56.12 57.33 47.81 32.25 34.86 25.32 27.02 73.00 91.61 43.37 8.33 99.50 61.66 51.94 49.20
H2O 2K 1.74 21.15 25.33 26.11 24.15 8.78 2.17 2.70 16.78 44.00 29.36 7.62 2.25 5.88 40.15 12.14 16.89
4K 4.07 36.16 36.00 33.52 32.87 17.78 6.66 5.95 24.09 55.00 47.65 17.41 4.00 24.50 54.85 21.43 26.37
6K 8.52 43.31 44.80 40.03 42.46 21.68 11.85 8.78 26.03 62.00 56.39 25.72 5.75 45.50 58.62 29.53 33.19
8K 13.85 44.94 47.81 43.64 44.90 23.65 18.78 11.35 26.49 69.50 69.05 33.41 5.25 62.50 59.74 36.26 38.20
TOVA 2K 22.57 37.26 39.43 45.74 34.48 14.77 28.87 21.17 26.95 62.50 90.73 42.74 0.00 18.00 62.68 52.48 37.52
4K 22.68 44.55 47.87 46.76 44.54 20.56 30.95 22.13 26.96 61.50 90.56 43.27 3.00 43.50 61.62 53.40 41.49
6K 24.59 45.93 53.92 55.09 47.43 25.07 32.33 24.10 27.00 68.50 90.81 43.89 4.25 67.00 61.50 52.39 45.24
8K 24.86 46.78 54.83 54.52 49.00 26.40 33.44 24.76 27.00 71.00 91.11 43.29 6.25 87.00 61.49 51.79 47.09
Sink 2K 21.83 34.27 29.24 38.64 29.50 12.59 28.51 20.21 26.62 65.00 89.46 42.20 2.00 25.50 64.95 59.54 36.88
4K 22.94 43.01 39.08 44.04 41.39 19.09 31.08 21.57 26.78 70.00 91.53 42.29 3.00 38.50 62.12 58.84 40.95
6K 25.41 47.40 44.13 47.39 45.73 21.90 32.53 22.19 26.87 72.00 91.25 43.41 3.08 52.50 62.22 56.24 43.39
8K 23.53 46.63 48.68 49.61 47.16 21.14 33.10 23.20 26.92 72.00 91.29 43.79 3.25 66.00 62.18 56.43 44.68
SnapKV 2K 21.81 37.22 37.19 46.10 35.42 16.53 29.83 21.05 26.77 61.00 88.84 42.56 4.03 51.50 62.37 51.45 39.60
4K 24.79 44.22 47.30 48.49 46.73 20.55 32.19 22.68 26.95 67.50 90.98 43.14 5.17 89.50 61.44 51.20 45.18
6K 24.10 45.57 50.44 53.12 48.41 24.27 33.43 23.53 27.03 71.50 92.28 43.58 5.25 98.00 61.32 52.16 47.12
8K 25.15 46.55 53.39 56.00 48.75 27.82 33.67 24.85 27.01 72.50 91.78 43.54 5.08 100.00 61.48 51.41 48.06
KeyDiff 2K 26.64 41.73 50.99 51.59 46.47 22.84 29.02 23.86 26.76 66.50 85.92 39.26 3.17 96.00 59.17 39.42 44.33
4K 28.70 45.62 56.06 54.58 49.31 28.25 32.30 25.03 27.07 70.00 90.85 42.84 4.21 99.00 60.80 48.00 47.66
6K 29.90 46.33 55.11 56.80 49.50 31.52 33.44 24.58 26.98 72.00 90.99 43.10 5.27 99.50 61.40 49.70 48.51
8K 33.57 46.77 55.48 56.87 49.37 30.88 34.17 25.12 27.01 72.50 92.28 42.81 5.83 99.50 61.48 50.90 49.03
Llama3.2-3B 23.76 40.23 50.09 50.69 42.29 26.84 33.09 24.30 25.21 72.50 90.11 42.58 3.00 96.50 56.22 56.52 45.87
H2O 2K 1.63 19.96 20.20 18.02 19.56 2.88 0.78 1.55 15.97 41.00 21.97 9.83 0.50 0.50 39.71 13.91 14.25
4K 2.92 31.94 33.23 24.49 28.08 7.55 5.44 6.30 22.77 53.00 38.85 20.33 1.50 7.50 51.23 22.94 22.38
6K 4.62 38.81 39.06 34.66 35.52 15.21 10.51 10.01 24.25 61.50 53.23 27.37 0.50 13.00 54.55 32.29 28.44
8K 9.65 39.66 43.20 38.09 40.41 21.46 17.80 13.28 24.67 70.00 64.30 32.19 2.00 24.50 55.00 39.09 33.46
TOVA 2K 17.14 30.14 32.44 35.96 30.05 13.08 26.15 19.70 25.04 56.50 87.81 40.48 2.50 11.50 55.51 52.36 33.52
4K 20.52 39.53 42.47 44.12 38.42 18.22 29.36 21.36 24.96 63.50 88.98 41.50 3.00 23.50 55.72 56.66 38.24
6K 20.22 39.78 45.86 49.08 41.54 20.43 30.50 22.17 25.11 66.50 89.00 42.50 4.00 46.50 55.57 57.53 41.02
8K 21.08 40.67 49.07 48.69 41.93 23.05 31.64 22.85 25.21 69.00 89.25 42.19 2.50 71.00 55.77 57.47 43.21
Sink 2K 16.85 30.69 26.58 33.26 25.27 13.82 26.74 19.15 25.15 65.00 86.17 40.79 1.50 19.50 56.65 52.73 33.74
4K 19.46 38.61 36.22 41.97 35.84 13.37 29.34 20.19 25.06 71.00 88.06 41.31 2.50 35.50 56.48 52.43 37.96
6K 19.33 40.29 37.95 46.48 40.29 15.31 30.43 21.35 25.14 71.50 88.93 42.04 3.50 47.00 56.55 54.11 40.01
8K 20.15 40.02 41.94 48.15 42.24 16.01 31.64 22.10 25.20 73.00 89.26 42.37 3.50 62.50 56.86 56.63 41.97
SnapKV 2K 17.38 31.37 31.48 37.77 30.05 11.54 27.03 19.93 24.97 59.00 88.13 40.48 3.50 32.50 56.32 55.91 35.46
4K 19.85 39.22 39.86 46.70 37.98 16.64 29.79 21.21 25.01 65.50 89.35 40.95 2.50 62.50 55.74 56.88 40.60
6K 20.83 39.65 44.48 49.30 40.18 20.28 31.27 22.73 25.09 69.00 89.95 41.47 4.00 85.00 55.69 57.82 43.55
8K 20.49 40.80 48.16 48.78 41.65 24.79 31.81 23.46 25.17 70.00 90.17 41.99 5.00 94.00 55.77 57.29 44.96
KeyDiff 2K 18.29 36.65 45.44 46.09 35.41 13.79 28.16 21.45 25.01 60.00 85.24 37.00 1.00 60.50 54.13 42.01 38.14
4K 22.34 40.60 49.15 50.14 40.30 21.65 31.38 23.44 25.06 66.50 87.92 41.41 2.50 88.50 55.55 52.24 43.67
6K 22.29 40.68 50.14 51.74 42.19 24.83 32.39 23.53 25.19 71.00 90.02 42.00 3.00 95.00 55.86 54.39 45.27
8K 22.41 40.77 50.10 49.83 43.58 28.09 32.78 23.60 25.17 72.00 90.17 42.46 3.50 96.50 55.85 55.65 45.78
4.2 LongBench
LongBench (Bai et al., 2024) is a bilingual, multi-task benchmark suite for LLMs, providing a comprehensive stress test for long prompt inputs. LongBench is useful to evaluate cache eviction methods in a resource constrained environments with a fixed memory budget: 51% of prompts are longer than the largest KV cache size of 8K. For cache budgets of 6k and 8k tokens, prompts in LongBench are compressed by 33% and 23% respectively on average, (see Section B.3 for more detail.)
Table 1 summarizes the evaluation results of Llama 3.1-8B-Instruct and Llama 3.2-3B-Instruct on the English subset of LongBench with 2K, 4K, 6K, and 8K cache budgets using various eviction policies with block prompt processing enabled with $B=128$. We present the full evaluation results in Table 11, and those of Qwen 2.5-3B/7B-Instruct in Table 9.
As shown in Table 1, KeyDiff outperforms other eviction strategies across most tasks, even demonstrating better performance with smaller cache budgets. KeyDiff shows significant a improvement on the PassageRetrieval-en (PR-en) dataset, which tests whether long-term dependencies within a long prompt can be correctly recognized (Bai et al., 2024), while achieving near full-context model performance even with the smallest budget. We observed similar trends in the full LongBench task suite as shown in Table 11.
Results for Full Prompt Processing
We also evaluate the methods in more standard inference environments without block prompt processing (i.e., setting the block size of the prompt to $B=\infty$) and summarize the results in Table 7. KeyDiff exhibits similar or better performance compared to competing methods. Notably, the attention-based methods (e.g., H2O, TOVA, and SnapKV) show significant performance improvements over the $B=128$ case. This result supports our hypothesis: an eviction scheme robust to changes in the scope of comparison among tokens is essential in memory constrained environments where token-wise attention weight can’t be fully materialized.
4.3 Ablation Study
We evaluate the design choices of KeyDiff, including the similarity metrics and the choice of the anchor vector, and validate the efficacy of KeyDiff.
Table 2: Ablation study results about anchor vector. Average of Full Longbench results for Llama 3.2-3B-Instruct. KeyDiff results match Table 11. (Higher is better)
2K 4K 6K 8K
KeyDiff 35.09 40.28 41.88 42.40
$\operatorname{Pairwise}$ 35.24 40.61 41.87 42.45
$\operatorname{Median}$ 35.43 40.67 41.89 42.26
Selecting the Anchor Vector
We have mainly evaluated KeyDiff using the method described in Equation 8. Scores to determine eviction are measured via cosine similarity with an anchor vector which can be computed in several ways. We run LongBench on Llama3.2-3B-Instruct with eviction policies using the following anchor choices: pairwise cosine similarity from Equation 7, denoted $\operatorname{Pairwise}$; KeyDiff, using the mean of all normalized keys as an anchor, and using the median of keys as an anchor, denoted $\operatorname{Median}$. Table 2 summarizes the average LongBench accuracy for the different methods to Llama 3.2-3B-Instruct. KeyDiff shows similar average scores to $\operatorname{Pairwise}$. Additionally, KeyDiff and $\operatorname{Median}$ show similar scores, demonstrating that KeyDiff is robust to the selection of the anchor design.
Table 3: Ablation study results about distance metric. Average of Full Longbench results of Llama 3.2-3B-Instruct. KeyDiff results match Table 11. (Higher is better)
2K 4K 6K 8K
KeyDiff 35.09 40.28 41.88 42.40
$\operatorname{DotProd}$ 30.14 38.01 41.09 42.23
$\operatorname{Euclidean}$ 13.68 21.06 25.91 29.53
Selecting the Similarity Metric
We use cosine similarity as the scoring metric for eviction in KeyDiff based on our discussion in Section 2.2. This could be replaced with other metrics like the dot product or Euclidean distance. We evaluate KeyDiff variants using dot product and Euclidean distance as the similarity metric, denoted $\operatorname{DotProd}$ and $\operatorname{Euclidean}$ respectively, and report the results in Table 3. KeyDiff and $\operatorname{DotProd}$ show similar performance for 6K and 8K budgets. However, KeyDiff outperforms $\operatorname{DotProd}$ for smaller cache sizes. This implies that considering both the direction and the magnitude of the keys to compute similarity are important for identifying the tokens to evict. On the other hand, $\operatorname{Euclidean}$ shows a significant performance drop relative to KeyDiff.
5 Related Work
Sparse Attention
LLMs often exhibit high attention sparsity, where a small subset of keys receives a significant proportion of attention scores. This characteristic allows sparse approximation techniques to reduce the computational cost of attention. Similar to PagedAttention (Kwon et al., 2023), Tang et al. (2024) estimates the importance of a page (a contiguous set of keys) to a given query, whereas Rehg (2024) further refined the budgets in a per-head manner. On the contrary, sample-based methods (Zhu et al., 2024; Ribar et al., ) attempt to approximate token importance by inspecting the attention scores from the last few queries or certain query channel dimensions. Despite their effectiveness in reducing computational costs, these methods do not address the memory overhead of the KV cache, which typically retains all KVs.
KV Cache Compression
Different approaches to compress the KV cache include architecture modification such as GQA (Ainslie et al., 2023), which shares a KV cache across a small number of heads. Other techniques to compress the KV cache include quantization such as in (Hooper et al., 2024; Liu et al., 2024b; Zhang et al., 2024a) in which the authors use various techniques to take advantage of existing patterns to efficiently quantize and compress the KV cache. More related to our work (Yang et al., 2024b) uses a scoring mechanism to determine the precision of the quantization for different tokens.
Token Eviction Methods
Unlike the sparse attention and KV cache compression methods, eviction methods evict KVs from the cache to reduce the size of the KV cache. As discussed in Section 2.3, the majority of the token eviction methods employ their own rules to decide the importance of the tokens by manipulating the attention score $A$. For example, by appropriately choosing the aggregation functions $\phi(A)$ of Equation 4, we can obtain existing attention-based eviction methods as discussed in Section A.1. Attention-based eviction may be a better choice when the entire prompt is being processed at once, as the eviction can be done by assessing the importance of all tokens simultaneously. However, computing the full attention score of long prompts could be prohibitively expensive in resource-constrained environments.
6 Conclusion
Inspired by our observation that distinctive keys tend to have high attention scores, we propose KeyDiff, a training-free KV cache eviction method based on key similarity that enables large language models to operate in memory and compute constrained environments. We justify KeyDiff by showing that it computes the optimal solution for an optimization problem minimizing the pairwise cosine similarity among keys in the KV cache. KeyDiff significantly outperforms state-of-the-art KV cache eviction methods under similar memory constraints, with only a 1.5% and 0.04% accuracy drop from the non-evicting baseline while achieving 33% and 23% KV cache memory reduction on LongBench.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Supplementary Material
[…]
Appendix A Extended Related Work
A.1 Attention-based eviction methods
In this section, we provide a unified framework to understand prominent attention-based eviction methods. As mentioned in Equation 5, we can specify attention-based eviction methods under the unified framework with proper selection of the aggregation function $\phi(A)$ as follows:
  • TOVA (Oren et al., 2024): $\phi^{\mathrm{TOVA}}(A)=A_{-1,:}$,
  • H2O (Zhang et al., 2024b): $\phi^{\mathrm{H2O}}(A)=A_{\mathrm{prev}}+A^{\top}1$,
  • SnapKV (Xiao et al., 2024): $\phi^{\mathrm{SnapKV}}(A)=(A^{\top}1)\ast K$, where $K$ is a vector of $\frac{1}{k}$ and $k$ is the kernel size of average smoothing.
A.2 Runtime and Memory Complexity
We analyze the runtime and memory complexity for the prominent KV cache eviction algorithms in Table 4. For a given block size $B$ and cache budget $N$, KeyDiff requires $\mathcal{O}(N+B)$ runtime and memory. The same holds true for TOVA, since it only requires computing the bottom row of $A$. Sink attention retains the $k$ first tokens in the input sequence, followed by a sliding window of size $L$, resulting in $\mathcal{O}(k+L)=O(N)$ memory and runtime, since $k+L$ equals the chosen cache budget. SnapKV computes attention over a sliding window of size $L$ against $N+B$ keys from the incoming block and the key cache, so the memory and runtime complexity is $\mathcal{O}((N+B)L)$. H2O accumulates attention weights over all tokens, and computes the attention over the current block, so it will require $\mathcal{O}(NB+B^{2})$ memory overhead and runtime. We summarize these details in Table 4
Table 4: Runtime and memory complexity of token eviction methods.
Runtime Complexity Memory Complexity
KeyDiff $\mathcal{O}(N+B)$ $\mathcal{O}(N+B)$
TOVA $\mathcal{O}(N+B)$ $\mathcal{O}(N+B)$
H2O $\mathcal{O}(NB+B^{2})$ $\mathcal{O}(NB+B^{2})$
SnapKV $\mathcal{O}((N+B)L)$ $\mathcal{O}((N+B)L)$
Sink $\mathcal{O}(N)$ $\mathcal{O}(1)$
A.2.1 FLOP count of KeyDiff
Regarding the FLOP count of KeyDiff: the bulk of the computation in KeyDiff (neglecting the ‘topk‘ operator) is the following two expressions:
  • $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\frac{k_{i}}{\|k_{i}\|}$
  • $s_{i}=\textrm{CosSim}(\mu(\hat{K}),k_{i})=\frac{\mu(\hat{K})\cdot k_{i}}{\max(\|\mu(\hat{K})\|\cdot\|k_{i}\|,\,\epsilon)},\quad i=1,\dots,n$
We will count the total number of additions, multiplications, square roots and divisions required by KeyDiff separately, since division and square root implementation are hardware-dependent, then assign weights to each operation at the end for a final count. Norms are assumed to be 2-norms. We count the FLOPs required for each operation as follows:
  • $\|k_{i}\|=\sqrt{k_{i}\cdot k_{i}}$: since $k_{i}\in\mathbb{R}^{d}$: $d$ multiplications, $d-1$ additions, one square root. Repeating for each $i$, this contributes $nd$ multiplications, $n(d-1)$ additions, and $n$ square roots.
  • $\frac{k_{i}}{c}$: naively, $d$ divisions, but can be rewritten as one division and $d$ multiplications. Repeating for each $i$, this contributes $nd$ multiplications and $n$ divisions.
  • $\frac{1}{n}\sum_{i=1}^{n}c_{i}$, for $c\in\mathbb{R}^{d}$: one division, $(n-1)d$ additions.
Combining the above, we can compute the anchor vector using $2nd$ multiplications, $2nd-n-d$ additions, $n$ square roots and $n+1$ divisions.
To compute the cosine similarity score, we have from above that $\mu(\hat{K})\cdot k_{i}$ requires $nd$ multiplications and $n(d-1)$ additions. Also from above, we have that computing $\|\mu(\hat{K})\|$ requires $d$ multiplications, $d-1$ additions and one square root. We reuse the computation of $\|k_{i}\|$ from the previous step and compute $\|k_{i}\|\|\mu(\hat{K})\|$ in $n$ multiplications and $\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)$ with more $n$ additions (assuming boolean comparison is equals addition in cost). We can then divide through to compute $\frac{\mu(\hat{K})\cdot k_{i}}{\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)}$ with $n$ divisions. Therefore, computing the cosine similarity between the anchor and each key requires $nd+d+n$ multiplications, $nd+d-1$ additions, $n+1$ divisions and one square root.
Adding everything up, KeyDiff requires:
  1. 1.
    $3nd+d+n$ multiplications,
  2. 2.
    $13nd-n-$ 1 additions,
  3. 3.
    $22n+$ 2 divisions,
  4. 4.
    $n+1$ square roots,
If we declare additions and multiplications are roughly one FLOP (i.e. can be computed in one cycle), division is roughly 5 FLOPs and square root is roughly 8 FLOPs, we arrive at a final FLOP count of:
$(3nd+d+n)+(3nd-n-1)+5*(2n+2)+8*(n+1)=(6d+18)n+d+17.$ (A.1)
This is linear in the number of keys $n$ with a small constant, relative to the quadratic complexity of the attention operator.
A.3 Additional PCA Visualizations
In order to demonstrate the phenomena in Figure 4 persists across all heads and layers, we have included several more visualizations as seen in Figure 7, Figure 8, and Figure 9.
(a) Sink Attention
(b) TOVA
(c) KeyDiff
(d) Retained keys only
Figure 7: (a, b, and c) PCA Visualizations in two dimensions of a key cache managed with Sink, TOVA, and KeyDiff. Retained tokens are blue, while evicted tokens are orange. Keys are taken from layer $27$ and head $4$ of Llama3.2-3B-Instruct, and generated using the NarrativeQA dataset. (d) PCA visualization of the retained keys for each KV cache eviction method
(a) Sink Attention
(b) TOVA
(c) KeyDiff
(d) Retained keys only
Figure 8: Keys taken from layer $20$ and head 00 of Llama3.2-3B-Instruct
(a) Sink Attention
(b) TOVA
(c) KeyDiff
(d) Retained keys only
Figure 9: Keys taken from layer $8$ and head $1$ of Llama3.2-3B-Instruct
Appendix B Additional discussion for LongBench
B.1 Empirical Motivation for KeyDiff Setup
To generate Figure 4, we used the first sample from the test split of the narrativeqa task in LongBench to prefill the KV cache of Llama3.2-3B-Instruct with a block size of $B=128$. The KV cache had a maximum size of $4096$ while the sample was much longer, requiring KV eviction. We applied PCA to the key cache and repeated the process for sink attention, TOVA and KeyDiff.
To construct Figure 5, we sample 100 prompts from the Qasper dataset in LongBench (Bai et al., 2024), compute the log determinant of $KK^{T}$ of the keys in the KV caches of each head and layer of Llama 3.2-3B-Instruct using a cache budget of $N=2048$ and a block size of $B=128$, and plot the distribution in Figure 5. We show this key distribution for sink attention, TOVA and KeyDiff.
B.2 Experiment Setup
In this subsection, we provide the experimental setup for KeyDiff and the baselines for the LongBench experiments. The LongBench evaluation is conducted using the default parameters of the LongBench evaluator with predefined prompt templates. The prompt block size is set to $B=128$, and the same $B$ value is used for all experiments.
For TOVA, H2O, ans SnapKV, the set of attention weights computed from a single key cache due to grouped query attention (Ainslie et al., 2023) is aggregated by taking the average over the attention weights. Additionally, only for SnapKV, we apply average smoothing to the attention score with a kernel size of 7 and keep the most recent 32 tokens in the cache, following the suggestion of the original paper. For Sink, we used the first four tokens as the attention sink, following the suggestion of the original paper.
B.3 Longbench dataset statistics
In this section, we provide the length statistics of the Longbench Benchmark and in-depth analysis of compression ratios for the given KV cache budgets, such as 2K, 4K, 6K, and 8K.
Prompt lengths
We measure the number of tokens in the samples using LLama tokenizer (Touvron et al., 2023). As shown in Figure 10, LongBench exhibits variability in sample length from the datasets.
Figure 10: Histograms of sample lengths measured by number of tokens
Compression ratio
The majority of other KV cache eviction studies assume an unconstrained memory footprint. Before they compress the cache by applying an eviction policy, they first set the target compression ratio and evict the appropriate number of KV pairs to satisfy the compression ratio (Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024). On the other hand, we fix the KV cache size and ensure the number of cached tokens is less than or equal to the predefined cache size. Due to these differences, it is less straightforward to set appropriate KV cache budgets to satisfy the target compression ratios. Instead, we provide the average compression ratio, which is defined as:
Average Compression $\text{Average Compression Ratio}=\frac{1}{I}\sum_{i=1}^{I}\frac{N}{L_{i}},$
where $N$ is the KV cache budget, and $L_{i}$ is the length of the $i$-th prompt (sample). We replace the summand with 1 whenever $N\geq L_{i}$, as compression doesn’t occur in that setting.
As summarized in Table 6, 2K cache budgets have a 0.31 average compression ratio, which indicates 69% of input prompts are compressed. Our largest cache budget, 8K, exhibits a 0.77 average compression ratio.
$\leq$ 2K 2K $\leq L\leq$ 4K 4K $\leq L\leq$ 6K 6K $\leq L\leq$ 8K $\geq$ 8K Total
NarrativeQA 0 0 0 8 192 200
Qasper 1 77 83 25 14 200
MultifidelityQA-En 9 31 21 32 57 150
MultifidelityQA-Zh 14 69 47 33 37 200
HotPotQA 1 4 12 12 171 200
2wikimqa 8 17 68 54 53 200
musique 0 0 0 3 197 200
dureader 0 0 0 16 184 200
gov report 0 20 29 45 106 200
qmsum 0 1 17 15 167 200
multi news 99 71 19 6 5 200
vcsum 4 20 18 32 126 200
trec 4 43 41 39 73 200
triviaqa 4 21 15 21 139 200
samsum 6 29 34 17 114 200
lsht 0 0 3 8 189 200
passage count 0 0 3 13 184 200
passage-retrieval-En 0 0 0 0 200 200
passage-retrieval-Zh 0 0 160 10 0 170
lcc 80 86 21 4 9 200
repobench-p 0 25 37 25 113 200
Table 5: Distribution of sample length measured by Llama2 tokenizer
2K 4K 6K 8K
NarrativeQA 0.10 0.20 0.30 0.40
Qasper 0.47 0.82 0.96 0.98
MultifidelityQA-En 0.40 0.66 0.82 0.93
MultifidelityQA-Zh 0.49 0.78 0.91 0.98
HotPotQA 0.19 0.37 0.52 0.66
2wikimqa 0.36 0.65 0.85 0.92
musique 0.13 0.27 0.40 0.53
dureader 0.17 0.34 0.52 0.68
gov report 0.28 0.52 0.70 0.81
qmsum 0.18 0.36 0.52 0.66
multi news 0.81 0.96 0.98 0.53
vcsum 0.27 0.49 0.65 0.68
trec 0.38 0.66 0.84 0.94
triviaqa 0.26 0.46 0.60 0.72
samsum 0.32 0.55 0.71 0.82
lsht 0.13 0.27 0.39 0.52
passage count 0.15 0.31 0.46 0.60
passage-retrieval-En 0.16 0.33 0.49 0.66
passage-retrieval-Zh 0.37 0.74 0.99 1.00
lcc 0.78 0.95 0.98 0.99
repobench-p 0.27 0.52 0.67 0.78
average 0.31 0.53 0.67 0.77
Table 6: Compression ratio of prompts w.r.t. various KV cache budgets
B.4 Additional Results
Table 7: Resource unrestricted LongBench Results (Higher is better). All methods processes the input prompt in parallel (i.e., block size =absent=$\infty$= ∞) and make an eviction decision with all token information in the input. The token eviction is made at every step of generation if the budget exceeds. We highlight the methods showing the best performance within a given budget with boldface. We omit NarrativeQA from evaluation due to higher chance of OOM errors.
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code Qasper MF-en MF-zh HotpotQA 2WikiMQA Musique DuReader GovReport QMSum MultiNews VCSum TREC TriviaQA SAMSum LSHT PCount PR-en PR-zh Lcc RB-P Avg. Llama3.1-8B 47.00 56.12 59.86 57.33 47.81 32.25 35.64 34.86 25.32 27.02 17.28 73.00 91.61 43.37 45.50 8.33 99.50 99.00 61.66 51.94 50.72 H2O 2K 22.75 32.73 25.93 43.56 29.49 0.00 5.35 3.70 4.73 17.42 4.44 46.19 54.88 10.39 12.20 16.13 100.00 37.75 42.44 16.44 26.33 4K 35.49 43.14 39.80 52.82 36.60 10.00 6.11 9.34 7.54 24.61 7.38 54.31 65.83 23.18 16.67 16.39 100.00 61.50 55.96 28.68 34.77 6K 44.05 47.74 47.88 52.67 45.90 11.11 7.97 15.55 13.77 26.58 9.13 60.11 78.42 32.35 13.64 12.04 100.00 94.00 60.23 39.82 40.65 8K 45.93 51.12 55.72 53.09 48.83 13.58 15.63 25.89 17.39 27.18 12.39 67.71 86.91 40.05 13.33 13.73 100.00 99.00 60.82 45.17 44.67 TOVA 2K 44.37 55.47 58.07 59.16 48.26 16.67 26.58 30.54 24.37 26.81 16.66 71.00 91.93 45.29 28.89 9.84 100.00 96.58 61.65 51.83 48.20 4K 46.45 56.23 58.82 60.72 49.96 15.28 29.93 33.52 25.56 27.18 17.40 72.08 91.40 44.49 29.79 13.06 100.00 99.00 61.95 52.19 49.25 6K 46.66 54.26 59.31 55.99 50.26 16.22 34.54 34.35 25.21 27.22 17.34 72.96 91.56 44.82 30.43 14.72 100.00 99.00 62.18 53.33 49.52 8K 46.57 55.44 59.16 59.87 51.77 13.58 35.28 35.21 25.98 27.27 16.80 73.23 90.70 44.71 34.78 13.59 100.00 99.00 61.96 54.32 49.96 Sink 2K 33.33 33.74 34.73 45.37 38.46 20.97 18.31 26.08 21.41 24.98 16.08 67.50 90.00 40.99 21.25 2.50 36.00 18.00 57.08 53.81 35.03 4K 38.57 41.15 46.38 48.07 40.87 22.51 17.50 29.28 21.89 25.12 16.84 71.50 90.52 41.32 24.75 2.50 49.00 22.50 57.94 53.64 38.09 6K 40.41 43.74 52.60 50.08 42.57 22.32 17.71 30.54 22.38 25.25 17.37 72.50 90.77 41.94 26.25 2.50 57.00 21.00 57.14 54.07 39.41 8K 40.53 44.58 54.77 49.13 42.38 23.75 20.35 31.35 22.61 25.28 17.45 72.50 90.77 42.27 27.75 3.00 70.00 20.50 57.13 55.35 40.57 SnapKV 2K 45.01 52.53 56.15 57.54 50.17 25.00 32.35 32.99 25.38 27.24 18.14 70.85 89.48 39.71 26.53 15.86 98.57 87.04 60.56 51.92 48.15 4K 46.28 55.05 59.57 55.49 49.70 24.69 35.09 33.78 26.63 27.15 17.48 72.68 90.28 42.59 28.57 13.38 98.08 98.50 61.14 52.05 49.41 6K 46.82 55.83 59.07 59.68 50.27 22.22 35.82 34.89 25.17 27.24 17.51 72.31 90.76 44.31 23.26 14.00 100.00 99.00 62.13 54.43 49.74 8K 46.72 55.33 59.79 57.18 51.51 15.28 34.83 35.28 25.71 27.24 17.40 71.96 90.46 45.08 23.26 10.92 100.00 98.99 61.76 54.80 49.17 KeyDiff 2K 44.58 53.88 53.65 57.40 47.66 14.89 33.44 29.97 25.88 26.95 16.10 72.00 92.27 43.51 31.82 11.38 100.00 95.92 59.26 45.66 47.81 4K 46.05 54.87 57.52 59.04 49.97 13.58 36.51 33.17 25.92 27.12 16.85 73.00 90.13 43.87 31.82 13.28 100.00 97.67 60.57 50.49 49.07 6K 46.61 54.49 59.19 59.70 50.03 12.99 36.04 34.42 26.67 27.27 17.47 73.33 90.94 44.62 33.33 11.83 100.00 99.00 61.29 53.20 49.62 8K 46.55 55.11 59.28 57.48 50.51 13.58 35.01 34.81 25.87 27.22 17.15 72.59 91.22 44.69 27.27 11.20 100.00 99.00 61.95 53.62 49.21 Llama3.2-3B 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 43.66 H2O 2K 20.21 24.23 18.98 28.28 23.51 9.88 10.40 1.43 4.13 16.14 3.05 48.00 41.36 12.23 16.00 4.81 1.52 0.50 40.16 17.58 17.12 4K 32.58 33.87 35.46 35.29 27.46 17.84 12.25 6.96 9.26 23.10 5.17 56.00 59.11 21.77 13.70 4.95 13.71 7.25 51.49 29.58 24.84 6K 38.61 44.28 46.68 42.34 36.93 11.61 15.57 13.52 13.03 24.44 7.74 63.50 72.81 30.46 12.86 4.76 63.00 19.00 54.28 39.85 32.76 8K 40.23 44.83 52.90 46.13 39.78 14.74 20.79 22.65 17.06 24.77 10.55 70.50 83.55 35.08 15.07 4.04 81.91 20.00 55.35 45.83 37.29 TOVA 2K 38.22 48.83 54.09 48.08 42.21 14.81 27.89 28.71 23.27 24.94 15.64 70.50 89.47 42.97 22.22 5.88 95.92 18.50 56.14 55.68 41.20 4K 40.55 50.77 56.10 54.26 42.68 17.08 31.66 31.09 23.44 25.37 16.00 71.50 89.77 43.00 21.92 5.88 95.50 18.50 56.33 56.52 42.40 6K 40.57 50.12 56.62 52.67 43.12 19.44 34.78 32.91 23.88 25.30 15.87 72.50 89.07 42.73 24.00 4.35 94.97 20.00 56.30 57.38 42.83 8K 40.86 49.79 56.31 52.91 42.03 15.34 36.74 33.34 24.11 25.30 16.15 72.50 89.40 42.39 24.32 5.77 95.50 20.00 56.27 56.85 42.79 Sink 2K 33.33 33.74 34.73 45.37 38.46 20.97 18.31 26.08 21.41 24.98 16.08 67.50 90.00 40.99 21.25 2.50 36.00 18.00 57.08 53.81 35.03 4K 38.57 41.15 46.38 48.07 40.87 22.51 17.50 29.28 21.89 25.12 16.84 71.50 90.52 41.32 24.75 2.50 49.00 22.50 57.94 53.64 38.09 6K 40.41 43.74 52.60 50.08 42.57 22.32 17.71 30.54 22.38 25.25 17.37 72.50 90.77 41.94 26.25 2.50 57.00 21.00 57.14 54.07 39.41 8K 40.53 44.58 54.77 49.13 42.38 23.75 20.35 31.35 22.61 25.28 17.45 72.50 90.77 42.27 27.75 3.00 70.00 20.50 57.13 55.35 40.57 KeyDiff 2K 38.15 49.58 51.15 48.73 42.04 20.24 33.47 29.13 23.91 25.10 14.47 70.50 88.31 42.28 20.55 4.90 89.00 17.50 55.21 48.73 40.65 4K 40.55 51.38 55.88 51.56 41.88 18.78 35.93 31.63 24.15 25.38 15.80 71.50 89.98 42.55 19.18 4.95 95.85 21.00 55.79 55.55 42.46 6K 40.54 51.04 55.42 51.78 41.71 15.26 36.60 32.80 24.67 25.34 16.32 72.50 90.58 43.19 22.22 5.21 97.42 20.50 55.87 56.28 42.76 8K 40.66 50.93 56.16 53.52 42.07 17.79 37.08 33.31 24.27 25.31 16.00 72.50 89.48 42.52 25.33 5.00 96.46 20.00 56.20 56.80 43.07
Table 8: Full Llama-3.1-8B/3.2-3B-Instruct LongBench Results with $B=128$ (Higher is better). We highlight the methods showing the best performance within a given budget with boldface. †: A subset of samples were evaluated due to OOM errors (183/200 samples are evaluated).
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code Narrative QA Qasper MF-en MF-zh HotpotQA 2WikiMQA Musique DuReader GovReport QMSum MultiNews VCSum TREC TriviaQA SAMSum LSHT PCount PR-en PR-zh Lcc RB-P Avg. Llama3.1-8B 30.05 47.00 56.12 59.86 57.33 47.81 32.25 35.64 34.86 25.32 27.02 17.28 73.00 91.61 43.37 45.50 8.33 99.50 99.00 61.66 51.94 49.74 H2O 2K 1.74 21.15 25.33 21.65 26.11 24.15 8.78 5.90 2.17 2.70 16.78 3.97 44.00 29.36 7.62 14.50 2.25 5.88 4.00 40.15 12.14 15.25 4K 4.07 36.16 36.00 38.02 33.52 32.87 17.78 5.68 6.66 5.95 24.09 6.03 55.00 47.65 17.41 18.50 4.00 24.50 31.25 54.85 21.43 24.83 6K 8.52 43.31 44.80 46.24 40.03 42.46 21.68 7.33 11.85 8.78 26.03 7.82 62.00 56.39 25.72 18.00 5.75 45.50 90.00 58.62 29.53 33.35 8K 13.85 44.94 47.81 56.14 43.64 44.90 23.65 11.01 18.78 11.35 26.49 9.96 69.50 69.05 33.41 19.50 5.25 62.50 98.67 59.74 36.26 38.40 TOVA 2K 22.57 37.26 39.43 36.96 45.74 34.48 14.77 16.98 28.87 21.17 26.95 16.21 62.50 90.73 42.74 18.75 0.00 18.00 32.00 62.68 52.48 34.35 4K 22.68 44.55 47.87 51.16 46.76 44.54 20.56 22.50 30.95 22.13 26.96 16.75 61.50 90.56 43.27 25.25 3.00 43.50 84.00 61.62 53.40 41.12 6K 24.59 45.93 53.92 55.45 55.09 47.43 25.07 27.68 32.33 24.10 27.00 16.91 68.50 90.81 43.89 29.00 4.25 67.00 98.67 61.50 52.39 45.31 8K 24.86 46.78 54.83 57.95 54.52 49.00 26.40 31.15 33.44 24.76 27.00 17.33 71.00 91.11 43.29 33.25 6.25 87.00 98.67 61.49 51.79 47.23 Sink 2K 21.83 34.27 29.24 32.82 38.64 29.50 12.59 16.18 28.51 20.21 26.62 15.54 65.00 89.46 42.20 22.25 2.00 25.50 32.50 64.95 59.54 33.78 4K 22.94 43.01 39.08 46.16 44.04 41.39 19.09 16.54 31.08 21.57 26.78 16.73 70.00 91.53 42.29 29.25 3.00 38.50 71.00 62.12 58.84 39.76 6K 25.41 47.40 44.13 52.78 47.39 45.73 21.90 17.55 32.53 22.19 26.87 17.05 72.00 91.25 43.41 33.75 3.08 52.50 98.00 62.22 56.24 43.49 8K 23.53 46.63 48.68 55.77 49.61 47.16 21.14 19.54 33.10 23.20 26.92 16.91 72.00 91.29 43.79 37.00 3.25 66.00 99.00 62.18 56.43 44.91 SnapKV 2K 21.81 37.22 37.19 38.29 46.10 35.42 16.53 16.37 29.83 21.05 26.77 16.16 61.00 88.84 42.56 21.75 4.03 51.50 81.17 62.37 51.45 38.45 4K 24.79 44.22 47.30 50.27 48.49 46.73 20.55 22.04 32.19 22.68 26.95 16.95 67.50 90.98 43.14 25.00 5.17 89.50 96.67 61.44 51.20 44.46 6K 24.10 45.57 50.44 55.27 53.12 48.41 24.27 27.46 33.43 23.53 27.03 16.84 71.50 92.28 43.58 27.00 5.25 98.00 99.00 61.32 52.16 46.65 8K 25.15 46.55 53.39 57.65 56.00 48.75 27.82 32.66 33.67 24.85 27.01 17.37 72.50 91.78 43.54 33.75 5.08 100.00 98.67 61.48 51.41 48.05 KeyL2Norm 2K 8.66 36.63 41.70 37.70 33.75 32.25 5.39 17.73 19.64 14.96 26.69 11.02 63.00 58.94 28.45 22.50 3.05 17.75 20.13 52.40 25.63 27.52 4K 15.38 44.06 51.75 47.52 50.22 45.56 18.44 27.61 29.50 22.27 26.93 13.44 69.50 79.41 37.50 27.00 4.50 58.00 80.50 58.82 35.08 40.14 6K 21.75 45.63 55.06 53.77 52.93 47.70 25.63 32.17 32.66 24.85 26.94 15.12 70.00 86.89 40.51 34.50 5.25 75.00 98.00 60.98 43.14 45.17 8K 25.12 45.70 56.02 56.57 58.14 47.77 30.29 34.09 33.81 24.89 26.94 15.89 71.50 89.26 41.34 39.00 7.25 87.00 99.00 62.05 48.28 47.61 KeyDiff 2K 26.64 41.73 50.99 51.18 51.59 46.47 22.84 32.37 29.02 23.86 26.76 14.81 66.50 85.92 39.26 42.25 3.17 96.00 96.25 59.17 39.42 45.06 4K 28.70 45.62 56.06 56.83 54.58 49.31 28.25 33.06 32.30 25.03 27.07 16.32 70.00 90.85 42.84 44.50 4.21 99.00 97.67 60.80 48.00 48.14 6K 29.90 46.33 55.11 59.00 56.80 49.50 31.52 34.97 33.44 24.58 26.98 16.80 72.00 90.99 43.10 47.00 5.27 99.50 99.00 61.40 49.70 49.19 8K 33.57 46.77 55.48 59.16 56.87 49.37 30.88 34.54 34.17 25.12 27.01 17.13 72.50 92.28 42.81 46.50 5.83 99.50 98.67 61.48 50.90 49.55 Llama3.2-3B 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 H2O 2K 1.63 19.96 20.20 15.20 18.02 19.56 2.88 6.47 0.78 1.55 15.97 3.11 41.00 21.97 9.83 11.75 0.50 0.50 0.00 39.71 13.91 12.60 4K 2.92 31.94 33.23 33.25 24.49 28.08 7.55 10.10 5.44 6.30 22.77 4.81 53.00 38.85 20.33 15.50 1.50 7.50 6.25 51.23 22.94 20.38 6K 4.62 38.81 39.06 45.17 34.66 35.52 15.21 13.36 10.51 10.01 24.25 6.66 61.50 53.23 27.37 15.25 0.50 13.00 19.50 54.55 32.29 26.43 8K 9.65 39.66 43.20 52.60 38.09 40.41 21.46 18.55 17.80 13.28 24.67 9.12 70.00 64.30 32.19 17.00 2.00 24.50 21.50 55.00 39.09 31.15 TOVA 2K 17.14 30.14 32.44 31.64 35.96 30.05 13.08 9.62 26.15 19.70 25.04 15.47 56.50 87.81 40.48 16.75 2.50 11.50 6.50 55.51 52.36 29.35 4K 20.52 39.53 42.47 45.80 44.12 38.42 18.22 17.76 29.36 21.36 24.96 16.60 63.50 88.98 41.50 18.75 3.00 23.50 15.00 55.72 56.66 34.56 6K 20.22 39.78 45.86 52.93 49.08 41.54 20.43 24.78 30.50 22.17 25.11 16.37 66.50 89.00 42.50 21.00 4.00 46.50 20.50 55.57 57.53 37.71 8K 21.08 40.67 49.07 55.17 48.69 41.93 23.05 31.02 31.64 22.85 25.21 16.55 69.00 89.25 42.19 22.50 2.50 71.00 21.50 55.77 57.47 39.91 Sink 2K 16.85 30.69 26.58 27.32 33.26 25.27 13.82 9.38 26.74 19.15 25.15 15.88 65.00 86.17 40.79 19.50 1.50 19.50 8.50 56.65 52.73 29.54 4K 19.46 38.61 36.22 41.68 41.97 35.84 13.37 9.86 29.34 20.19 25.06 16.44 71.00 88.06 41.31 21.75 2.50 35.50 16.00 56.48 52.43 33.96 6K 19.33 40.29 37.95 49.68 46.48 40.29 15.31 11.10 30.43 21.35 25.14 16.64 71.50 88.93 42.04 23.50 3.50 47.00 19.50 56.55 54.11 36.22 8K 20.15 40.02 41.94 53.57 48.15 42.24 16.01 14.76 31.64 22.10 25.20 16.50 73.00 89.26 42.37 26.25 3.50 62.50 20.50 56.86 56.63 38.25 SnapKV 2K 17.38 31.37 31.48 29.65 37.77 30.05 11.54 9.66 27.03 19.93 24.97 15.97 59.00 88.13 40.48 16.25 3.50 32.50 9.00 56.32 55.91 30.85 4K 19.85 39.22 39.86 47.33 46.70 37.98 16.64 16.88 29.79 21.21 25.01 16.74 65.50 89.35 40.95 18.25 2.50 62.50 22.50 55.74 56.88 36.73 6K 20.83 39.65 44.48 51.84 49.30 40.18 20.28 25.32 31.27 22.73 25.09 16.81 69.00 89.95 41.47 18.75 4.00 85.00 20.50 55.69 57.82 39.52 8K 20.49 40.80 48.16 55.44 48.78 41.65 24.79 30.40 31.81 23.46 25.17 16.44 70.00 90.17 41.99 22.00 5.00 94.00 21.50 55.77 57.29 41.20 KeyL2Norm 2K 7.67 30.39 31.85 30.64 29.47 25.76 7.41 14.17 15.36 12.42 24.20 7.91 48.00 50.99 23.09 17.50 2.00 7.50 5.00 48.92 26.32 22.22 4K 12.92 37.59 44.71 43.85 38.89 33.42 12.41 22.42 24.63 19.27 24.77 11.37 63.00 72.51 31.75 19.00 3.87 9.50 11.00 55.82 40.08 30.13 6K 13.02 40.55 48.17 50.87 43.18 40.17 17.10 31.29 28.99 21.47 25.08 13.30 65.00 79.61 37.16 21.50 2.50 46.50 18.50 56.49 47.05 35.60 8K 15.72 40.54 47.88 54.29 49.29 43.79 22.22 33.18 31.86 22.50 25.19 14.28 70.00 84.92 39.45 23.00 1.50 69.00 20.50 56.82 50.73 38.89 KeyDiff 2K 18.29 36.65 45.44 47.47 46.09 35.41 13.79 28.89 28.16 21.45 25.01 13.56 60.00 85.24 37.00 24.88 1.00 60.50 12.00 54.13 42.01 35.09 4K 22.34 40.60 49.15 52.56 50.14 40.30 21.65 32.46 31.38 23.44 25.06 15.28 66.50 87.92 41.41 27.50 2.50 88.50 19.50 55.55 52.24 40.28 6K 22.29 40.68 50.14 54.51 51.74 42.19 24.83 34.64 32.39 23.53 25.19 15.88 71.00 90.02 42.00 28.75 3.00 95.00 21.50 55.86 54.39 41.88 8K 22.41 40.77 50.10 55.62 49.83 43.58 28.09 34.30 32.78 23.60 25.17 15.77 72.00 90.17 42.46 30.75 3.50 96.50 21.50 55.85 55.65 42.40
Table 9: Full Qwen-2.5-7B/3B-Instruct LongBench Results with $B=128$ (Higher is better). We highlight the best and second best methods within a given budget with bold and underline.
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code Narrative QA Qasper MF-en MF-zh HotpotQA 2WikiMQA Musique DuReader GovReport QMSum MultiNews VCSum TREC TriviaQA SAMSum LSHT PCount PR-en PR-zh Lcc RB-P Avg. Qwen2.5-7B 15.75 16.94 32.38 14.87 11.89 11.88 7.95 30.56 34.33 19.91 22.67 15.28 65.50 87.05 44.75 39.47 4.22 93.08 68.79 57.74 61.84 36.04 H2O 2K 2.39 7.29 12.42 11.73 8.55 11.06 2.73 6.07 3.62 6.60 15.69 3.44 42.50 28.21 10.63 16.00 0.65 0.00 1.50 35.10 18.77 11.66 4K 1.99 11.92 19.88 14.72 10.24 10.12 4.73 7.51 9.08 10.14 20.85 6.15 51.00 37.37 20.57 15.75 3.16 6.43 27.62 52.14 29.09 17.64 6K 3.34 14.79 23.94 15.33 11.45 11.30 5.52 9.30 14.63 14.27 22.06 8.68 55.75 51.99 28.01 18.50 1.39 9.41 54.53 54.68 38.32 22.25 8K 6.10 15.55 28.29 14.99 12.37 14.65 6.24 16.10 20.78 17.22 22.44 11.12 59.00 58.74 33.05 24.92 1.82 15.73 55.16 55.63 44.56 25.45 TOVA 2K 8.49 14.01 21.04 11.55 14.00 11.51 5.09 14.45 27.43 17.84 22.83 15.75 56.50 79.56 40.55 20.50 2.43 9.29 20.45 55.99 56.15 25.02 4K 12.83 17.03 27.01 14.14 16.80 13.37 8.05 21.15 29.21 19.05 22.73 15.81 58.50 82.67 42.71 27.75 1.67 15.00 43.53 56.69 56.59 28.68 6K 15.77 15.33 30.31 14.58 19.30 13.78 9.11 25.70 30.40 19.95 22.91 15.10 61.50 83.47 42.90 27.60 1.15 21.75 55.16 57.68 57.99 30.54 8K 15.69 15.55 33.09 14.78 18.37 13.99 11.26 27.92 31.33 20.17 22.82 15.27 62.00 84.49 43.01 33.21 2.78 30.33 55.16 57.45 58.96 31.79 Sink 2K 7.68 14.68 19.36 12.98 8.58 9.34 3.97 10.66 27.75 17.96 22.33 14.26 62.00 75.26 42.76 23.00 1.07 7.50 21.70 50.11 49.57 23.93 4K 7.68 17.18 23.46 14.65 9.09 9.38 4.39 10.57 30.23 18.62 22.79 15.48 64.50 83.39 44.19 29.81 2.74 18.08 64.95 55.23 51.30 28.46 6K 7.37 16.61 25.73 14.74 11.29 11.27 5.69 11.49 31.47 18.72 22.86 15.62 64.50 84.86 44.47 31.07 3.59 41.48 71.21 55.89 55.99 30.76 8K 8.22 16.15 28.63 15.52 11.59 11.11 6.44 17.29 32.56 18.49 22.91 15.45 65.00 83.95 44.15 35.75 4.14 48.72 71.71 56.82 56.42 31.95 SnapKV 2K 11.60 12.45 23.66 12.54 12.38 10.64 7.03 14.40 27.57 18.27 22.85 15.23 58.00 81.78 41.13 23.67 3.76 19.42 35.09 55.83 56.53 26.85 4K 14.35 13.45 28.28 13.76 16.33 11.74 8.12 21.96 29.71 19.18 22.82 15.20 57.00 83.80 43.27 25.51 2.41 39.83 55.28 58.12 58.67 30.42 6K 14.34 16.35 31.12 14.16 17.56 14.10 8.74 25.56 31.09 20.16 22.84 15.04 60.00 83.80 42.99 30.81 2.91 54.17 55.16 57.48 60.26 32.32 8K 15.60 15.81 33.47 14.77 18.02 14.49 10.53 27.45 31.99 20.09 22.84 15.20 61.00 84.08 43.01 34.28 4.58 64.25 55.16 57.46 60.59 33.56 KeyDiff 2K 7.17 10.06 24.28 12.96 10.03 10.81 5.71 23.59 17.09 18.03 22.71 11.73 52.00 53.98 32.22 30.00 3.52 33.33 34.37 53.13 32.05 23.75 4K 13.16 12.00 32.08 14.15 13.04 13.68 5.39 27.83 25.61 20.42 22.76 13.37 54.00 70.90 40.23 39.62 3.37 58.42 54.86 55.95 42.27 30.15 6K 13.42 14.90 35.11 14.62 18.70 14.09 8.34 30.74 29.83 21.08 22.86 14.39 60.50 77.03 42.00 38.08 4.13 69.83 54.33 56.76 51.50 32.96 8K 14.90 15.77 34.32 14.90 19.02 13.93 9.27 30.65 31.29 20.90 22.79 14.69 60.00 83.01 43.65 40.70 3.87 74.13 55.16 57.33 52.80 33.96 Qwen2.5-3B 18.08 22.49 39.72 28.99 27.86 20.45 18.93 32.95 32.80 23.74 24.89 10.95 67.50 85.05 43.88 37.50 5.00 40.97 20.61 51.91 47.53 33.42 H2O 2K 1.80 9.18 11.62 12.62 8.54 7.31 2.70 8.57 5.93 6.97 16.89 4.15 38.00 21.87 7.69 16.00 1.00 3.00 2.69 37.36 22.90 11.75 4K 2.82 17.34 23.27 21.55 10.18 10.47 3.03 11.05 11.06 10.73 22.93 5.77 50.75 34.93 18.03 16.25 4.35 7.32 16.64 47.74 29.42 17.89 6K 5.52 18.62 27.93 27.26 12.61 15.07 4.26 14.65 14.92 13.89 24.21 7.55 58.00 45.94 24.93 16.00 2.91 9.10 21.32 49.50 34.54 21.37 8K 6.16 19.84 32.32 29.66 16.01 17.74 4.99 19.42 20.21 16.49 24.54 9.22 64.00 56.10 32.56 20.25 3.13 11.61 21.32 50.61 38.80 24.52 TOVA 2K 11.69 14.94 25.33 19.90 17.29 12.58 5.91 15.34 26.67 21.49 24.78 16.58 51.50 68.80 41.79 17.75 0.23 6.00 8.68 49.79 48.60 24.08 4K 12.19 18.31 32.56 27.33 20.58 13.80 7.74 21.11 28.82 22.27 24.98 15.82 59.00 80.66 43.05 21.25 1.11 9.56 19.18 49.93 46.74 27.43 6K 13.62 19.56 34.64 28.72 21.67 16.25 8.47 27.26 30.17 23.10 24.94 14.53 63.50 81.88 42.97 26.25 1.16 10.58 21.32 51.30 47.70 29.03 8K 14.66 20.93 37.77 29.72 22.57 17.08 9.63 29.10 31.12 23.17 24.83 13.48 67.00 84.11 43.55 28.25 2.06 13.08 21.32 51.32 47.64 30.11 Sink 2K 9.71 13.75 22.11 20.97 11.63 14.67 4.43 11.89 27.39 19.45 24.36 13.00 56.00 58.77 42.37 22.75 2.50 8.75 4.33 48.27 49.72 23.18 4K 11.46 18.28 30.40 24.02 15.50 14.62 6.97 11.48 30.08 20.12 24.86 13.35 63.00 68.77 43.11 29.50 3.00 11.75 16.75 51.76 50.47 26.63 6K 13.01 20.03 32.59 27.06 18.62 15.77 9.37 13.35 30.98 20.70 24.97 13.05 66.50 75.39 42.77 30.00 4.00 14.92 20.44 52.32 50.35 28.39 8K 10.26 21.27 35.15 29.49 24.31 17.60 9.40 17.59 31.81 21.14 24.99 12.07 68.50 79.17 43.32 34.50 1.00 24.00 20.44 51.47 49.38 29.85 SnapKV 2K 11.70 13.91 24.28 20.75 14.80 10.89 7.42 15.09 27.40 21.63 24.64 15.71 54.50 75.35 42.72 22.38 2.50 18.33 19.06 49.65 50.59 25.87 4K 12.98 22.21 31.77 26.57 18.33 14.41 10.83 21.14 29.14 22.38 24.89 15.88 61.00 84.17 42.63 21.17 3.75 25.42 22.46 50.22 48.77 29.05 6K 14.16 20.09 36.15 28.41 19.14 15.59 12.70 26.21 30.35 22.75 24.91 14.96 65.00 83.92 43.52 25.50 5.00 32.20 21.32 51.04 47.49 30.50 8K 12.76 20.88 37.10 30.10 22.49 18.19 13.83 29.54 31.33 23.37 24.80 13.75 65.50 84.88 44.49 28.00 5.20 35.83 21.32 51.31 47.82 31.55 KeyDiff 2K 3.99 10.20 22.71 15.77 8.93 13.12 5.51 24.79 17.35 16.56 24.31 10.53 57.50 41.19 27.43 25.25 3.88 11.32 10.68 46.44 34.33 20.56 4K 9.39 18.61 31.37 23.64 18.96 18.10 7.86 27.01 25.64 22.28 24.70 10.93 65.00 63.02 37.74 30.50 4.00 20.08 26.85 49.27 39.24 27.34 6K 10.51 19.71 35.51 28.89 26.92 18.28 11.47 31.83 29.32 23.63 24.90 11.46 64.50 75.36 41.51 34.50 3.57 31.41 21.32 50.60 42.95 30.39 8K 12.24 20.49 38.52 29.60 23.05 19.41 15.95 30.92 31.10 23.89 24.83 11.80 67.50 79.05 41.73 36.50 3.08 40.21 21.32 51.05 45.88 31.82
Table 10: Llama-3.2-3B-Instruct LongBench Results with prompt block size $B\in[64,256]$ (Higher is better). We highlight the best and second best methods within a given budget with bold and underline.
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code Narrative QA Qasper MF-en MF-zh HotpotQA 2WikiMQA Musique DuReader GovReport QMSum MultiNews VCSum TREC TriviaQA SAMSum LSHT PCount PR-en PR-zh Lcc RB-P Avg. $B=64$ 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 H2O 2K 1.30 18.23 16.96 14.25 12.26 16.84 0.72 6.88 0.78 1.29 16.24 2.97 35.00 18.07 9.78 13.50 0.50 1.50 0.25 39.98 12.68 11.43 4K 2.30 31.92 31.59 32.44 25.02 22.58 4.89 9.29 5.36 5.57 23.02 4.80 50.50 33.05 18.76 13.75 1.00 2.50 3.75 50.11 21.57 18.75 6K 3.12 38.87 37.63 44.58 34.38 35.35 12.13 12.82 10.43 9.38 24.31 6.63 63.00 45.46 26.08 14.25 0.00 8.50 17.50 53.84 30.66 25.19 8K 9.11 40.09 45.04 52.58 39.24 38.25 15.88 17.82 17.97 12.96 24.70 9.26 70.50 60.57 32.71 16.00 0.50 21.50 19.50 54.72 39.26 30.39 TOVA 2K 17.24 30.03 31.04 32.07 36.58 28.97 12.17 10.50 26.35 19.78 25.07 15.20 60.50 87.45 41.07 15.50 1.00 10.50 6.00 55.30 52.36 29.27 4K 19.59 39.27 42.16 44.54 44.58 37.63 18.62 17.44 28.82 21.46 25.18 16.49 62.50 89.48 41.89 17.50 3.50 24.00 15.00 55.14 56.58 34.35 6K 21.53 40.32 46.16 52.81 49.44 40.35 18.73 25.74 30.47 22.30 25.23 16.23 66.00 90.00 42.48 21.00 3.00 47.00 18.50 55.15 58.03 37.64 8K 21.32 40.87 50.20 54.84 49.35 42.11 24.52 30.71 31.60 23.05 25.20 16.57 69.00 90.50 41.80 23.25 5.50 74.50 19.50 55.45 58.34 40.39 Sink 2K 15.68 29.91 26.61 27.42 33.16 25.43 13.36 9.37 26.70 19.25 25.01 16.04 64.50 86.33 41.04 19.00 1.50 19.00 8.50 56.48 52.91 29.39 4K 19.35 37.77 36.91 41.61 41.46 35.26 12.88 10.30 29.59 20.27 25.01 16.09 69.00 88.06 41.88 21.25 2.50 36.00 16.50 55.85 52.51 33.81 6K 19.36 40.01 38.16 49.57 46.39 39.01 14.20 10.41 30.54 21.64 25.07 16.93 71.00 88.50 42.09 23.50 3.50 47.50 19.50 56.02 53.97 36.04 8K 20.49 39.89 42.21 53.08 47.72 41.44 16.37 15.08 31.80 22.04 25.07 16.89 72.00 89.26 42.20 25.25 3.50 61.50 20.50 56.21 56.42 38.04 SnapKV 2K 18.07 31.21 30.60 31.32 37.31 30.69 11.71 9.98 26.98 19.87 25.13 16.05 61.00 87.85 40.36 16.25 3.00 32.00 8.00 56.78 55.77 30.95 4K 19.30 39.01 40.81 44.86 47.83 37.74 16.75 16.94 29.90 21.37 25.25 16.70 65.00 88.88 40.99 16.75 3.50 59.50 18.00 55.15 57.20 36.26 6K 20.85 40.32 45.58 52.80 48.03 41.63 18.47 24.98 30.68 22.32 25.12 16.71 68.50 90.00 41.49 18.25 4.50 85.00 18.50 55.46 57.30 39.36 8K 20.64 41.10 47.89 54.70 48.49 41.79 21.58 31.46 32.03 23.32 25.22 16.65 71.00 90.00 41.44 21.00 4.00 95.00 19.50 55.44 57.54 40.94 KeyDiff 2K 17.40 38.12 45.25 47.09 45.28 34.23 13.97 27.96 28.34 21.09 24.94 13.45 56.00 83.29 38.53 24.25 1.00 63.50 12.50 54.41 40.00 34.79 4K 22.38 42.00 50.84 53.16 47.34 40.56 21.43 32.82 30.96 23.32 25.08 15.35 67.50 87.09 42.53 27.50 2.00 89.00 18.50 54.39 53.09 40.33 6K 22.25 41.55 50.32 54.91 51.61 42.02 24.62 34.60 32.40 23.63 25.26 16.18 71.00 88.42 41.90 29.25 3.50 95.00 19.50 55.69 55.27 41.85 8K 21.57 41.24 50.12 55.33 49.98 43.78 27.45 34.30 32.43 23.67 25.21 15.99 71.50 90.84 42.32 30.00 3.00 96.50 19.50 55.54 56.56 42.23 $B=256$ 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 H2O 2K 1.87 19.19 23.35 16.06 20.95 17.91 2.33 7.31 0.83 1.73 16.29 3.28 42.50 25.82 9.83 14.50 2.75 0.00 1.00 38.59 14.57 13.36 4K 5.58 31.49 33.28 32.36 26.85 28.34 8.61 10.83 5.46 6.27 23.13 4.98 54.00 41.73 19.14 15.00 1.00 8.00 6.50 51.39 24.24 20.87 6K 7.49 37.99 41.54 44.71 39.53 36.18 15.46 13.87 10.77 10.46 24.60 6.94 61.50 58.64 27.20 15.50 1.00 15.00 19.00 54.30 32.45 27.34 8K 9.92 39.71 43.84 52.15 39.14 39.23 19.28 18.04 17.78 13.60 24.98 8.99 71.00 67.64 32.83 17.25 2.50 30.00 21.00 55.25 39.33 31.59 TOVA 2K 18.46 30.80 33.74 32.24 39.73 32.18 14.10 10.17 26.32 20.17 25.18 15.67 62.00 89.36 40.60 16.25 2.00 16.00 4.50 55.98 53.42 30.42 4K 20.36 38.18 42.53 46.18 46.83 36.60 17.81 17.09 28.99 20.54 25.23 16.34 63.50 89.13 41.55 19.12 3.50 29.50 14.50 55.55 55.91 34.71 6K 20.71 40.46 45.82 52.95 51.33 40.92 21.47 24.80 30.71 22.37 25.48 16.40 67.00 88.50 41.91 20.25 3.50 49.00 20.00 55.85 56.74 37.91 8K 20.84 40.79 48.02 54.82 50.12 40.71 25.17 30.85 31.47 22.98 25.49 16.41 71.00 89.00 41.99 22.25 2.00 76.50 21.00 55.93 57.59 40.23 Sink 2K 15.48 30.74 26.93 27.76 33.86 25.63 13.30 9.56 26.70 19.41 25.15 15.56 66.00 86.44 41.17 18.50 1.00 19.00 8.50 56.47 52.37 29.50 4K 18.91 38.38 37.28 41.76 41.81 35.17 13.07 9.96 29.23 20.61 25.01 16.38 70.50 88.06 42.07 21.75 1.50 36.00 17.00 56.05 51.98 33.93 6K 19.37 40.30 38.14 49.77 46.14 39.60 14.18 11.12 30.48 21.54 25.20 16.28 71.00 88.80 41.79 23.50 3.50 47.00 20.00 55.65 53.53 36.04 8K 20.35 40.19 42.96 53.25 47.82 41.10 17.87 14.83 31.32 22.14 25.17 16.56 72.50 89.26 42.47 26.25 5.00 62.50 20.00 56.00 56.03 38.27 SnapKV 2K 17.04 31.55 31.75 31.87 37.25 34.03 12.17 9.80 27.17 20.16 25.26 16.27 61.00 87.63 40.95 17.50 4.00 35.00 10.00 55.93 54.39 31.46 4K 19.67 39.34 40.95 45.81 44.27 38.78 16.17 16.61 29.99 21.15 25.49 16.71 66.50 89.15 40.76 19.75 3.50 66.00 19.00 55.44 56.45 36.74 6K 22.98 40.13 44.80 52.39 50.61 39.28 20.31 24.84 31.28 22.17 25.42 16.55 70.00 89.79 41.91 19.75 3.50 86.50 20.00 55.93 57.49 39.79 8K 20.23 40.75 47.40 54.77 49.74 41.47 24.74 31.19 31.95 22.96 25.45 16.27 71.50 90.17 42.12 23.25 4.00 93.50 21.00 55.93 57.36 41.23 KeyDiff 2K 18.99 37.20 46.57 46.61 45.14 33.12 15.15 29.54 28.30 21.92 25.41 14.44 58.50 86.30 37.77 23.75 1.00 67.00 13.50 54.07 41.59 35.52 4K 21.00 41.48 48.69 52.94 47.07 40.26 22.69 33.07 31.11 23.35 25.22 15.57 65.50 88.72 41.79 26.50 2.00 91.00 19.00 55.79 51.50 40.20 6K 21.61 40.78 49.68 55.08 50.64 41.38 24.57 34.62 32.02 23.36 25.56 15.86 71.50 89.42 42.37 28.75 3.00 94.00 21.00 55.96 55.96 41.77 8K 22.24 40.83 50.23 55.45 50.50 43.28 28.37 33.88 32.67 23.85 25.48 15.83 72.50 90.34 42.31 30.50 3.00 95.50 21.00 56.01 55.58 42.35
[…]
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code Narrative QA Qasper MF-en MF-zh HotpotQA 2WikiMQA Musique DuReader GovReport QMSum MultiNews VCSum TREC TriviaQA SAMSum LSHT PCount PR-en PR-zh Lcc RB-P Avg. Llama3.1-8B 30.05 47.00 56.12 59.86 57.33 47.81 32.25 35.64 34.86 25.32 27.02 17.28 73.00 91.61 43.37 45.50 8.33 99.50 99.00 61.66 51.94 49.74 Sink 2K 21.83 34.27 29.24 32.82 38.64 29.50 12.59 16.18 28.51 20.21 26.62 15.54 65.00 89.46 42.20 22.25 2.00 25.50 32.50 64.95 59.54 33.78 4K 22.94 43.01 39.08 46.16 44.04 41.39 19.09 16.54 31.08 21.57 26.78 16.73 70.00 91.53 42.29 29.25 3.00 38.50 71.00 62.12 58.84 39.76 6K 25.41 47.40 44.13 52.78 47.39 45.73 21.90 17.55 32.53 22.19 26.87 17.05 72.00 91.25 43.41 33.75 3.08 52.50 98.00 62.22 56.24 43.49 8K 23.53 46.63 48.68 55.77 49.61 47.16 21.14 19.54 33.10 23.20 26.92 16.91 72.00 91.29 43.79 37.00 3.25 66.00 99.00 62.18 56.43 44.91 SnapKV 2K 21.81 37.22 37.19 38.29 46.10 35.42 16.53 16.37 29.83 21.05 26.77 16.16 61.00 88.84 42.56 21.75 4.03 51.50 81.17 62.37 51.45 38.45 4K 24.79 44.22 47.30 50.27 48.49 46.73 20.55 22.04 32.19 22.68 26.95 16.95 67.50 90.98 43.14 25.00 5.17 89.50 96.67 61.44 51.20 44.46 6K 24.10 45.57 50.44 55.27 53.12 48.41 24.27 27.46 33.43 23.53 27.03 16.84 71.50 92.28 43.58 27.00 5.25 98.00 99.00 61.32 52.16 46.65 8K 25.15 46.55 53.39 57.65 56.00 48.75 27.82 32.66 33.67 24.85 27.01 17.37 72.50 91.78 43.54 33.75 5.08 100.00 98.67 61.48 51.41 48.05 KeyDiff + 10% Recent Tokens 2K 27.36 40.79 50.63 50.21 49.44 44.77 25.04 30.25 29.41 23.28 26.76 15.78 64.00 85.90 43.76 44.00 4.25 98.50 95.83 62.97 48.70 45.79 4K 27.42 45.67 53.62 57.40 53.77 47.30 27.48 33.40 31.90 24.18 26.83 16.51 70.00 88.89 43.95 47.00 5.21 99.00 98.67 62.20 52.98 48.26 6K 30.58 46.24 55.58 57.82 56.44 47.74 29.03 34.33 33.15 24.67 26.91 16.98 71.50 92.22 43.96 47.00 4.23 99.50 99.00 62.56 52.98 49.16 8K 32.17 46.66 55.65 58.65 57.24 48.64 30.54 33.44 33.85 24.93 26.94 17.12 72.50 91.72 43.70 46.00 5.47 99.50 99.00 62.56 51.75 49.43 KeyDiff + 20% Recent Tokens 2K 26.73 41.88 49.79 48.46 49.68 42.51 26.90 30.90 28.83 23.34 26.76 16.11 62.50 87.39 43.99 42.50 4.58 94.00 95.58 63.33 50.01 45.51 4K 26.05 45.11 55.51 55.87 54.14 47.41 25.52 33.51 31.87 24.40 26.95 16.09 70.00 90.22 43.76 43.50 4.00 99.50 98.67 62.06 52.27 47.92 6K 28.39 46.43 55.12 58.27 57.02 48.94 28.98 34.31 33.32 24.62 26.96 17.20 71.50 91.72 44.02 46.50 5.13 99.50 99.00 62.39 52.57 49.14 8K 31.35 46.74 54.71 58.60 58.19 48.14 31.77 34.02 33.55 24.98 26.95 17.11 72.50 91.72 44.17 46.00 7.38 99.50 99.00 62.39 52.04 49.56 KeyDiff 2K 26.64 41.73 50.99 51.18 51.59 46.47 22.84 32.37 29.02 23.86 26.76 14.81 66.50 85.92 39.26 42.25 3.17 96.00 96.25 59.17 39.42 45.06 4K 28.70 45.62 56.06 56.83 54.58 49.31 28.25 33.06 32.30 25.03 27.07 16.32 70.00 90.85 42.84 44.50 4.21 99.00 97.67 60.80 48.00 48.14 6K 29.90 46.33 55.11 59.00 56.80 49.50 31.52 34.97 33.44 24.58 26.98 16.80 72.00 90.99 43.10 47.00 5.27 99.50 99.00 61.40 49.70 49.19 8K 33.57 46.77 55.48 59.16 56.87 49.37 30.88 34.54 34.17 25.12 27.01 17.13 72.50 92.28 42.81 46.50 5.83 99.50 98.67 61.48 50.90 49.55 Llama3.2-3B 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 Sink 2K 16.85 30.69 26.58 27.32 33.26 25.27 13.82 9.38 26.74 19.15 25.15 15.88 65.00 86.17 40.79 19.50 1.50 19.50 8.50 56.65 52.73 29.54 4K 19.46 38.61 36.22 41.68 41.97 35.84 13.37 9.86 29.34 20.19 25.06 16.44 71.00 88.06 41.31 21.75 2.50 35.50 16.00 56.48 52.43 33.96 6K 19.33 40.29 37.95 49.68 46.48 40.29 15.31 11.10 30.43 21.35 25.14 16.64 71.50 88.93 42.04 23.50 3.50 47.00 19.50 56.55 54.11 36.22 8K 20.15 40.02 41.94 53.57 48.15 42.24 16.01 14.76 31.64 22.10 25.20 16.50 73.00 89.26 42.37 26.25 3.50 62.50 20.50 56.86 56.63 38.25 SnapKV 2K 17.38 31.37 31.48 29.65 37.77 30.05 11.54 9.66 27.03 19.93 24.97 15.97 59.00 88.13 40.48 16.25 3.50 32.50 9.00 56.32 55.91 30.85 4K 19.85 39.22 39.86 47.33 46.70 37.98 16.64 16.88 29.79 21.21 25.01 16.74 65.50 89.35 40.95 18.25 2.50 62.50 22.50 55.74 56.88 36.73 6K 20.83 39.65 44.48 51.84 49.30 40.18 20.28 25.32 31.27 22.73 25.09 16.81 69.00 89.95 41.47 18.75 4.00 85.00 20.50 55.69 57.82 39.52 8K 20.49 40.80 48.16 55.44 48.78 41.65 24.79 30.40 31.81 23.46 25.17 16.44 70.00 90.17 41.99 22.00 5.00 94.00 21.50 55.77 57.29 41.20 KeyDiff + 10% Recent Tokens 2K 19.92 37.58 45.99 46.41 44.24 34.81 15.20 29.22 28.23 22.49 24.99 14.71 59.00 84.08 41.38 24.38 3.00 70.50 15.50 57.49 51.91 36.72 4K 21.97 40.70 49.76 52.53 47.79 41.92 20.71 32.83 31.12 23.37 25.07 15.77 67.50 87.47 41.28 25.00 2.50 91.00 20.00 57.14 56.25 40.56 6K 23.76 40.56 50.79 54.23 50.54 42.15 25.20 34.12 31.95 23.10 25.21 15.91 71.50 88.17 41.67 26.50 3.50 96.00 20.50 56.80 55.18 41.78 8K 23.65 40.58 49.96 55.71 51.52 44.10 25.95 34.41 32.80 23.77 25.25 15.75 73.50 89.67 41.83 28.75 3.00 96.00 20.50 57.00 56.12 42.37 KeyDiff + 20% Recent Tokens 2K 19.27 34.86 45.26 44.94 42.81 34.15 14.27 27.31 28.10 22.13 25.08 14.96 61.50 85.07 41.62 24.17 2.00 71.50 16.00 57.56 53.58 36.48 4K 22.56 41.28 48.37 52.00 47.50 42.64 19.64 32.28 31.17 22.83 25.08 15.68 68.50 88.17 41.38 24.25 2.00 90.50 19.50 57.04 55.05 40.35 6K 22.88 40.74 50.37 55.02 49.90 42.34 25.02 34.95 31.97 23.28 25.24 16.32 71.50 89.61 41.43 27.00 3.00 95.50 20.50 56.82 55.77 41.86 8K 23.82 40.58 50.02 55.71 51.25 43.64 24.52 34.33 33.02 23.82 25.20 16.08 73.50 89.67 41.99 29.25 3.00 97.00 20.50 57.05 56.48 42.40 KeyDiff 2K 18.29 36.65 45.44 47.47 46.09 35.41 13.79 28.89 28.16 21.45 25.01 13.56 60.00 85.24 37.00 24.88 1.00 60.50 12.00 54.13 42.01 35.09 4K 22.34 40.60 49.15 52.56 50.14 40.30 21.65 32.46 31.38 23.44 25.06 15.28 66.50 87.92 41.41 27.50 2.50 88.50 19.50 55.55 52.24 40.28 6K 22.29 40.68 50.14 54.51 51.74 42.19 24.83 34.64 32.39 23.53 25.19 15.88 71.00 90.02 42.00 28.75 3.00 95.00 21.50 55.86 54.39 41.88 8K 22.41 40.77 50.10 55.62 49.83 43.58 28.09 34.30 32.78 23.60 25.17 15.77 72.00 90.17 42.46 30.75 3.50 96.50 21.50 55.85 55.65 42.40
Appendix C Attention Sinks and Approximate Collinearity
(a) Keys
(b) Keys
(c) Queries
(d) Queries
Figure 11: Angular distributions of keys and queries measured with respect to to mean key and mean query
(a) Layer 25
(b) Layer 26
Figure 12: Cosine Similarity of mean key and mean query for each head in several layers.
To better understand why there exists a negative correlation between cosine similarity between keys and attention scores, we look to recent research that seeks to the importance of attention sinks in decoder-based LLMs. The authors in (Barbero et al., 2025) show that attention sinks emerge from training decoder-based LLMs since they can denoise the model and prevent rank collapse by limiting over mixing in attention heads. Moreover, attention patterns in decoder based models demonstrate that most attention logits are quite small (and almost always negative) for most keys and queries. This allows the attention sink to have high attention activation, preventing over mixing in addition to allowing the heads to specialize and selectively identify important tokens.
At the same time, the results in (Godey et al., 2024) suggest that hidden states, keys and queries are all approximately collinear in the sense that $\textrm{CosSim}(x_{i},x_{j})\gg 0$. In geometric terms, this means that most key tokens and query tokens lie in the same direction in Euclidean space. Our own results, seen in Figure 11 demonstrate that this is the case for both keys and queries. Our experiments show that most keys and queries lie within a small angular distance from the mean key and query. More than this, we see that across all heads, the mean key and mean query have negative cosine similarity Figure 12. Moreover, as seen in Figure 13, we find that the norms of keys and queries are tightly clustered around a relatively fixed value. This means that variations in the norm of key and query tokens have less impact on the magnitude of attention scores than their direction. These three observations indicate that most keys and queries combine to create uniformly small attention logits, and that larger attention weights are constructed by projecting keys closer to the direction of the mean query. This appears to be the fundamental mechanism through which over mixing is prevented: if most attention activations are very small, each head can increase the activations of a small number of keys across most queries selectively projecting them to be more aligned with the distribution of queries. This hypothesis is further supported by the fact that sink tokens themselves often have small norm, which results in an approximate no-op in the attention head as in (Barbero et al., 2025), however in this case, the only way for a key corresponding to a sink token to have high attention scores is if it is as parallel as possible to the set of query tokens.
(a) Keys
(b) Keys
(c) Queries
(d) Queries
Figure 13: Distribution of L2 norms for keys and queries in several layers.
Figure 14: Average cosine similarity per head between high scoring keydiff key tokens and mean query.
To verify the above hypothesis, we show Figure 14 that keys which have maximum angular difference from the mean key are aligned with the mean query, resulting in very large attention weights. This demonstrates how LLMs exploit the geometry of the hidden states and projections to limit over mixing, and selectively identify important tokens.
To summarize, we have for all attention heads in decoder based transformer models:
  • The majority of keys and queries are approximately collinear with their mean.
  • Mean keys and mean queries have negative cosine similarity across all heads.
  • Most keys and queries have a similar L2 norm.
  • Decoder based attention heads can selectively increase attention weights for a fixed key by aligning it with the mean query.
  • Key token importance can hence be measured by the angular distance between a key and the mean key.
We can show mathematically with some reasonable assumptions based on the above observations that key tokens with persistently high attention scores must be geometrically aligned with queries.
Theorem C.1.
Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$. Without loss of generality suppose $||q||=1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$:
$w=\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}.$
Then
$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1\leq\textrm{CosSim}({k^{*}},q)$
Proof.
To show this we have that
$w$ $=\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}$
$w\left(\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)\right)$ $=\exp({k^{*}}^{\top}q)$
$w\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)$ $=(1-w)\exp({k^{*}}^{\top}q)$
Note that $-M\leq k_{i}^{\top}q\leq M$ and hence
$w\sum\limits_{i=1}^{n}\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
$wn\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
$\frac{wn\exp(-M)}{1-w}$ $\leq\exp({k^{*}}^{\top}q)$
$\ \log(wn)-M-\log(1-w)$ $\leq M\textrm{CosSim}({k^{*}},q)$
$\frac{\log(wn)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
$\frac{\log(\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}n)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
$\frac{\log(\frac{\exp({k^{*}}^{\top}q)}{(n+1)\exp(M)}n)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
$\frac{-M\textrm{CosSim}(k^{*},q)+\log(\frac{n}{n+1})-2M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
$\frac{\log(\frac{n}{n+1})-2M-\log(1-w)}{M}$ $\leq 2\textrm{CosSim}({k^{*}},q)$
$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1$ $\leq\textrm{CosSim}({k^{*}},q)$
The above proof demonstrates that as long as the norms of the keys are bounded, in order for an attention head to be able to freely allocate $w$ attention weight to a given key ${k^{*}}$, the cosine similarity between ${k^{*}}$ and $q$ must be high, therefore ${k^{*}}$ and $q$ must be approximately collinear.
Using this result, we can also show that as long as the cosine similarity between ${k^{*}}$ and $q$ is high, while the cosine similarity between $\bar{k}$ and $q$ is low, $\textrm{CosSim}({k^{*}},\bar{k})$ is small. Note that, since empirical results demonstrate that most keys have high cosine similarity with their mean $\bar{k}$, a key with high importance, approximately collinear to $q$, will also have low cosine similarity to $\bar{k}$. Generally, this also suggests that key tokens with low cosine similarity to $\bar{k}$ have greater importance.
In order to show this, we need the following auxiliary result.
Lemma C.2.
Suppose $\{x_{1},...,x_{n}\}$ is an orthonormal basis of $\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$. Define $\alpha_{i}=\textrm{CosSim}(x_{i},y)$. Then $\sum\limits_{i=1}^{n}\alpha_{i}^{2}=1$.
Proof.
Note that $\alpha_{i}=\frac{y^{\top}x_{i}}{||y||||x_{i}||}=\frac{y^{\top}x_{i}}{||y||}$. If we expand $\frac{y}{||y||}$ in the basis $\{x_{1},...,x_{n}\}$ we see that
$\frac{y}{||y||}$ $=\sum\limits_{i=1}^{n}\left({\frac{y}{||y||}}^{\top}x_{i}\right)x_{i}$
$=\sum\limits_{i=1}^{n}\alpha_{i}x_{i}$
But then, we know that since $\left|\left|\frac{y}{||y||}\right|\right|_{2}^{2}=1$, then $\left\langle\sum\limits_{i=1}^{n}\alpha_{i}x_{i},\sum\limits_{i=1}^{n}\alpha_{i}x_{i}\right\rangle=1$. But we have
$\left\langle\sum\limits_{i=1}^{n}\alpha_{i}x_{i},\sum\limits_{i=1}^{n}\alpha_{i}x_{i}\right\rangle$ $=\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\langle\alpha_{i}x_{i},\alpha_{j}x_{j}\rangle$
$=\sum\limits_{i=1}^{n}\langle\alpha_{i}x_{i},\alpha_{i}x_{i}\rangle$
$=\sum\limits_{i=1}^{n}\alpha_{i}^{2}$
proving the result. ∎
Theorem C.3.
Consider tokens ${k^{*}}$, $q$, $\bar{k}$ as above where $\bar{k}$ is the average of the keys tokens. Suppose $\textrm{CosSim}({k^{*}},q)=\beta_{q}>0$ and $\textrm{CosSim}(\bar{k},q)=\alpha_{q}<0$. Then $\textrm{CosSim}(\bar{k},{k^{*}})\leq 1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$.
Proof.
Consider the cosine similarity of $\bar{k}$ and ${k^{*}}$:
$\textrm{CosSim}(\bar{k},{k^{*}})$ $=\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$
expand $\bar{k}$ in an orthonormal basis which contains $q$, $\{q,r_{1},...,r_{n-1}\}$ such that
$\bar{k}=||\bar{k}||\left(\alpha_{q}q+\sum\limits_{i=1}^{n-1}\alpha_{i}r_{i}\right)$
where $\alpha_{i}=\textrm{CosSim}(\bar{k},r_{i})$. Additionally, define $\beta_{i}=\textrm{CosSim}({k^{*}},r_{i})$ and note that by the definition of an orthonormal basis and the cosine similarity operation, using the result from Lemma C.2 we have that $\alpha_{q}^{2}+\sum\limits_{i=1}^{n-1}\alpha_{i}^{2}=1$ and that $\beta_{q}^{2}+\sum\limits_{i=1}^{n-1}\beta_{i}^{2}=1$. Now we have that
$\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$ $=\frac{{k^{*}}^{\top}\left(||\bar{k}||\alpha_{q}q+\sum\limits_{i=1}^{n-1}||\bar{k}||\alpha_{i}r_{i}\right)}{||{k^{*}}||||\bar{k}||}$
$=\alpha_{q}\beta_{q}+\frac{1}{||{k^{*}}||}\sum\limits_{i=1}^{n-1}\alpha_{i}{k^{*}}^{\top}r_{i}$
$=\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}\alpha_{i}\beta_{i}$
$\leq\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}|\alpha_{i}||\beta_{i}|$
Applying Young’s inequality we obtain
$\leq\alpha_{q}\beta_{q}+\frac{1}{2}\sum\limits_{i=1}^{n-1}\alpha_{q}^{2}+\beta_{q}^{2}$
$=\alpha_{q}\beta_{q}+\frac{1}{2}(1-\alpha_{q}^{2})+\frac{1}{2}(1-\beta_{q}^{2})$
$=1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$
Note that on the domain $\alpha_{q}\in[-1,0)$, $\beta_{q}\in(0,1]$ the function $1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$ is bounded above by $1$ and decreasing to $-\frac{1}{2}$ as $\alpha\rightarrow-1$. Hence, the smaller $\textrm{CosSim}(\bar{k},q)=\alpha_{q}$ is, the smaller $\textrm{CosSim}(\bar{k},{k^{*}})$ must be.

### v3

@@
 Abstract
-In this work, we demonstrate that distinctive keys during LLM inference tend to have high attention scores.
-We explore this phenomenon and propose KeyDiff, a training-free KV cache eviction method based on key similarity.
-This method facilitates the deployment of LLM-based application requiring long input prompts in resource-constrained environments with limited memory and compute budgets.
+We demonstrate that geometrically distinctive keys during LLM inference tend to have high attention scores.
+Based on the phenomenon we propose KeyDiff, a training-free KV cache eviction method based solely on key similarity.
 Unlike other KV cache eviction methods, KeyDiff can process arbitrarily long prompts within strict resource constraints and efficiently generate responses.
-We demonstrate that KeyDiff computes the optimal solution to a KV cache selection problem that maximizes key diversity, providing a theoretical understanding of KeyDiff.
-Notably, KeyDiff does not rely on attention scores, allowing the use of optimized attention mechanisms like FlashAttention.
-We demonstrate the effectiveness of KeyDiff across diverse tasks and models, illustrating a performance gap of less than 0.04% with 8K cache budget ($\sim$ 23% KV cache reduction) from the non-evicting baseline on the LongBench benchmark for Llama 3.1-8B and Llama 3.2-3B.
-Machine Learning, ICML
+We provide a theoretical basis for KeyDiff by relating key diversity with attention scores.
+These results imply KeyDiff can efficiently identify the most important tokens to retain.
+Notably KeyDiff does not rely on attention scores, allowing the use of optimized attention mechanisms like FlashAttention.
+Under a strict memory allowance, we demonstrate the effectiveness of KeyDiff for the Llama and Qwen model families by observing a performance gap of less than 0.04% with 8K cache budget ($\sim$ 23% KV cache reduction) from the non-evicting baseline on LongBench for Llama 3.1-8B and Llama 3.2-3B.
+We also observe near baseline performance for Deepseek-R1-Distill-Llama-8B on the Math500 reasoning benchmark and decrease end-to-end inference latency by up to 30% compared to the other token-eviction methods.
 1 Introduction
-Large language models (LLMs) have revolutionized natural language processing, enabling tasks like document summarization, code generation, question answering (Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023; Dubey et al., 2024).
-Many exciting applications of LLMs, such as retrieval augmented generation (Lewis et al., 2020) and reasoning tasks (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2024), require very long input prompts, or long context support, to provide a valuable end-user experience.
-However, deploying these models in production environments has grown in importance and managing long input prompts remains a challenge due to large compute and memory requirements.
-Due to cost and privacy concerns, there has been a recent focus on deploying LLMs on the edge, i.e.
-directly on devices such as mobile phones or personal computers (Alizadeh et al., 2023; Liu et al., 2024c; van Baalen et al., 2024; Xu et al., 2024a).
-This imposes tight resource constraints, such as limited or shared compute resources, limited available DRAM, and even power limits in the mobile setting.
-Many techniques, such as system optimization (Alizadeh et al., 2023; Liu et al., 2024c; Xu et al., 2024a), model quantization (van Baalen et al., 2024), and efficient text generation (Goel et al., 2024) help to enable on-device LLM inference, but the main challenge for long context application on-device deployment is effectively managing key and value states.
-In a naïve implementation of autoregressive text generation, key and value states (KVs) must be recomputed for each new generated token.
-In a decoder-only transformer model, one can instead store the previously computed KVs in a KV cache for later use, allowing for a significant decrease in LLM inference latency.
-This comes at the cost of maintaining the KV cache, whose memory footprint grows linearly with the input sequence length (Yang et al., 2024a) and will violate any imposed memory limit with a sufficiently large input.
-One solution to bound this memory overhead is to fix the number of KVs in the cache, or cache budget, and use a cache eviction policy to remove KVs from the cache when the budget has been exceeded, bounding the KV cache memory usage.
-However, state-of-the-art cache eviction policies retain KVs with high attention weights (Xiao et al., 2024; Oren et al., 2024; Zhang et al., 2024b), which requires explicitly materializing the attention matrix of the entire input prompt.
-This results in large intermediate memory overhead during cache eviction, often violating memory limits.
-This intermediate cache memory overhead can be mitigated by segmenting the input prompt into blocks, sequentially processing each block by the LLM and caching the intermediate KVs (Kwon et al., 2023; Agrawal et al., 2023; Xu et al., 2024b).
-The cache eviction policy can be applied block-wise to satisfy imposed memory constraints (Figure 1).
-However, our experiments show that applying existing attention-based eviction methods with block prompt processing results in performance degradation of downstream tasks (Table 1).
-This is due to the limited scope of the attention computation during block prompt processing: attention scores are only computed for a subsequence of the original prompt and only contain local information.
-These partial scores may not be representative of a token’s global importance across the entire prompt and can result in incorrect cache eviction.
-This means that we need an eviction method that is independent from attention and cheap to compute during inference while maintaining model capability.
-The “attention sink” phenomenon (Xiao et al., 2024; Sun et al., 2024) describes the observation that, in decoder-only transformer-based LLMs, high attention scores are assigned to the first few tokens in the KV cache regardless of the input.
-This implies that eviction can be determined by only inspecting KVs and that an attention-free cache eviction criteria exists.
-In fact, we observe that keys with lower average pairwise cosine similarity also have higher attention outputs, as shown in Figure 2.
-Motivated by these observations, we propose an eviction method called KeyDiff, that evicts keys from the cache with the highest cosine similarity to a cache representative called an anchor.
-We also show that KeyDiff computes the optimal solution of the KV cache selection problem that minimizes pairwise key cosine similarity.
-Our experiments demonstrate that KeyDiff can achieve an accuracy drop of $\leq$ .04% on LongBench (Bai et al., 2024) with 8K token budget ($\sim$ 23% KV cache reduction), and can match or outperform state-of-the-art eviction methods, especially when using block prompt processing.
-Our contributions are as follows:
-  • We present the empirical observation that low pairwise cosine similarity among keys is correlated with higher attention output.
-  • We design an eviction policy based on this observation, called KeyDiff, that enables long context LLM inference with limited memory overhead.
-  • We show that KeyDiff computes the optimal solution to a KV cache selection problem that maximizes key diversity measured by pairwise cosine similarity.
-  • We demonstrate the effectiveness of KeyDiff on the LongBench task suite, achieving only 1.5% and 0.04% performance gap from the eviction-free baseline, using 6K and 8K cache budgets, respectively.
+Key-Value (KV) caching is a standard technique to accelerate large language model (LLM) inference that reuses key and value states (KVs) from previously processed tokens, enabling efficient autoregressive generation.
+This is crucial for long-context applications such as document summarization, code generation, question answering [7, 25, 31, 10], retrieval augmented generation [19] and reasoning [34, 17, 41].
+However, the memory footprint of the stored KV cache grows linearly with input length, which becomes a bottleneck in memory-constrained environments.
+This challenge is particularly acute for LLM inference on edge device, where compute, memory, and power resources are limited [3, 22, 32, 36].
+While KV cache eviction policies have been proposed to bound memory overhead by removing unimportant KVs (often measured by attention scores) [35, 24, 43], they typically process the entire prompt at once and violate memory constraints during intermediate computation.
+To enforce strict memory bounds throughout the prompt prefill and token generation inference phases, we adopt a block-wise inference strategy: the input prompt is divided into smaller blocks which are processed sequentially by the model, similar to [18, 1, 37].
+After processing each block, we evict some of the cached KVs by according to an eviction policy that scores each KV, as illustrated in Figure 1.
+Unlike previous approaches that apply eviction after processing the entire prompt, this strategy satisfies memory constraints throughout the full inference process.
+However, we observe a degradation in accuracy when applying existing eviction methods in this setting (Table 1).
+We hypothesize that the performance drop stems from a mismatch in design: existing eviction methods assume access to full-prompt attention, where key importance is computed over the entire input.
+During block prompt processing, however, attention is computed using only the current block’s tokens without access to future blocks.
+As a result, attention scores based on a limited context often fail to reflect a token’s true importance across the full prompt.
+To this end, we observe that keys with lower average pairwise cosine similarity tend to receive higher attention scores across a variety of inputs.
+This suggests that key diversity serves as a strong proxy for global token importance, even without access to future tokens.
+This insight enables an attention-free approach to cache eviction that is based on the geometry of the cached keys.
+Motivated by these observations, we propose KeyDiff, an attention-free cache eviction method that removes redundancy among cached keys, operates effectively during block-wise inference, and avoids excessive memory overhead.
+Our contributions are summarized as follows:
+  • Insight. We observe that lower pairwise cosine similarity among keys correlates with higher attention scores, suggesting its utility as a proxy for token importance. (Section 3.1)
+  • Method. We introduce KeyDiff, an eviction strategy that selects keys based on their similarity to other cached entries without relying on attention scores or future tokens. (Section 3.2)
+  • Theory.
+Through our analysis of key and query geometry we provide a theoretical understanding how/why KeyDiff works.
+We also show that KeyDiff solves an optimal subset selection problem that maximizes key diversity.
+(Section 3.3)
+  • Performance.
+KeyDiff achieves $\leq 1.5\%$ and $\leq 0.04\%$ accuracy drop on LongBench with 6K and 8K cache budgets, respectively, outperforming state-of-the-art eviction methods across Llama and Qwen models (Section 4.2), and near non-evicting baseline performance for Deepseek-R1-Distill-Llama-8B on the Math-500 reasoning benchmark.
+(Section 4.3)
+  • Efficiency. We observe up to $30\%$ end-to-end inference latency reduction using KeyDiff compared to existing KV cache eviction methods. (Section 4.5)
 2 Background
@@
 2.1 Transformers
-The Transformer architecture (Vaswani, 2017) processes input data using a sequence of transformer blocks.
-A transformer block $f$ takes a sequence $X=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T\times d}$ as input and applies the causal self-attention operator $\operatorname{Attention}$ followed by a feed-forward network $\operatorname{FF}$ with optional gating (Shazeer, 2020) to produce the output $X^{\prime}=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{T})\in\mathbb{R}^{T\times d}$:
+The Transformer architecture [33] processes input data using a sequence of transformer blocks.
+A transformer block $f$ takes a sequence $X=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T\times d}$ as input and applies the causal self-attention operator $\operatorname{Attention}$ followed by a feed-forward network $\operatorname{FF}$ with optional gating [28] to produce the output $X^{\prime}=(x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{T})\in\mathbb{R}^{T\times d}$:
 $X^{\prime}=f(X)=\operatorname{FF}(\operatorname{Attention}(X)),$ (1)
 The causal $\operatorname{Attention}$ operator projects each input token $x_{t}$ with matrices $W_{q},W_{k},W_{v}\in\mathbb{R}^{d\times d}$ into key, query, and value matrices ($K=XW_{k}$, $Q=XW_{Q}$, $V=XW_{V}$, respectively) then applies the following relation to produce the attention output  [^1] The multi-head extension and output projections are omitted for brevity.:
-$O^{\operatorname{attn}}=\operatorname{Softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+M\right)V=AV$ (2)
+$O^{\operatorname{attn}}=\operatorname{Softmax}\left(\Large\nicefrac{{QK^{\top}}}{{\sqrt{d}}}+M\right)V=AV$ (2)
 where $O^{\operatorname{attn}}\in\mathbb{R}^{T\times d}$, and the causal attention mask $M$ is an upper triangular matrix with nonzero values of $\infty$- ∞.
@@
 We can apply Equation 2 to an existing KV cache $\mathcal{C}$ as follows:
-$o^{\operatorname{attn}}_{T+1}=\operatorname{Softmax}\left(\frac{q_{T+1}[K\|k_{T+1}]^{\top}}{\sqrt{d}}+M\right)[V\|v_{T+1}],$ (3)
+$o^{\operatorname{attn}}_{T+1}=\operatorname{Softmax}\left(\Large\nicefrac{{q_{T+1}[K\|k_{T+1}]^{\top}}}{{\sqrt{d}}}+M\right)[V\|v_{T+1}],$ (3)
 where $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ are the key, query, and value states of $x_{T+1}$, and $[X\|x_{T+1}]$ represents the concatenation of $x_{T+1}$ to an existing tensor $X$ along the time dimension, $M$ is the causal attention mask accounting for both the KV cache and $x_{T+1}$.
 KV caching dramatically reduces the latency of $\operatorname{Attention}$ by only computing $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ for each token $x_{T+1}$ and reusing the KVs in $\mathcal{C}$.
-However, the size of the KV cache increases linearly with the number of processed tokens and dominates the memory footprint in long-context applications (Yang et al., 2024a).
+However, the size of the KV cache increases linearly with the number of processed tokens and dominates the memory footprint in long-context applications [39].
 2.3 KV Cache Eviction Methods
@@
 $S$ $=\texttt{topk}(\phi(A),N)$ (5)
-$K^{\prime}$ $=\texttt{gather}(K,S)$
-$V^{\prime}$ $=\texttt{gather}(V,S)$
+$K^{\prime}$ $=\texttt{gather}(K,S),\quad V^{\prime}=\texttt{gather}(V,S)$
 where $\texttt{topk}(x,N)$ returns the indices of $N$-largest values of $x$ and $\texttt{gather}(X,S)$ gathers columns indexed by $S$.
@@
 Existing eviction policies like Zhang et al.
-(2024b); Oren et al.
-(2024) focus on processing the entire input prompt at once: KVs are computed for each token in the prompt and stored in a cache $\mathcal{C}$, then the eviction policy $\pi_{N}$ is applied to reduce the number of tokens in $\mathcal{C}^{\prime}$ to $N$, before token generation.
+[43], Oren et al.
+[24] focus on processing the entire input prompt at once: KVs are computed for each token in the prompt and stored in a cache $\mathcal{C}$, then the eviction policy $\pi_{N}$ is applied to reduce the number of tokens in $\mathcal{C}^{\prime}$ to $N$, before token generation.
 However, the intermediate cache $\mathcal{C}$ before eviction will grow to the size of the input prompt.
 This can often exceed model’s allocated memory limit when deploying long context applications in resource constrained environments.
-As demonstrated in efficient LLM inference frameworks (Agrawal et al., 2023; Holmes et al., 2024; Kwon et al., 2023; Xu et al., 2024a), one solution is to apply $\pi_{N}$ more frequently by segmenting $X$ into non-overlapping blocks $X=[X_{0},X_{1},\ldots,X_{m-1}]$, where $X_{i}=[x_{Bi},\ldots,x_{B(i+1)-1}]$, $B$ is the block size, and $m=\lceil T/B\rceil$, and iteratively updating the cache by exploiting causality, applying Equation 4 in a block-wise fashion:
+As demonstrated in efficient LLM inference frameworks [1, 14, 18, 36], one solution is to apply $\pi_{N}$ more frequently by segmenting $X$ into non-overlapping blocks $X=[X_{0},X_{1},\ldots,X_{m-1}]$, where $X_{i}=[x_{Bi},\ldots,x_{B(i+1)-1}]$, $B$ is the block size, and $m=\lceil T/B\rceil$, and iteratively updating the cache by exploiting causality, applying Equation 4 in a block-wise fashion:
 $\mathcal{C}_{i}$ $\leftarrow([K_{i-1}\|k_{Bi:B(i+1)-1}],[V_{i-1}\|v_{Bi:B(i+1)-1}])$ (6)
@@
 where $k_{Bi:B(i+1)-1}$ and $v_{Bi:B(i+1)-1}$ are the keys and values selected from $X_{i}$ respectively, and $\mathcal{C}_{i}=(K_{i},V_{i})$ is the KV cache after processing the first $i$ prompt blocks.
-As in Equation 4, we concatenate the $B$ new KVs to the current cache, apply the eviction policy $\pi_{N}$ and update the cache in Equation 6.
+As in Equation 4, we concatenate the $B$ new KVs to the current cache, apply $\pi_{N}$ and update the cache in Equation 6.
 We refer to this as block prompt processing.
@@
 3 Method
-We demonstrate a negative correlation between attention scores and the similarity among keys (Section 3.1) and leverage this observation to develop KeyDiff (Section 3.2), followed by a theoretical justification of the method and preliminary evidence of its efficacy (Section 3.3).
+We demonstrate a negative correlation between attention scores and the cosine similarity among keys (Section 3.1) and leverage this observation to develop KeyDiff (Section 3.2), followed by a theoretical justification of KeyDiff (Section 3.3) and preliminary evidence of its efficacy (Section C.2).
 3.1 Correlation of Attention Scores and Key Dissimilarity
 To address the shortcomings of attention-based KV cache eviction in Section 2.4, we develop an alternative attention-free scoring metric that retains significant KVs across blocks while being resource efficient.
-We recall the “attention sink” phenomenon: LLMs often assign high attention weight to the first few tokens, regardless of the input (Xiao et al., 2024; Sun et al., 2024); these highly weighted tokens are called sink tokens.
+We recall the “attention sink" phenomenon: LLMs often assign high attention weight to the first few tokens, regardless of the input [35, 29]; these highly weighted tokens are called sink tokens.
 However, the index of the sink tokens can vary across heads and layers and reside deeper in the sequence than the first few tokens.
@@
 Pairwise cosine similarity of keys is solely a function of the keys in the cache, which are independent of input queries; the surprising aspect of Figure 2 is the negative correlation with attention weights.
-These distinctive keys essentially recover the attention sink phenomenon (Xiao et al., 2024).
+These distinctive keys essentially recover the attention sink phenomenon [35].
 3.2 KeyDiff
@@
 If the cache $\mathcal{C}$ has intermediate size $n$ and budget $N$ where $n>N$, $\pi^{\textsc{KeyDiff}{}}_{N}$ is defined as:
-$S$ $=\texttt{topk}(-\operatorname{CosSim}(K)\mathbf{1},N)$ (7)
-$K^{\prime}$ $=\texttt{gather}(K,S)$
-$V^{\prime}$ $=\texttt{gather}(V,S)$
-where $K\in\mathbb{R}^{n\times d}$ and $V\in\mathbb{R}^{n\times d}$ are the cached keys and values, $\operatorname{CosSim}(K)\in\mathbb{R}^{n\times n}$ is the pairwise cosine similarity matrix of keys in $K$ with $\operatorname{CosSim}(K)_{ij}=\frac{\langle k_{i},k_{j}\rangle}{\|k_{i}\|\|k_{j}\|}$, and $\mathbf{1}\in\mathbb{R}^{n}$ is a vector of ones.
+$S$ $=\texttt{topk}(-\operatorname{CosSim}(K)\mathbf{1},N),$ (7)
+$K^{\prime}$ $=\texttt{gather}(K,S),\quad V^{\prime}=\texttt{gather}(V,S)$
+where $K\in\mathbb{R}^{n\times d}$ and $V\in\mathbb{R}^{n\times d}$ are the cached keys and values, $\operatorname{CosSim}(K)\in\mathbb{R}^{n\times n}$ is the pairwise cosine similarity matrix of keys in $K$ with $\operatorname{CosSim}(K)_{ij}=\frac{k_{i}\cdot k_{j}}{\|k_{i}\|\|k_{j}\|}$, and $\mathbf{1}\in\mathbb{R}^{n}$ is a vector of ones.
 Figure 3: An overview of KeyDiff.
@@
 Efficient Variant of KeyDiff
-Unlike attention-based eviction policies, KeyDiff does not require access to the attention weights $A$, facilitating optimized attention kernels that do not materialize $A$, such as FlashAttention (Dao et al., 2022).
+Unlike attention-based eviction policies, KeyDiff does not require access to the attention weights $A$, facilitating optimized attention kernels that do not materialize $A$ such as FlashAttention [8].
 However, computing the pairwise cosine similarities runs in $\mathcal{O}(n^{2})$ time.
-Fortunately, the operation
-$\texttt{topk}(-\operatorname{CosSim}(K)\mathbf{1},N)$
-can be computed in $\mathcal{O}(n)$ operations on average by computing the score of each token as
-$s_{i}=-\operatorname{CosSim}(\mu(\hat{K}),\hat{k}_{i})$
-where $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\hat{k}_{i}$ and $\hat{k}_{i}=\frac{k_{i}}{||k_{i}||}$, and retaining the tokens with the $N$ largest scores with a variation of the quick select algorithm.
-This simplification is derived in Section 3.3.
-The eviction score of the efficient expression of KeyDiff is defined as follows:
-$s$ $=-\operatorname{CosSim}(\mu(\hat{K}),K)$ (8)
-$S$ $=\texttt{topk}(s,N)$
+Fortunately, we can compute the score of each token in Equation 7 in $\mathcal{O}(n)$ as follows:
+$S$ $=\texttt{topk}(-\operatorname{CosSim}(\mu(\hat{K}),\hat{k}_{i}),N)$ (8)
+where $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\hat{k}_{i}$ and $\hat{k}_{i}=\frac{k_{i}}{||k_{i}||}$.
 We refer to $\mu(\hat{K})$ as the anchor vector.
-Experimentation has shown that the anchor vector $\mu(\hat{K})$ can be replaced with $\mu(K)$ without losing accuracy (see Table 2).
-We evaluate this efficient version of KeyDiff described in Figure 3 in the following sections.
-3.3 Derivation of KeyDiff from an Optimization Perspective
-[…]
-This is a combinatorial optimization problem, which is difficult to solve efficiently, particularly during inference.
-However, we can relax Section 3.3 to produce a more tractable approximation to the original problem.
-First, we rewrite Section 3.3 by normalizing keys $k_{i}$ such that $\hat{k}_{i}=\frac{k_{i}}{\|k_{i}\|}$ resulting in:
-$\sum\limits_{i\in S}\sum\limits_{j\in S}\hat{k}_{i}\cdot\hat{k}_{j}$ $=\sum\limits_{i\in S}\hat{k}_{i}\cdot\left(\sum\limits_{j\in S}\hat{k}_{j}\right)=\sum\limits_{i\in S}\langle\hat{k}_{i},N\mu(\hat{K}_{S})\rangle$
-[…]
-Key Cache Diversity and KeyDiff
-To empirically verify KeyDiff’s ability to retain diverse keys, we apply PCA to the keys computed in an attention block of Llama 3.2-3B-Instruct after evaluating a long context prompt.
-We visualize the distribution of retained and evicted keys by applying sink attention (Xiao et al., 2024), TOVA (Oren et al., 2024), and KeyDiff as eviction policies in Figure 4.
-Visual observation reveals the tokens retained by sink attention and TOVA tend to tightly cluster together while KeyDiff’s retained tokens are more evenly distributed.
-This observation generalizes across heads and layers, as shown in Section A.3.
-We also visualize the log determinant of the Gram matrix of the key cache $\log(\det(KK^{T}))$ generated using different eviction policies in Figure 5.
-This quantity corresponds to the volume of space spanned by the keys in $\mathcal{C}$.
-The distribution of volumes for KeyDiff attains higher values, indicating that the retained keys are more distinctive than TOVA and sink, corroborating the results of Figure 4.
-Details of how these plots were generated are discussed in Section B.1.
+We show this formulation retains the same KVs of Equation 7 under a mild condition.
+(see Section C.2).
+Our experimentation has shown that the anchor vector $\mu(\hat{K})$ can be replaced with $\mu(K)$ without losing accuracy (see Table 14).
+We evaluate the efficient KeyDiff described in Figure 3 using unnormalized keys $k$ in all subsequent sections.
+Figures 5, 8, 9 and 10 visualize the keys retained and evicted by sink attention [35], TOVA [24] and KeyDiff via PCA.
+KeyDiff retains more varied keys.
+A full complexity and FLOP analysis can be found in Appendices B and B.1
+KeyDiff with Sliding Window
+In tasks such as reasoning and coding, where the most recent tokens are often important, we can augment KeyDiff and its efficient variant to use a percentage of the cache budget for a sliding window [6], which we call KeyDiff with sliding window.
+This extension introduces no complexity or memory overhead and we observe better results on certain tasks than vanilla KeyDiff (Tables 13 and E).
+3.3 Why KeyDiff Works: A Theoretical Perspective
+To solidify the theoretical foundation of KeyDiff and show that KeyDiff ultimately selects keys most aligned with queries, we prove the following two results.
+We first validate the relationship between cosine similarity and attention scores observed in Figure 2 by bounding the attention score of a new incoming key $k^{\star}$ in terms cosine similarity with a fixed query $q$:
+Lemma 3.1.
+Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$.
+Without loss of generality suppose $||q||=1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$.
+Then, for $n\to\infty$,
+$\frac{-\log(1-w)}{2M}-1\leq\textrm{CosSim}({k^{*}},q)$
+We then establish a relationship between the cosine similarities of $k^{\star}$, $q$, and the mean of prior keys $\bar{k}$:
+Theorem 3.2.
+Consider tokens ${k^{*}}$, $q$ as above, and the average of the keys tokens $\bar{k}$. Suppose $\textrm{CosSim}({k^{*}},q)=\beta_{q}>0$ and $\textrm{CosSim}(\bar{k},q)=\alpha_{q}<0$. Then
+$\textrm{CosSim}(\bar{k},{k^{*}})\leq 1+\alpha_{q}\beta_{q}-0.5\alpha_{q}^{2}-0.5\beta_{q}^{2}.$ (9)
+Figure 4: PCA embedding of keys and queries from Llama 3.2 3B
+By combining Lemmas 3.1 and 3.2, we establish a relationship between the attention weight $w$ and the KeyDiff score $\mathrm{CosSim}(\bar{k},k^{*})$.
+As $\mathrm{CosSim}(\bar{k},q)$ decreases and $\mathrm{CosSim}(k^{*},q)$ increases (along with the attention weight $w$), then $\mathrm{CosSim}(\bar{k},k^{*})$ tends to $-1$: this means KeyDiff selects distinct keys most aligned with $q$.
+We visualize this in Figure 4 with a PCA embedding of keys and queries from a single head of Llama 3.2 3B, highlighting the relationship between top scoring keys via KeyDiff, the anchor vector and queries.
+Similar trends are found from the other layers and heads as shown in Figure 11.
+The proofs of Lemmas 3.1 and 3.2 are in Section C.3, along with empirical motivation for the chosen assumptions.
 (a) Sink Attention
@@
 Retained tokens are blue, while evicted tokens are orange.
-The retained tokens using Sink and TOVA cluster together, while KeyDiff retains more diverse keys.
 Keys are taken from layer $5$ and head $3$ of Llama3.2-3B-Instruct, and generated using the NarrativeQA dataset.
 (d) PCA visualization of the retained keys for each KV cache eviction method.
-We present distributions of additional heads and layers in Section A.3.
-Figure 5: Distribution of $\log\left(\textrm{det}\left(KK^{T}\right)\right)$ from the Qasper dataset in LongBench.
-Larger values mean more of the key space is spanned by the key cache.
-KeyDiff retains keys that span a greater volume of the ambient space than TOVA or sink attention.
-4 Experiment
+4 Experiments
 In this section, we empirically demonstrate the effectiveness of KeyDiff.
-We begin with a description of competing, state-of-the-art eviction methods, followed by a detailed description of the evaluation setup.
-We then test each eviction policy on the LongBench task suite (Section 4.2) and report results in Table 1 followed by an ablation study (Section 4.3) of important properties of KeyDiff.
+We begin with a description of competing, state-of-the-art eviction methods, followed by a detailed description of the evaluation setup, then present our experimental results.
 Our findings can be summarized as follows:
-  • KeyDiff outperforms competing eviction policies on the Needle-In-A-Haystack benchmark (Kamradt, 2023) (Section 4.1).
+  • Needle-In-a-Haystack. KeyDiff outperforms competing eviction policies on the Needle-In-A-Haystack benchmark (Section 4.1).
+  • LongBench.
 KeyDiff outperforms competing eviction policies with block size $B=128$ on LongBench, achieving an 1.5% accuracy drop with a 6K cache budget ($\sim$33compression rate) and $\leq$ .04% with a 8k cache budget ($\sim$23compression rate) with Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Section 4.2).
-  • KeyDiff, utilizing negative cosine similarity as the eviction criteria and the mean of cached keys as the anchor vector, generally outperforms other variants of KeyDiff that use different similarity metrics or anchor vectors.
+  • Reasoning.
+KeyDiff performs competitively on the Math-500 reasoning benchmark with other eviction methods using the DeepSeek-R1-Distill-Qwen-7B and Llama-8B, and shows near eviction-free baseline performance when augmented with a sliding window (Section 4.3) for DeepSeek-R1-Distill-Llama-8B.
+  • Ablation Study.
+We perform an ablation study on the main parameters of KeyDiff and show that utilizing negative cosine similarity as the eviction criteria and the mean of cached keys as the anchor vector performs best.
 (Section 4.4).
+  • Efficiency. We compare the end-to-end inference latency of KeyDiff, [20] and [24] and observe a 30% latency improvement with KeyDiff (Section 4.5).
 Experimental Setup
-We apply several cache eviction methods to several decoder-only transformer-based language models, including Llama $3.1$-$8$B-Instruct (Dubey et al., 2024), Llama $3.2$-$3$B-Instruct, and Qwen $2.5$-$3$B/7B-Instruct.
-We evaluate these models using H2O (Zhang et al., 2024b), TOVA (Oren et al., 2024), SnapKV (Li et al., 2024), and StreamingLLM (Xiao et al., 2024), (or “sink attention”) cache eviction policies, along with the eviction-free model performance as a baseline.
+We apply several cache eviction methods to several decoder-only transformer-based language models, including Llama $3.1$-$8$B-Instruct [10], Llama $3.2$-$3$B-Instruct [10], and Qwen $2.5$-$3$B/7B-Instruct [38].
+We evaluate these models using H2O [43], TOVA [24], SnapKV [20], and StreamingLLM [35], (or “sink attention") cache eviction policies, along with the eviction-free model as a baseline.
 We simulate a resource constrained environment by processing prompts and generating responses using Equation 6, with a block size of $B=128$ for prompt processing and $B=1$ for token generation using greedy decoding for all experiments.
 We denote the cache budgets of 2048, 4096, 6144 and 8192 as 2K, 4K, 6K and 8K, respectively.
-4.1 Needle In a Haystack
-To compare the impact of various cache eviction policies on fact retrieval, we conduct the “Needle In a Haystack” test (Liu et al., 2024a; Kamradt, 2023).
-This test embeds specific information (“needle”) at different points within a body of unrelated text (“haystack”); finding and retaining the needle is challenging for eviction policies, which can’t know during eviction what information must be retained.
-The results are shown in Figure 6, where we show the recall accuracy of Llama3.23.23.23.2-$3$B-Instruct across different document lengths (x-axis) and needle depths (y-axis) with a cache size of $6$K.
-KeyDiff performs similarly to TOVA and sink attention for shorter documents and outperforms both methods as the document length increases.
-(a) Sink
-(b) TOVA
+(a) TOVA
+(b) SnapKV
 (c) KeyDiff
 Figure 6: Accuracy across document length and needle depth for needle in a haystack test. Cache size is $6$K with $B=128$.
+4.1 Needle In a Haystack
+To compare the impact of various cache eviction policies on fact retrieval, we conduct the “Needle In a Haystack" test [21, 16].
+This test embeds specific information (“needle") at different points within a body of unrelated text (“haystack"); finding and retaining the needle is challenging for eviction policies, which can’t know what information must be retained during block prompt processing.
+The results are shown in Figure 6 and Figure 19, where we show the recall accuracy of Llama3.23.23.23.2-$3$B-Instruct across different document lengths (x-axis) and needle depths (y-axis) with a cache size of $6$K.
+KeyDiff performs similarly to TOVA, SnapKV and sink attention for shorter documents and outperforms all three methods as the document length increases.
 Table 1: Llama-3.1-8B/3.2-3B-Instruct LongBench results with $B=128$ (Higher is better).
@@
 4.2 LongBench
-LongBench (Bai et al., 2024) is a bilingual, multi-task benchmark suite for LLMs, providing a comprehensive stress test for long prompt inputs.
-LongBench is useful to evaluate cache eviction methods in a resource constrained environments with a fixed memory budget: 51% of prompts are longer than the largest KV cache size of 8K.
-For cache budgets of 6k and 8k tokens, prompts in LongBench are compressed by 33% and 23% respectively on average, (see Section B.3 for more detail.)
+LongBench [4] is a bilingual, multi-task benchmark suite for LLMs, providing a comprehensive stress test for long prompt inputs.
+LongBench is useful for evaluating cache eviction methods in a resource constrained environments with a fixed memory budget: 51% of prompts are longer than the largest KV cache size of 8K.
+For cache budgets of 6k and 8k tokens, prompts in LongBench are compressed by 33% and 23% respectively on average, (see Section F.3 for more detail.)
 Table 1 summarizes the evaluation results of Llama 3.1-8B-Instruct and Llama 3.2-3B-Instruct on the English subset of LongBench with 2K, 4K, 6K, and 8K cache budgets using various eviction policies with block prompt processing enabled with $B=128$.
-We present the full evaluation results in Table 11, and those of Qwen 2.5-3B/7B-Instruct in Table 9.
 As shown in Table 1, KeyDiff outperforms other eviction strategies across most tasks, even demonstrating better performance with smaller cache budgets.
-KeyDiff shows significant a improvement on the PassageRetrieval-en (PR-en) dataset, which tests whether long-term dependencies within a long prompt can be correctly recognized (Bai et al., 2024), while achieving near full-context model performance even with the smallest budget.
-We observed similar trends in the full LongBench task suite as shown in Table 11.
-Results for Full Prompt Processing
-We also evaluate the methods in more standard inference environments without block prompt processing (i.e., setting the block size of the prompt to $B=\infty$) and summarize the results in Table 7.
+KeyDiff shows significant a improvement on the PassageRetrieval-en (PR-en) dataset, which tests whether long-term dependencies within a long prompt can be correctly recognized [4], while achieving near full-context model performance even with the smallest budget.
+Adding a sliding window to KeyDiff improves coding task performance (Table 13).
+We observed similar trends in the full LongBench task suite as shown in Table 10 and in the additional results in Appendix F.
 KeyDiff exhibits similar or better performance compared to competing methods.
@@
 This result supports our hypothesis: an eviction scheme robust to changes in the scope of comparison among tokens is essential in memory constrained environments where token-wise attention weight can’t be fully materialized.
+Additional Results
+We present the full evaluation results in Table 10 and more complete comparisons on LongBench in Appendix F, such as: standard prompt processing with a single large block (i.e.
+$B=\infty$) in Table 9; eviction method performance with Qwen 2.5-3B/7B-Instruct in Table 11; performance behavior with block sizes $B=[64,256]$ in Table 12; and performance on KeyDiff combined with a sliding window as described in Section 3.2.
+We also compare against the $L_{2}$-norm minimizing eviction method of [9] in Table 10.
+4.3 Math-500 Reasoning Benchmark
+Reasoning is an important long-context task for LLMs.
+Unlike other long-context use cases, reasoning typically involves a relatively short prompt followed by a long generation, which presents unique challenges for token eviction methods.
+To evaluate the effectiveness of token eviction methods, we apply KeyDiff and SnapKV to the DeepSeek-R1-Distill-Qwen-7B and Llama-8B distilled models [12], and assess their performance on the Math-500 reasoning benchmark [13].
+Surprisingly, we found that Llama equipped with KeyDiff and a moderate KV cache budget performs comparably to, or slightly better than, the eviction-free baseline, while also outperforming SnapKV.
+We kindly refer the reader to Appendix E for additional details on the reasoning task evaluation.
 4.4 Ablation Study
 We evaluate the design choices of KeyDiff, including the similarity metrics and the choice of the anchor vector, and validate the efficacy of KeyDiff.
-Table 2: Ablation study results about anchor vector. Average of Full Longbench results for Llama 3.2-3B-Instruct. KeyDiff results match Table 11. (Higher is better)
-2K 4K 6K 8K
-KeyDiff 35.09 40.28 41.88 42.40
-$\operatorname{Pairwise}$ 35.24 40.61 41.87 42.45
-$\operatorname{Median}$ 35.43 40.67 41.89 42.26
-Selecting the Anchor Vector
-We have mainly evaluated KeyDiff using the method described in Equation 8.
-Scores to determine eviction are measured via cosine similarity with an anchor vector which can be computed in several ways.
-We run LongBench on Llama3.2-3B-Instruct with eviction policies using the following anchor choices: pairwise cosine similarity from Equation 7, denoted $\operatorname{Pairwise}$; KeyDiff, using the mean of all normalized keys as an anchor, and using the median of keys as an anchor, denoted $\operatorname{Median}$.
-Table 2 summarizes the average LongBench accuracy for the different methods to Llama 3.2-3B-Instruct.
-KeyDiff shows similar average scores to $\operatorname{Pairwise}$.
-Additionally, KeyDiff and $\operatorname{Median}$ show similar scores, demonstrating that KeyDiff is robust to the selection of the anchor design.
-Table 3: Ablation study results about distance metric. Average of Full Longbench results of Llama 3.2-3B-Instruct. KeyDiff results match Table 11. (Higher is better)
-2K 4K 6K 8K
-KeyDiff 35.09 40.28 41.88 42.40
-$\operatorname{DotProd}$ 30.14 38.01 41.09 42.23
-$\operatorname{Euclidean}$ 13.68 21.06 25.91 29.53
-Selecting the Similarity Metric
-We use cosine similarity as the scoring metric for eviction in KeyDiff based on our discussion in Section 2.2.
-This could be replaced with other metrics like the dot product or Euclidean distance.
-We evaluate KeyDiff variants using dot product and Euclidean distance as the similarity metric, denoted $\operatorname{DotProd}$ and $\operatorname{Euclidean}$ respectively, and report the results in Table 3.
-KeyDiff and $\operatorname{DotProd}$ show similar performance for 6K and 8K budgets.
-However, KeyDiff outperforms $\operatorname{DotProd}$ for smaller cache sizes.
-This implies that considering both the direction and the magnitude of the keys to compute similarity are important for identifying the tokens to evict.
-On the other hand, $\operatorname{Euclidean}$ shows a significant performance drop relative to KeyDiff.
+We provide a full description of the test setup in Appendix G and summarize the findings here:
+  • KeyDiff anchor choice does not greatly impact benchmark accuracies (See Table 14).
+  • KeyDiff using cosine similarity as the distance metric outperforms other metrics (See Table 15)
+4.5 Latency and Complexity
+Additionally, in order to demonstrate that KeyDiff decreases end-to-end inference latency, we measured time to first token for the Llama 3.2 3B instruct model using different block prompt processing sizes and cache strategies.
+These results are visualized in Figures 17 and 7.
+Since KeyDiff does not require attention weight materialization, FlashAttention [8] can be used, resulting in up to 30% lower latency than TOVA and SnapKV.
+We compare the complexity of KeyDiff with competitors in Appendix B and perform a complete FLOP count in Section B.1.
+Figure 7: Time-to-first-token (TTFT) for Llama 3.2-3B using Flash Attention with different eviction strategies with block prompt processing sizes $64$, $128$, and $256$.
 5 Related Work
@@
 This characteristic allows sparse approximation techniques to reduce the computational cost of attention.
-Similar to PagedAttention (Kwon et al., 2023), Tang et al.
-(2024) estimates the importance of a page (a contiguous set of keys) to a given query, whereas Rehg (2024) further refined the budgets in a per-head manner.
-On the contrary, sample-based methods (Zhu et al., 2024; Ribar et al., ) attempt to approximate token importance by inspecting the attention scores from the last few queries or certain query channel dimensions.
+Similar to PagedAttention [18], Tang et al.
+[30] estimates the importance of a page (a contiguous set of keys) to a given query, whereas Rehg [26] further refined the budgets in a per-head manner.
+On the contrary, sample-based methods [44, 27] attempt to approximate token importance by inspecting the attention scores from the last few queries or certain query channel dimensions.
 Despite their effectiveness in reducing computational costs, these methods do not address the memory overhead of the KV cache, which typically retains all KVs.
 KV Cache Compression
-Different approaches to compress the KV cache include architecture modification such as GQA (Ainslie et al., 2023), which shares a KV cache across a small number of heads.
-Other techniques to compress the KV cache include quantization such as in (Hooper et al., 2024; Liu et al., 2024b; Zhang et al., 2024a) in which the authors use various techniques to take advantage of existing patterns to efficiently quantize and compress the KV cache.
-More related to our work (Yang et al., 2024b) uses a scoring mechanism to determine the precision of the quantization for different tokens.
+Different approaches to compress the KV cache include architecture modification such as GQA [2], which shares a KV cache across a small number of heads.
+Other techniques to compress the KV cache include quantization such as in [15, 23, 42] in which the authors use various techniques to take advantage of existing patterns to efficiently quantize and compress the KV cache.
+More related to our work [40] uses a scoring mechanism to determine the precision of the quantization for different tokens.
 Token Eviction Methods
@@
 Inspired by our observation that distinctive keys tend to have high attention scores, we propose KeyDiff, a training-free KV cache eviction method based on key similarity that enables large language models to operate in memory and compute constrained environments.
-We justify KeyDiff by showing that it computes the optimal solution for an optimization problem minimizing the pairwise cosine similarity among keys in the KV cache.
+We justify KeyDiff by showing that it minimizes the pairwise cosine similarity among keys in the KV cache, maximizing the aforementioned diversity.
 KeyDiff significantly outperforms state-of-the-art KV cache eviction methods under similar memory constraints, with only a 1.5% and 0.04% accuracy drop from the non-evicting baseline while achieving 33% and 23% KV cache memory reduction on LongBench.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+Similar to other token eviction methods, KeyDiff is primarily designed and evaluated for the GQA attention mechanism used in models such as Llama and Qwen.
+In future work, we plan to extend KeyDiff for seamless integration with other attention variants, such as Multi-Head Latent Attention [12].
 Supplementary Material
@@
 As mentioned in Equation 5, we can specify attention-based eviction methods under the unified framework with proper selection of the aggregation function $\phi(A)$ as follows:
-  • TOVA (Oren et al., 2024): $\phi^{\mathrm{TOVA}}(A)=A_{-1,:}$,
-  • H2O (Zhang et al., 2024b): $\phi^{\mathrm{H2O}}(A)=A_{\mathrm{prev}}+A^{\top}1$,
-  • SnapKV (Xiao et al., 2024): $\phi^{\mathrm{SnapKV}}(A)=(A^{\top}1)\ast K$, where $K$ is a vector of $\frac{1}{k}$ and $k$ is the kernel size of average smoothing.
-A.2 Runtime and Memory Complexity
+  • TOVA [24]: $\phi^{\mathrm{TOVA}}(A)=A_{-1,:}$,
+  • H2O [43]: $\phi^{\mathrm{H2O}}(A)=A_{\mathrm{prev}}+A^{\top}1$,
+  • SnapKV [35]: $\phi^{\mathrm{SnapKV}}(A)=(A^{\top}1)\ast K$, where $K$ is a vector of $\frac{1}{k}$ and $k$ is the kernel size of average smoothing.
+Appendix B Runtime and Memory Complexity
 We analyze the runtime and memory complexity for the prominent KV cache eviction algorithms in Table 2.
@@
 Sink $\mathcal{O}(N)$ $\mathcal{O}(1)$
-A.2.1 FLOP count of KeyDiff
-Regarding the FLOP count of KeyDiff: the bulk of the computation in KeyDiff (neglecting the ‘topk‘ operator) is the following two expressions:
+B.1 FLOP count of KeyDiff
+The bulk of the computation in KeyDiff (neglecting the ‘topk‘ operator) is the following two expressions:
   • $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\frac{k_{i}}{\|k_{i}\|}$
@@
 Also from above, we have that computing $\|\mu(\hat{K})\|$ requires $d$ multiplications, $d-1$ additions and one square root.
-We reuse the computation of $\|k_{i}\|$ from the previous step and compute $\|k_{i}\|\|\mu(\hat{K})\|$ in $n$ multiplications and $\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)$ with more $n$ additions (assuming boolean comparison is equals addition in cost).
+We reuse the computation of $\|k_{i}\|$ from the previous step and compute $\|k_{i}\|\|\mu(\hat{K})\|$ in $n$ multiplications and $\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)$ with more $n$ additions (assuming boolean comparison equals addition in cost).
 We can then divide through to compute $\frac{\mu(\hat{K})\cdot k_{i}}{\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)}$ with $n$ divisions.
@@
     $n+1$ square roots,
-If we declare additions and multiplications are roughly one FLOP (i.e.
-can be computed in one cycle), division is roughly 5 FLOPs and square root is roughly 8 FLOPs, we arrive at a final FLOP count of:
-$(3nd+d+n)+(3nd-n-1)+5*(2n+2)+8*(n+1)=(6d+18)n+d+17.$ (A.1)
+If, based on x86 instruction tables, we declare additions and square roots cost one FLOP (i.e.
+can be computed in one cycle), multiplications cost three FLOPs, division is roughly 47 FLOPs, we arrive at a final FLOP count of:
+$943(3nd+d+n)+(3nd-n-1)+47*(2n+2)+(n+1)=(12d+97)n+3d+$ 94 . (A.1)
 This is linear in the number of keys $n$ with a small constant, relative to the quadratic complexity of the attention operator.
-A.3 Additional PCA Visualizations
+Appendix C KeyDiff: A Theoretical Perspective
+C.1 Additional PCA Visualizations
 In order to demonstrate the phenomena in Figure 11 persists across all heads and layers, we have included several more visualizations as seen in Figure 8, Figure 9, and Figure 10.
@@
 Figure 10: Keys taken from layer $8$ and head $1$ of Llama3.2-3B-Instruct
-Appendix B Additional discussion for LongBench
-B.1 Empirical Motivation for KeyDiff Setup
+(a) Layer 0 Head 1
+(b) Layer 10 Head 1
+(c) Layer 20 Head 1
+(d) Layer 25 Head 1
+Figure 11: PCA plots of Query and Keys from Llama-3.2-3B-Instruct
+C.2 Derivation of KeyDiff from an Optimization Perspective
+[…]
+This is a combinatorial optimization problem, which is difficult to solve efficiently, particularly during inference.
+However, we can relax Section C.2 to produce a more tractable approximation to the original problem.
+First, we rewrite Section C.2 by normalizing keys $k_{i}$ such that $\hat{k}_{i}=\frac{k_{i}}{\|k_{i}\|}$ resulting in:
+$\sum\limits_{i\in S}\sum\limits_{j\in S}\hat{k}_{i}\cdot\hat{k}_{j}$ $=\sum\limits_{i\in S}\hat{k}_{i}\cdot\left(\sum\limits_{j\in S}\hat{k}_{j}\right)=\sum\limits_{i\in S}\langle\hat{k}_{i},N\mu(\hat{K}_{S})\rangle$
+[…]
+Key Cache Diversity and KeyDiff
+To empirically verify KeyDiff’s ability to retain diverse keys, we apply PCA to the keys computed in an attention block of Llama 3.2-3B-Instruct after evaluating a long context prompt.
+We visualize the distribution of retained and evicted keys by applying sink attention [35], TOVA [24], and KeyDiff as eviction policies in Figure 11.
+Visual observation reveals the tokens retained by sink attention and TOVA tend to tightly cluster together while KeyDiff’s retained tokens are more evenly distributed.
+This observation generalizes across heads and layers, as shown in Section C.1.
+We also visualize the log determinant of the Gram matrix of the key cache $\log(\det(KK^{T}))$ generated using different eviction policies in Figure 12.
+This quantity corresponds to the volume of space spanned by the keys in $\mathcal{C}$.
+The distribution of volumes for KeyDiff attains higher values, indicating that the retained keys are more distinctive than TOVA and sink, corroborating the results of Figure 11.
+Details of how these plots were generated are discussed in Section F.1.
+Figure 12: Distribution of $\log\left(\textrm{det}\left(KK^{T}\right)\right)$ from the Qasper dataset in LongBench.
+Larger values mean more of the key space is spanned by the key cache.
+KeyDiff retains keys that span a greater volume of the ambient space than TOVA or sink attention.
+C.3 Attention Sinks and Approximate Collinearity
+(a) Keys
+(b) Queries
+Figure 13: Cosine similarity between keys and their mean (left) and queries and their mean (right) across heads and layers
+Figure 14: Cosine similarity of mean key and mean query for each head and layer.
+To better understand why there exists a negative correlation between cosine similarity between keys and attention scores, we look to recent research that seeks to the importance of attention sinks in decoder-based LLMs.
+The authors in [5] show that attention sinks emerge from training decoder-based LLMs since they can denoise the model and prevent rank collapse by limiting over mixing in attention heads.
+Moreover, attention patterns in decoder based models demonstrate that most attention logits are quite small (and almost always negative) for most keys and queries.
+This allows the attention sink to have high attention activation, preventing over mixing in addition to allowing the heads to specialize and selectively identify important tokens.
+At the same time, the results in [11] suggest that hidden states, keys and queries are all approximately collinear in the sense that $\textrm{CosSim}(x_{i},x_{j})\gg 0$.
+In geometric terms, this means that most key tokens and query tokens lie in the same direction in Euclidean space.
+Our own results, seen in Figure 13 demonstrate that this is the case for both keys and queries.
+Our experiments show that most keys and queries lie within a small angular distance from the mean key and query.
+More than this, we see that across all heads, the mean key and mean query have negative cosine similarity Figure 14.
+Moreover, as seen in Figure 15, we find that the norms of keys and queries are tightly clustered around a relatively fixed value.
+This means that variations in the norm of key and query tokens have less impact on the magnitude of attention scores than their direction.
+These three observations indicate that most keys and queries combine to create uniformly small attention logits, and that larger attention weights are constructed by projecting keys closer to the direction of the mean query.
+This appears to be the fundamental mechanism through which over mixing is prevented: if most attention activations are very small, each head can increase the activations of a small number of keys across most queries selectively projecting them to be more aligned with the distribution of queries.
+This hypothesis is further supported by the fact that sink tokens themselves often have small norm, which results in an approximate no-op in the attention head as in [5], however in this case, the only way for a key corresponding to a sink token to have high attention scores is if it is as parallel as possible to the set of query tokens.
+Figure 15: Distribution of $L^{2}$ norms for keys and queries across heads and layers.
+Figure 16: Cosine similarity between highest KeyDiff scoring key token with mean query.
+To verify the above hypothesis, we show Figure 16 that keys which have maximum angular difference from the mean key are aligned with the mean query, resulting in very large attention weights.
+This demonstrates how LLMs exploit the geometry of the hidden states and projections to limit over mixing, and selectively identify important tokens.
+To summarize, we have for all attention heads in decoder based transformer models:
+  • The majority of keys and queries are approximately collinear with their mean.
+  • Mean keys and mean queries have negative cosine similarity across all heads.
+  • Most keys and queries have a similar L2 norm.
+  • Decoder based attention heads can selectively increase attention weights for a fixed key by aligning it with the mean query.
+  • Key token importance can hence be measured by the angular distance between a key and the mean key.
+We can show mathematically with some reasonable assumptions based on the above observations that key tokens with persistently high attention scores must be geometrically aligned with queries.
+Theorem C.1.
+Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$.
+Without loss of generality suppose $||q||=1$, the scaling parameter is $1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$:
+$w=\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}.$
+Then
+$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1\leq\textrm{CosSim}({k^{*}},q)$
+Proof.
+To show this we have that
+$w$ $=\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}$
+$w\left(\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)\right)$ $=\exp({k^{*}}^{\top}q)$
+$w\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)$ $=(1-w)\exp({k^{*}}^{\top}q)$
+Note that $-M\leq k_{i}^{\top}q\leq M$ and hence
+$w\sum\limits_{i=1}^{n}\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
+$wn\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
+$\frac{wn\exp(-M)}{1-w}$ $\leq\exp({k^{*}}^{\top}q)$
+$\ \log(wn)-M-\log(1-w)$ $\leq M\textrm{CosSim}({k^{*}},q)$
+$\frac{\log(wn)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
+$\frac{\log(\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}n)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
+$\frac{\log(\frac{\exp({k^{*}}^{\top}q)}{(n+1)\exp(M)}n)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
+$\frac{-M\textrm{CosSim}(k^{*},q)+\log(\frac{n}{n+1})-2M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
+$\frac{\log(\frac{n}{n+1})-2M-\log(1-w)}{M}$ $\leq 2\textrm{CosSim}({k^{*}},q)$
+$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1$ $\leq\textrm{CosSim}({k^{*}},q)$
+Taking the limit as $n\to\infty$ produces Lemma 3.1
+The above proof demonstrates that as long as the norms of the keys are bounded, in order for an attention head to be able to freely allocate $w$ attention weight to a given key ${k^{*}}$, the cosine similarity between ${k^{*}}$ and $q$ must be high, therefore ${k^{*}}$ and $q$ must be approximately collinear.
+Using this result, we can also show that as long as the cosine similarity between ${k^{*}}$ and $q$ is high, while the cosine similarity between $\bar{k}$ and $q$ is low, $\textrm{CosSim}({k^{*}},\bar{k})$ is small.
+Note that, since empirical results demonstrate that most keys have high cosine similarity with their mean $\bar{k}$, a key with high importance, approximately collinear to $q$, will also have low cosine similarity to $\bar{k}$.
+Generally, this also suggests that key tokens with low cosine similarity to $\bar{k}$ have greater importance.
+In order to show this, we need the following auxiliary result.
+Lemma C.2.
+Suppose $\{x_{1},...,x_{n}\}$ is an orthonormal basis of $\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$. Define $\alpha_{i}=\textrm{CosSim}(x_{i},y)$. Then $\sum\limits_{i=1}^{n}\alpha_{i}^{2}=1$.
+Proof.
+Note that $\alpha_{i}=\frac{y^{\top}x_{i}}{||y||||x_{i}||}=\frac{y^{\top}x_{i}}{||y||}$. If we expand $\frac{y}{||y||}$ in the basis $\{x_{1},...,x_{n}\}$ we see that
+$\frac{y}{||y||}$ $=\sum\limits_{i=1}^{n}\left({\frac{y}{||y||}}^{\top}x_{i}\right)x_{i}$
+$=\sum\limits_{i=1}^{n}\alpha_{i}x_{i}$
+But then, we know that since $\left|\left|\frac{y}{||y||}\right|\right|_{2}^{2}=1$, then $\left\langle\sum\limits_{i=1}^{n}\alpha_{i}x_{i},\sum\limits_{i=1}^{n}\alpha_{i}x_{i}\right\rangle=1$.
+But we have
+$\left\langle\sum\limits_{i=1}^{n}\alpha_{i}x_{i},\sum\limits_{i=1}^{n}\alpha_{i}x_{i}\right\rangle$ $=\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\langle\alpha_{i}x_{i},\alpha_{j}x_{j}\rangle$
+$=\sum\limits_{i=1}^{n}\langle\alpha_{i}x_{i},\alpha_{i}x_{i}\rangle$
+$=\sum\limits_{i=1}^{n}\alpha_{i}^{2}$
+proving the result. ∎
+Theorem C.3.
+Consider tokens ${k^{*}}$, $q$, $\bar{k}$ as above where $\bar{k}$ is the average of the keys tokens.
+Suppose $\textrm{CosSim}({k^{*}},q)=\beta_{q}>0$ and $\textrm{CosSim}(\bar{k},q)=\alpha_{q}<0$.
+Then $\textrm{CosSim}(\bar{k},{k^{*}})\leq 1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$.
+Proof.
+Consider the cosine similarity of $\bar{k}$ and ${k^{*}}$:
+$\textrm{CosSim}(\bar{k},{k^{*}})$ $=\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$
+expand $\bar{k}$ in an orthonormal basis which contains $q$, $\{q,r_{1},...,r_{n-1}\}$ such that
+$\bar{k}=||\bar{k}||\left(\alpha_{q}q+\sum\limits_{i=1}^{n-1}\alpha_{i}r_{i}\right)$
+where $\alpha_{i}=\textrm{CosSim}(\bar{k},r_{i})$.
+Additionally, define $\beta_{i}=\textrm{CosSim}({k^{*}},r_{i})$ and note that by the definition of an orthonormal basis and the cosine similarity operation, using the result from Lemma C.2 we have that $\alpha_{q}^{2}+\sum\limits_{i=1}^{n-1}\alpha_{i}^{2}=1$ and that $\beta_{q}^{2}+\sum\limits_{i=1}^{n-1}\beta_{i}^{2}=1$.
+Now we have that
+$\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$ $=\frac{{k^{*}}^{\top}\left(||\bar{k}||\alpha_{q}q+\sum\limits_{i=1}^{n-1}||\bar{k}||\alpha_{i}r_{i}\right)}{||{k^{*}}||||\bar{k}||}$
+$=\alpha_{q}\beta_{q}+\frac{1}{||{k^{*}}||}\sum\limits_{i=1}^{n-1}\alpha_{i}{k^{*}}^{\top}r_{i}$
+$=\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}\alpha_{i}\beta_{i}$
+$\leq\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}|\alpha_{i}||\beta_{i}|$
+Applying Young’s inequality we obtain
+$\leq\alpha_{q}\beta_{q}+\frac{1}{2}\sum\limits_{i=1}^{n-1}\alpha_{q}^{2}+\beta_{q}^{2}$
+$=\alpha_{q}\beta_{q}+\frac{1}{2}(1-\alpha_{q}^{2})+\frac{1}{2}(1-\beta_{q}^{2})$
+$=1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$
+Note that on the domain $\alpha_{q}\in[-1,0)$, $\beta_{q}\in(0,1]$ the function $1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$ is bounded above by $1$ and decreasing to $-\frac{1}{2}$ as $\alpha\rightarrow-1$.
+Hence, the smaller $\textrm{CosSim}(\bar{k},q)=\alpha_{q}$ is, the smaller $\textrm{CosSim}(\bar{k},{k^{*}})$ must be.
+Appendix D TTFT Analysis
+We have measured end-to-end inference latency (measured as time to first token (TTFT)) for Llama 3.2-3B via the standard Huggingface API when the model is using eager attention and FlashAttention as in Figure 17 and Figure 7.
+Tests are performed on NVIDIA A100 80GB GPUs.
+We test various block sizes and KV cache budgets.
+KeyDiff outperforms TOVA and SnapKV with FlashAttention as well as with eager attention for large cache budgets.
+We can see inference latency with KeyDiff is independent of block size when using FlashAttention because of KeyDiff’s linear complexity and its lack of required materialized attention weights.
+Figure 17: TTFT for eager attention with different cache eviction strategies using block size $64$, $128$, $256$ for block prompt processing.
+Appendix E Math-500 reasoning benchmark
+In order to measure the effectiveness of different caching methods on reasoning tasks, we tested several different model using various caching algorithms on the Math 500 reasoning benchmark.
+Specifically, we test KeyDiff with a sliding window whose window size is $20\%$ of the KV cache budget, and SnapKV on the DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1-Distill-Qwen-7B models.
+We randomly sample 5 responses with $\textrm{TopP}=0.95$, $\textrm{Temperature}=0.95$ with the 4096, 6144, and 32,786 max generation lengths.
+The reported scores are the average of accuracies over the random samples.
+Table 3 summarizes the Math-500 evaluation results for Llama-8B.
+As shown in the table, token eviction methods generally perform well even with KV cache budgets that are strictly smaller than the maximum generation length.
+Surprisingly, we found that KeyDiff slightly outperforms baseline methods in certain configurations (e.g., KeyDiff with a 2K budget for 4K generation length, and KeyDiff with a 4K budget for 8K generation length).
+To further analyze the effectiveness of token eviction, we measure accuracy in cases where the context length of the baseline method (i.e., prompt length + generation length) exceeds the available KV cache budget.
+As shown in Table 4, for samples where eviction is actively triggered, KeyDiff continues to outperform the token eviction baseline (SnapKV), and often achieves accuracy close to or slightly better than the non-evicting baseline.
+We also conducted a similar evaluation on DeepSeek-R1-Distill-Qwen-7B and observed a slight performance degradation for token eviction methods compared to full KV cache baselines (See Tables 5 and 6.) However, KeyDiff still demonstrates comparable performance to SnapKV overall.
+This discrepancy may stem from architectural differences that Llama uses a lower GQA [2] ratio than Qwen, which results in more information compression in the KV cache.
+We hypothesize that models with more compression like Qwen are more sensitive to eviction since each evicted token contains more information in Qwen than Llama by design.
+Table 3: Math 500 results on DeepSeek-R1-Llama-8B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
+Method Max Gen. Length Budget Flex Match Exact Match Avg. Gen. Length
+Full 4K N/A 0.711 0.537 2769
+KeyDiff 4K 1024 0.695 0.531 2753
+2048 0.720 0.546 2740
+SnapKV 4K 1024 0.689 0.529 2759
+2048 0.714 0.544 2757
+Full 8K N/A 0.840 0.628 3812
+KeyDiff 8K 2048 0.819 0.617 3888
+4096 0.844 0.634 3805
+SnapKV 8K 2048 0.805 0.610 3898
+4096 0.828 0.627 3898
+Full 32K N/A 0.898 0.668 6869
+KeyDiff 32K 2048 0.883 0.662 7678
+4096 0.894 0.668 7312
+8192 0.894 0.667 7096
+SnapKV 32K 2048 0.849 0.641 7509
+4096 0.884 0.661 7218
+8192 0.893 0.665 7005
+Table 4: Math 500 results on DeepSeek-R1-Llama-8B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
+Max Gen. Length $=$ 4K
+Num Tokens $>$ 1K (497/500 samples)
+Budget Flex Exact
+Full N/A 0.709 0.534
+KeyDiff 1024 0.693 0.528
+SnapKV 1024 0.687 0.526
+Num Tokens $>$ 2K (240/500 samples)
+Full N/A 0.604 0.449
+KeyDiff 2048 0.618 0.463
+SnapKV 2048 0.610 0.458
+Max Gen. Length $=$ 8K
+Num Tokens $>$ 2K (353/500 samples)
+Budget Flex Exact
+Full N/A 0.783 0.584
+KeyDiff 2048 0.756 0.570
+SnapKV 2048 0.736 0.560
+Num Tokens $>$ 4K (195/500 samples)
+Full N/A 0.650 0.455
+KeyDiff 4096 0.662 0.470
+SnapKV 4096 0.637 0.453
+Max Gen. Length $=$ 32K
+Num Tokens $>$ 2K (353/500 samples)
+Budget Flex Exact
+Full N/A 0.783 0.584
+KeyDiff 2048 0.756 0.570
+SnapKV 2048 0.736 0.560
+Num Tokens $>$ 4K (195/500 samples)
+Full N/A 0.650 0.455
+KeyDiff 4096 0.662 0.470
+SnapKV 4096 0.637 0.453
+Num Tokens $>$ 8K (162/500 samples)
+Full N/A 0.793 0.550
+KeyDiff 8192 0.786 0.545
+SnapKV 8192 0.782 0.541
+Table 5: Math 500 results on DeepSeek-R1-Qwen-7B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
+Method Max Gen. Length Budget Flex Match Exact Match Avg. Gen. Length
+Full 4K N/A 0.764 0.579 2630
+KeyDiff 4K 1024 0.666 0.512 2692
+2048 0.749 0.570 2629
+SnapKV 4K 1024 0.692 0.533 2655
+2048 0.749 0.566 2637
+Full 8K N/A 0.877 0.658 3287
+KeyDiff 8K 2048 0.811 0.613 3348
+4096 0.867 0.647 3208
+SnapKV 8K 2048 0.812 0.612 3328
+4096 0.868 0.647 3214
+Full 32K N/A 0.923 0.682 4051
+KeyDiff 32K 2048 0.811 0.613 4322
+4096 0.897 0.647 3800
+8192 0.891 0.663 3741
+SnapKV 32K 2048 0.812 0.612 4279
+4096 0.868 0.647 3828
+8192 0.891 0.662 3808
+Table 6: Math 500 results on DeepSeek-R1-Qwen-7B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
+Max Gen. Length $=$ 4K
+Num Tokens $>$ 1K (497/500 samples)
+Budget Flex Exact
+Full N/A 0.762 0.578
+KeyDiff 1024 0.664 0.511
+SnapKV 1024 0.690 0.532
+Num Tokens $>$ 2K (328/500 samples)
+Full N/A 0.650 0.478
+KeyDiff 2048 0.628 0.464
+SnapKV 2048 0.629 0.459
+Max Gen. Length $=$ 8K
+Num Tokens $>$ 2K (317/500 samples)
+Budget Flex Exact
+Full N/A 0.816 0.603
+KeyDiff 2048 0.713 0.533
+SnapKV 2048 0.718 0.533
+Num Tokens $>$ 4K (179/500 samples)
+Full N/A 0.626 0.441
+KeyDiff 4096 0.611 0.419
+SnapKV 4096 0.615 0.422
+Max Gen. Length $=$ 32K
+Num Tokens $>$ 2K (360/500 samples)
+Budget Flex Exact
+Full N/A 0.889 0.642
+KeyDiff 2048 0.713 0.533
+SnapKV 2048 0.718 0.533
+Num Tokens $>$ 4K (141/500 samples)
+Full N/A 0.787 0.524
+KeyDiff 4096 0.611 0.419
+SnapKV 4096 0.615 0.422
+Num Tokens $>$ 8K (53/500 samples)
+Full N/A 0.550 0.275
+KeyDiff 8192 0.392 0.222
+SnapKV 8192 0.381 0.215
+Appendix F Additional discussion for LongBench
+F.1 Empirical Motivation for KeyDiff Setup
 To generate Figure 11, we used the first sample from the test split of the narrativeqa task in LongBench to prefill the KV cache of Llama3.2-3B-Instruct with a block size of $B=128$.
@@
 We applied PCA to the key cache and repeated the process for sink attention, TOVA and KeyDiff.
-To construct Figure 5, we sample 100 prompts from the Qasper dataset in LongBench (Bai et al., 2024), compute the log determinant of $KK^{T}$ of the keys in the KV caches of each head and layer of Llama 3.2-3B-Instruct using a cache budget of $N=2048$ and a block size of $B=128$, and plot the distribution in Figure 5.
+To construct Figure 12, we sample 100 prompts from the Qasper dataset in LongBench [4], compute the log determinant of $KK^{T}$ of the keys in the KV caches of each head and layer of Llama 3.2-3B-Instruct using a cache budget of $N=2048$ and a block size of $B=128$, and plot the distribution in Figure 12.
 We show this key distribution for sink attention, TOVA and KeyDiff.
-B.2 Experiment Setup
+F.2 Experiment Setup
 In this subsection, we provide the experimental setup for KeyDiff and the baselines for the LongBench experiments.
 The LongBench evaluation is conducted using the default parameters of the LongBench evaluator with predefined prompt templates.
-The prompt block size is set to $B=128$, and the same $B$ value is used for all experiments.
-For TOVA, H2O, ans SnapKV, the set of attention weights computed from a single key cache due to grouped query attention (Ainslie et al., 2023) is aggregated by taking the average over the attention weights.
+Tests are performed on NVIDIA A100 80GB GPUs.
+For TOVA, H2O, and SnapKV, the set of attention weights computed from a single key cache due to grouped query attention [2] is aggregated by taking the average over the attention weights.
 Additionally, only for SnapKV, we apply average smoothing to the attention score with a kernel size of 7 and keep the most recent 32 tokens in the cache, following the suggestion of the original paper.
 For Sink, we used the first four tokens as the attention sink, following the suggestion of the original paper.
-B.3 Longbench dataset statistics
+F.3 Longbench dataset statistics
 In this section, we provide the length statistics of the Longbench Benchmark and in-depth analysis of compression ratios for the given KV cache budgets, such as 2K, 4K, 6K, and 8K.
 Prompt lengths
-We measure the number of tokens in the samples using LLama tokenizer (Touvron et al., 2023). As shown in Figure 10, LongBench exhibits variability in sample length from the datasets.
+We measure the number of tokens in the samples using LLama tokenizer [31]. As shown in Figure 18, LongBench exhibits variability in sample length from the datasets.
 Figure 18: Histograms of sample lengths measured by number of tokens
@@
 The majority of other KV cache eviction studies assume an unconstrained memory footprint.
-Before they compress the cache by applying an eviction policy, they first set the target compression ratio and evict the appropriate number of KV pairs to satisfy the compression ratio (Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024).
+Before they compress the cache by applying an eviction policy, they first set the target compression ratio and evict the appropriate number of KV pairs to satisfy the compression ratio [43, 24, 20].
 On the other hand, we fix the KV cache size and ensure the number of cached tokens is less than or equal to the predefined cache size.
@@
 Table 8: Compression ratio of prompts w.r.t. various KV cache budgets
-B.4 Additional Results
+F.4 Additional Results
 Table 9: Resource unrestricted LongBench Results (Higher is better).
@@
 QA Summarization Fewshot Learning Synthetic Code Narrative QA Qasper MF-en MF-zh HotpotQA 2WikiMQA Musique DuReader GovReport QMSum MultiNews VCSum TREC TriviaQA SAMSum LSHT PCount PR-en PR-zh Lcc RB-P Avg.
-Llama3.1-8B 30.05 47.00 56.12 59.86 57.33 47.81 32.25 35.64 34.86 25.32 27.02 17.28 73.00 91.61 43.37 45.50 8.33 99.50 99.00 61.66 51.94 49.74 H2O 2K 1.74 21.15 25.33 21.65 26.11 24.15 8.78 5.90 2.17 2.70 16.78 3.97 44.00 29.36 7.62 14.50 2.25 5.88 4.00 40.15 12.14 15.25 4K 4.07 36.16 36.00 38.02 33.52 32.87 17.78 5.68 6.66 5.95 24.09 6.03 55.00 47.65 17.41 18.50 4.00 24.50 31.25 54.85 21.43 24.83 6K 8.52 43.31 44.80 46.24 40.03 42.46 21.68 7.33 11.85 8.78 26.03 7.82 62.00 56.39 25.72 18.00 5.75 45.50 90.00 58.62 29.53 33.35 8K 13.85 44.94 47.81 56.14 43.64 44.90 23.65 11.01 18.78 11.35 26.49 9.96 69.50 69.05 33.41 19.50 5.25 62.50 98.67 59.74 36.26 38.40 TOVA 2K 22.57 37.26 39.43 36.96 45.74 34.48 14.77 16.98 28.87 21.17 26.95 16.21 62.50 90.73 42.74 18.75 0.00 18.00 32.00 62.68 52.48 34.35 4K 22.68 44.55 47.87 51.16 46.76 44.54 20.56 22.50 30.95 22.13 26.96 16.75 61.50 90.56 43.27 25.25 3.00 43.50 84.00 61.62 53.40 41.12 6K 24.59 45.93 53.92 55.45 55.09 47.43 25.07 27.68 32.33 24.10 27.00 16.91 68.50 90.81 43.89 29.00 4.25 67.00 98.67 61.50 52.39 45.31 8K 24.86 46.78 54.83 57.95 54.52 49.00 26.40 31.15 33.44 24.76 27.00 17.33 71.00 91.11 43.29 33.25 6.25 87.00 98.67 61.49 51.79 47.23 Sink 2K 21.83 34.27 29.24 32.82 38.64 29.50 12.59 16.18 28.51 20.21 26.62 15.54 65.00 89.46 42.20 22.25 2.00 25.50 32.50 64.95 59.54 33.78 4K 22.94 43.01 39.08 46.16 44.04 41.39 19.09 16.54 31.08 21.57 26.78 16.73 70.00 91.53 42.29 29.25 3.00 38.50 71.00 62.12 58.84 39.76 6K 25.41 47.40 44.13 52.78 47.39 45.73 21.90 17.55 32.53 22.19 26.87 17.05 72.00 91.25 43.41 33.75 3.08 52.50 98.00 62.22 56.24 43.49 8K 23.53 46.63 48.68 55.77 49.61 47.16 21.14 19.54 33.10 23.20 26.92 16.91 72.00 91.29 43.79 37.00 3.25 66.00 99.00 62.18 56.43 44.91 SnapKV 2K 21.81 37.22 37.19 38.29 46.10 35.42 16.53 16.37 29.83 21.05 26.77 16.16 61.00 88.84 42.56 21.75 4.03 51.50 81.17 62.37 51.45 38.45 4K 24.79 44.22 47.30 50.27 48.49 46.73 20.55 22.04 32.19 22.68 26.95 16.95 67.50 90.98 43.14 25.00 5.17 89.50 96.67 61.44 51.20 44.46 6K 24.10 45.57 50.44 55.27 53.12 48.41 24.27 27.46 33.43 23.53 27.03 16.84 71.50 92.28 43.58 27.00 5.25 98.00 99.00 61.32 52.16 46.65 8K 25.15 46.55 53.39 57.65 56.00 48.75 27.82 32.66 33.67 24.85 27.01 17.37 72.50 91.78 43.54 33.75 5.08 100.00 98.67 61.48 51.41 48.05 KeyL2Norm 2K 8.66 36.63 41.70 37.70 33.75 32.25 5.39 17.73 19.64 14.96 26.69 11.02 63.00 58.94 28.45 22.50 3.05 17.75 20.13 52.40 25.63 27.52 4K 15.38 44.06 51.75 47.52 50.22 45.56 18.44 27.61 29.50 22.27 26.93 13.44 69.50 79.41 37.50 27.00 4.50 58.00 80.50 58.82 35.08 40.14 6K 21.75 45.63 55.06 53.77 52.93 47.70 25.63 32.17 32.66 24.85 26.94 15.12 70.00 86.89 40.51 34.50 5.25 75.00 98.00 60.98 43.14 45.17 8K 25.12 45.70 56.02 56.57 58.14 47.77 30.29 34.09 33.81 24.89 26.94 15.89 71.50 89.26 41.34 39.00 7.25 87.00 99.00 62.05 48.28 47.61 KeyDiff 2K 26.64 41.73 50.99 51.18 51.59 46.47 22.84 32.37 29.02 23.86 26.76 14.81 66.50 85.92 39.26 42.25 3.17 96.00 96.25 59.17 39.42 45.06 4K 28.70 45.62 56.06 56.83 54.58 49.31 28.25 33.06 32.30 25.03 27.07 16.32 70.00 90.85 42.84 44.50 4.21 99.00 97.67 60.80 48.00 48.14 6K 29.90 46.33 55.11 59.00 56.80 49.50 31.52 34.97 33.44 24.58 26.98 16.80 72.00 90.99 43.10 47.00 5.27 99.50 99.00 61.40 49.70 49.19 8K 33.57 46.77 55.48 59.16 56.87 49.37 30.88 34.54 34.17 25.12 27.01 17.13 72.50 92.28 42.81 46.50 5.83 99.50 98.67 61.48 50.90 49.55 Llama3.2-3B 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 H2O 2K 1.63 19.96 20.20 15.20 18.02 19.56 2.88 6.47 0.78 1.55 15.97 3.11 41.00 21.97 9.83 11.75 0.50 0.50 0.00 39.71 13.91 12.60 4K 2.92 31.94 33.23 33.25 24.49 28.08 7.55 10.10 5.44 6.30 22.77 4.81 53.00 38.85 20.33 15.50 1.50 7.50 6.25 51.23 22.94 20.38 6K 4.62 38.81 39.06 45.17 34.66 35.52 15.21 13.36 10.51 10.01 24.25 6.66 61.50 53.23 27.37 15.25 0.50 13.00 19.50 54.55 32.29 26.43 8K 9.65 39.66 43.20 52.60 38.09 40.41 21.46 18.55 17.80 13.28 24.67 9.12 70.00 64.30 32.19 17.00 2.00 24.50 21.50 55.00 39.09 31.15 TOVA 2K 17.14 30.14 32.44 31.64 35.96 30.05 13.08 9.62 26.15 19.70 25.04 15.47 56.50 87.81 40.48 16.75 2.50 11.50 6.50 55.51 52.36 29.35 4K 20.52 39.53 42.47 45.80 44.12 38.42 18.22 17.76 29.36 21.36 24.96 16.60 63.50 88.98 41.50 18.75 3.00 23.50 15.00 55.72 56.66 34.56 6K 20.22 39.78 45.86 52.93 49.08 41.54 20.43 24.78 30.50 22.17 25.11 16.37 66.50 89.00 42.50 21.00 4.00 46.50 20.50 55.57 57.53 37.71 8K 21.08 40.67 49.07 55.17 48.69 41.93 23.05 31.02 31.64 22.85 25.21 16.55 69.00 89.25 42.19 22.50 2.50 71.00 21.50 55.77 57.47 39.91 Sink 2K 16.85 30.69 26.58 27.32 33.26 25.27 13.82 9.38 26.74 19.15 25.15 15.88 65.00 86.17 40.79 19.50 1.50 19.50 8.50 56.65 52.73 29.54 4K 19.46 38.61 36.22 41.68 41.97 35.84 13.37 9.86 29.34 20.19 25.06 16.44 71.00 88.06 41.31 21.75 2.50 35.50 16.00 56.48 52.43 33.96 6K 19.33 40.29 37.95 49.68 46.48 40.29 15.31 11.10 30.43 21.35 25.14 16.64 71.50 88.93 42.04 23.50 3.50 47.00 19.50 56.55 54.11 36.22 8K 20.15 40.02 41.94 53.57 48.15 42.24 16.01 14.76 31.64 22.10 25.20 16.50 73.00 89.26 42.37 26.25 3.50 62.50 20.50 56.86 56.63 38.25 SnapKV 2K 17.38 31.37 31.48 29.65 37.77 30.05 11.54 9.66 27.03 19.93 24.97 15.97 59.00 88.13 40.48 16.25 3.50 32.50 9.00 56.32 55.91 30.85 4K 19.85 39.22 39.86 47.33 46.70 37.98 16.64 16.88 29.79 21.21 25.01 16.74 65.50 89.35 40.95 18.25 2.50 62.50 22.50 55.74 56.88 36.73 6K 20.83 39.65 44.48 51.84 49.30 40.18 20.28 25.32 31.27 22.73 25.09 16.81 69.00 89.95 41.47 18.75 4.00 85.00 20.50 55.69 57.82 39.52 8K 20.49 40.80 48.16 55.44 48.78 41.65 24.79 30.40 31.81 23.46 25.17 16.44 70.00 90.17 41.99 22.00 5.00 94.00 21.50 55.77 57.29 41.20 KeyL2Norm 2K 7.67 30.39 31.85 30.64 29.47 25.76 7.41 14.17 15.36 12.42 24.20 7.91 48.00 50.99 23.09 17.50 2.00 7.50 5.00 48.92 26.32 22.22 4K 12.92 37.59 44.71 43.85 38.89 33.42 12.41 22.42 24.63 19.27 24.77 11.37 63.00 72.51 31.75 19.00 3.87 9.50 11.00 55.82 40.08 30.13 6K 13.02 40.55 48.17 50.87 43.18 40.17 17.10 31.29 28.99 21.47 25.08 13.30 65.00 79.61 37.16 21.50 2.50 46.50 18.50 56.49 47.05 35.60 8K 15.72 40.54 47.88 54.29 49.29 43.79 22.22 33.18 31.86 22.50 25.19 14.28 70.00 84.92 39.45 23.00 1.50 69.00 20.50 56.82 50.73 38.89 KeyDiff 2K 18.29 36.65 45.44 47.47 46.09 35.41 13.79 28.89 28.16 21.45 25.01 13.56 60.00 85.24 37.00 24.88 1.00 60.50 12.00 54.13 42.01 35.09 4K 22.34 40.60 49.15 52.56 50.14 40.30 21.65 32.46 31.38 23.44 25.06 15.28 66.50 87.92 41.41 27.50 2.50 88.50 19.50 55.55 52.24 40.28 6K 22.29 40.68 50.14 54.51 51.74 42.19 24.83 34.64 32.39 23.53 25.19 15.88 71.00 90.02 42.00 28.75 3.00 95.00 21.50 55.86 54.39 41.88 8K 22.41 40.77 50.10 55.62 49.83 43.58 28.09 34.30 32.78 23.60 25.17 15.77 72.00 90.17 42.46 30.75 3.50 96.50 21.50 55.85 55.65 42.40
+Llama3.1-8B 30.05 47.00 56.12 59.86 57.33 47.81 32.25 35.64 34.86 25.32 27.02 17.28 73.00 91.61 43.37 45.50 8.33 99.50 99.00 61.66 51.94 49.74 H2O 2K 1.74 21.15 25.33 21.65 26.11 24.15 8.78 5.90 2.17 2.70 16.78 3.97 44.00 29.36 7.62 14.50 2.25 5.88 4.00 40.15 12.14 15.25 4K 4.07 36.16 36.00 38.02 33.52 32.87 17.78 5.68 6.66 5.95 24.09 6.03 55.00 47.65 17.41 18.50 4.00 24.50 31.25 54.85 21.43 24.83 6K 8.52 43.31 44.80 46.24 40.03 42.46 21.68 7.33 11.85 8.78 26.03 7.82 62.00 56.39 25.72 18.00 5.75 45.50 90.00 58.62 29.53 33.35 8K 13.85 44.94 47.81 56.14 43.64 44.90 23.65 11.01 18.78 11.35 26.49 9.96 69.50 69.05 33.41 19.50 5.25 62.50 98.67 59.74 36.26 38.40 TOVA 2K 22.57 37.26 39.43 36.96 45.74 34.48 14.77 16.98 28.87 21.17 26.95 16.21 62.50 90.73 42.74 18.75 0.00 18.00 32.00 62.68 52.48 34.35 4K 22.68 44.55 47.87 51.16 46.76 44.54 20.56 22.50 30.95 22.13 26.96 16.75 61.50 90.56 43.27 25.25 3.00 43.50 84.00 61.62 53.40 41.12 6K 24.59 45.93 53.92 55.45 55.09 47.43 25.07 27.68 32.33 24.10 27.00 16.91 68.50 90.81 43.89 29.00 4.25 67.00 98.67 61.50 52.39 45.31 8K 24.86 46.78 54.83 57.95 54.52 49.00 26.40 31.15 33.44 24.76 27.00 17.33 71.00 91.11 43.29 33.25 6.25 87.00 98.67 61.49 51.79 47.23 Sink 2K 21.83 34.27 29.24 32.82 38.64 29.50 12.59 16.18 28.51 20.21 26.62 15.54 65.00 89.46 42.20 22.25 2.00 25.50 32.50 64.95 59.54 33.78 4K 22.94 43.01 39.08 46.16 44.04 41.39 19.09 16.54 31.08 21.57 26.78 16.73 70.00 91.53 42.29 29.25 3.00 38.50 71.00 62.12 58.84 39.76 6K 25.41 47.40 44.13 52.78 47.39 45.73 21.90 17.55 32.53 22.19 26.87 17.05 72.00 91.25 43.41 33.75 3.08 52.50 98.00 62.22 56.24 43.49 8K 23.53 46.63 48.68 55.77 49.61 47.16 21.14 19.54 33.10 23.20 26.92 16.91 72.00 91.29 43.79 37.00 3.25 66.00 99.00 62.18 56.43 44.91 SnapKV 2K 21.81 37.22 37.19 38.29 46.10 35.42 16.53 16.37 29.83 21.05 26.77 16.16 61.00 88.84 42.56 21.75 4.03 51.50 81.17 62.37 51.45 38.45 4K 24.79 44.22 47.30 50.27 48.49 46.73 20.55 22.04 32.19 22.68 26.95 16.95 67.50 90.98 43.14 25.00 5.17 89.50 96.67 61.44 51.20 44.46 6K 24.10 45.57 50.44 55.27 53.12 48.41 24.27 27.46 33.43 23.53 27.03 16.84 71.50 92.28 43.58 27.00 5.25 98.00 99.00 61.32 52.16 46.65 8K 25.15 46.55 53.39 57.65 56.00 48.75 27.82 32.66 33.67 24.85 27.01 17.37 72.50 91.78 43.54 33.75 5.08 100.00 98.67 61.48 51.41 48.05 KeyL2Norm[9] 2K 8.66 36.63 41.70 37.70 33.75 32.25 5.39 17.73 19.64 14.96 26.69 11.02 63.00 58.94 28.45 22.50 3.05 17.75 20.13 52.40 25.63 27.52 4K 15.38 44.06 51.75 47.52 50.22 45.56 18.44 27.61 29.50 22.27 26.93 13.44 69.50 79.41 37.50 27.00 4.50 58.00 80.50 58.82 35.08 40.14 6K 21.75 45.63 55.06 53.77 52.93 47.70 25.63 32.17 32.66 24.85 26.94 15.12 70.00 86.89 40.51 34.50 5.25 75.00 98.00 60.98 43.14 45.17 8K 25.12 45.70 56.02 56.57 58.14 47.77 30.29 34.09 33.81 24.89 26.94 15.89 71.50 89.26 41.34 39.00 7.25 87.00 99.00 62.05 48.28 47.61 KeyDiff 2K 26.64 41.73 50.99 51.18 51.59 46.47 22.84 32.37 29.02 23.86 26.76 14.81 66.50 85.92 39.26 42.25 3.17 96.00 96.25 59.17 39.42 45.06 4K 28.70 45.62 56.06 56.83 54.58 49.31 28.25 33.06 32.30 25.03 27.07 16.32 70.00 90.85 42.84 44.50 4.21 99.00 97.67 60.80 48.00 48.14 6K 29.90 46.33 55.11 59.00 56.80 49.50 31.52 34.97 33.44 24.58 26.98 16.80 72.00 90.99 43.10 47.00 5.27 99.50 99.00 61.40 49.70 49.19 8K 33.57 46.77 55.48 59.16 56.87 49.37 30.88 34.54 34.17 25.12 27.01 17.13 72.50 92.28 42.81 46.50 5.83 99.50 98.67 61.48 50.90 49.55 Llama3.2-3B 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 H2O 2K 1.63 19.96 20.20 15.20 18.02 19.56 2.88 6.47 0.78 1.55 15.97 3.11 41.00 21.97 9.83 11.75 0.50 0.50 0.00 39.71 13.91 12.60 4K 2.92 31.94 33.23 33.25 24.49 28.08 7.55 10.10 5.44 6.30 22.77 4.81 53.00 38.85 20.33 15.50 1.50 7.50 6.25 51.23 22.94 20.38 6K 4.62 38.81 39.06 45.17 34.66 35.52 15.21 13.36 10.51 10.01 24.25 6.66 61.50 53.23 27.37 15.25 0.50 13.00 19.50 54.55 32.29 26.43 8K 9.65 39.66 43.20 52.60 38.09 40.41 21.46 18.55 17.80 13.28 24.67 9.12 70.00 64.30 32.19 17.00 2.00 24.50 21.50 55.00 39.09 31.15 TOVA 2K 17.14 30.14 32.44 31.64 35.96 30.05 13.08 9.62 26.15 19.70 25.04 15.47 56.50 87.81 40.48 16.75 2.50 11.50 6.50 55.51 52.36 29.35 4K 20.52 39.53 42.47 45.80 44.12 38.42 18.22 17.76 29.36 21.36 24.96 16.60 63.50 88.98 41.50 18.75 3.00 23.50 15.00 55.72 56.66 34.56 6K 20.22 39.78 45.86 52.93 49.08 41.54 20.43 24.78 30.50 22.17 25.11 16.37 66.50 89.00 42.50 21.00 4.00 46.50 20.50 55.57 57.53 37.71 8K 21.08 40.67 49.07 55.17 48.69 41.93 23.05 31.02 31.64 22.85 25.21 16.55 69.00 89.25 42.19 22.50 2.50 71.00 21.50 55.77 57.47 39.91 Sink 2K 16.85 30.69 26.58 27.32 33.26 25.27 13.82 9.38 26.74 19.15 25.15 15.88 65.00 86.17 40.79 19.50 1.50 19.50 8.50 56.65 52.73 29.54 4K 19.46 38.61 36.22 41.68 41.97 35.84 13.37 9.86 29.34 20.19 25.06 16.44 71.00 88.06 41.31 21.75 2.50 35.50 16.00 56.48 52.43 33.96 6K 19.33 40.29 37.95 49.68 46.48 40.29 15.31 11.10 30.43 21.35 25.14 16.64 71.50 88.93 42.04 23.50 3.50 47.00 19.50 56.55 54.11 36.22 8K 20.15 40.02 41.94 53.57 48.15 42.24 16.01 14.76 31.64 22.10 25.20 16.50 73.00 89.26 42.37 26.25 3.50 62.50 20.50 56.86 56.63 38.25 SnapKV 2K 17.38 31.37 31.48 29.65 37.77 30.05 11.54 9.66 27.03 19.93 24.97 15.97 59.00 88.13 40.48 16.25 3.50 32.50 9.00 56.32 55.91 30.85 4K 19.85 39.22 39.86 47.33 46.70 37.98 16.64 16.88 29.79 21.21 25.01 16.74 65.50 89.35 40.95 18.25 2.50 62.50 22.50 55.74 56.88 36.73 6K 20.83 39.65 44.48 51.84 49.30 40.18 20.28 25.32 31.27 22.73 25.09 16.81 69.00 89.95 41.47 18.75 4.00 85.00 20.50 55.69 57.82 39.52 8K 20.49 40.80 48.16 55.44 48.78 41.65 24.79 30.40 31.81 23.46 25.17 16.44 70.00 90.17 41.99 22.00 5.00 94.00 21.50 55.77 57.29 41.20 KeyL2Norm [9] 2K 7.67 30.39 31.85 30.64 29.47 25.76 7.41 14.17 15.36 12.42 24.20 7.91 48.00 50.99 23.09 17.50 2.00 7.50 5.00 48.92 26.32 22.22 4K 12.92 37.59 44.71 43.85 38.89 33.42 12.41 22.42 24.63 19.27 24.77 11.37 63.00 72.51 31.75 19.00 3.87 9.50 11.00 55.82 40.08 30.13 6K 13.02 40.55 48.17 50.87 43.18 40.17 17.10 31.29 28.99 21.47 25.08 13.30 65.00 79.61 37.16 21.50 2.50 46.50 18.50 56.49 47.05 35.60 8K 15.72 40.54 47.88 54.29 49.29 43.79 22.22 33.18 31.86 22.50 25.19 14.28 70.00 84.92 39.45 23.00 1.50 69.00 20.50 56.82 50.73 38.89 KeyDiff 2K 18.29 36.65 45.44 47.47 46.09 35.41 13.79 28.89 28.16 21.45 25.01 13.56 60.00 85.24 37.00 24.88 1.00 60.50 12.00 54.13 42.01 35.09 4K 22.34 40.60 49.15 52.56 50.14 40.30 21.65 32.46 31.38 23.44 25.06 15.28 66.50 87.92 41.41 27.50 2.50 88.50 19.50 55.55 52.24 40.28 6K 22.29 40.68 50.14 54.51 51.74 42.19 24.83 34.64 32.39 23.53 25.19 15.88 71.00 90.02 42.00 28.75 3.00 95.00 21.50 55.86 54.39 41.88 8K 22.41 40.77 50.10 55.62 49.83 43.58 28.09 34.30 32.78 23.60 25.17 15.77 72.00 90.17 42.46 30.75 3.50 96.50 21.50 55.85 55.65 42.40
 Table 11: Full Qwen-2.5-7B/3B-Instruct LongBench Results with $B=128$ (Higher is better). We highlight the best and second best methods within a given budget with bold and underline.
@@
 Llama3.1-8B 30.05 47.00 56.12 59.86 57.33 47.81 32.25 35.64 34.86 25.32 27.02 17.28 73.00 91.61 43.37 45.50 8.33 99.50 99.00 61.66 51.94 49.74 Sink 2K 21.83 34.27 29.24 32.82 38.64 29.50 12.59 16.18 28.51 20.21 26.62 15.54 65.00 89.46 42.20 22.25 2.00 25.50 32.50 64.95 59.54 33.78 4K 22.94 43.01 39.08 46.16 44.04 41.39 19.09 16.54 31.08 21.57 26.78 16.73 70.00 91.53 42.29 29.25 3.00 38.50 71.00 62.12 58.84 39.76 6K 25.41 47.40 44.13 52.78 47.39 45.73 21.90 17.55 32.53 22.19 26.87 17.05 72.00 91.25 43.41 33.75 3.08 52.50 98.00 62.22 56.24 43.49 8K 23.53 46.63 48.68 55.77 49.61 47.16 21.14 19.54 33.10 23.20 26.92 16.91 72.00 91.29 43.79 37.00 3.25 66.00 99.00 62.18 56.43 44.91 SnapKV 2K 21.81 37.22 37.19 38.29 46.10 35.42 16.53 16.37 29.83 21.05 26.77 16.16 61.00 88.84 42.56 21.75 4.03 51.50 81.17 62.37 51.45 38.45 4K 24.79 44.22 47.30 50.27 48.49 46.73 20.55 22.04 32.19 22.68 26.95 16.95 67.50 90.98 43.14 25.00 5.17 89.50 96.67 61.44 51.20 44.46 6K 24.10 45.57 50.44 55.27 53.12 48.41 24.27 27.46 33.43 23.53 27.03 16.84 71.50 92.28 43.58 27.00 5.25 98.00 99.00 61.32 52.16 46.65 8K 25.15 46.55 53.39 57.65 56.00 48.75 27.82 32.66 33.67 24.85 27.01 17.37 72.50 91.78 43.54 33.75 5.08 100.00 98.67 61.48 51.41 48.05 KeyDiff + 10% Recent Tokens 2K 27.36 40.79 50.63 50.21 49.44 44.77 25.04 30.25 29.41 23.28 26.76 15.78 64.00 85.90 43.76 44.00 4.25 98.50 95.83 62.97 48.70 45.79 4K 27.42 45.67 53.62 57.40 53.77 47.30 27.48 33.40 31.90 24.18 26.83 16.51 70.00 88.89 43.95 47.00 5.21 99.00 98.67 62.20 52.98 48.26 6K 30.58 46.24 55.58 57.82 56.44 47.74 29.03 34.33 33.15 24.67 26.91 16.98 71.50 92.22 43.96 47.00 4.23 99.50 99.00 62.56 52.98 49.16 8K 32.17 46.66 55.65 58.65 57.24 48.64 30.54 33.44 33.85 24.93 26.94 17.12 72.50 91.72 43.70 46.00 5.47 99.50 99.00 62.56 51.75 49.43 KeyDiff + 20% Recent Tokens 2K 26.73 41.88 49.79 48.46 49.68 42.51 26.90 30.90 28.83 23.34 26.76 16.11 62.50 87.39 43.99 42.50 4.58 94.00 95.58 63.33 50.01 45.51 4K 26.05 45.11 55.51 55.87 54.14 47.41 25.52 33.51 31.87 24.40 26.95 16.09 70.00 90.22 43.76 43.50 4.00 99.50 98.67 62.06 52.27 47.92 6K 28.39 46.43 55.12 58.27 57.02 48.94 28.98 34.31 33.32 24.62 26.96 17.20 71.50 91.72 44.02 46.50 5.13 99.50 99.00 62.39 52.57 49.14 8K 31.35 46.74 54.71 58.60 58.19 48.14 31.77 34.02 33.55 24.98 26.95 17.11 72.50 91.72 44.17 46.00 7.38 99.50 99.00 62.39 52.04 49.56 KeyDiff 2K 26.64 41.73 50.99 51.18 51.59 46.47 22.84 32.37 29.02 23.86 26.76 14.81 66.50 85.92 39.26 42.25 3.17 96.00 96.25 59.17 39.42 45.06 4K 28.70 45.62 56.06 56.83 54.58 49.31 28.25 33.06 32.30 25.03 27.07 16.32 70.00 90.85 42.84 44.50 4.21 99.00 97.67 60.80 48.00 48.14 6K 29.90 46.33 55.11 59.00 56.80 49.50 31.52 34.97 33.44 24.58 26.98 16.80 72.00 90.99 43.10 47.00 5.27 99.50 99.00 61.40 49.70 49.19 8K 33.57 46.77 55.48 59.16 56.87 49.37 30.88 34.54 34.17 25.12 27.01 17.13 72.50 92.28 42.81 46.50 5.83 99.50 98.67 61.48 50.90 49.55 Llama3.2-3B 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 Sink 2K 16.85 30.69 26.58 27.32 33.26 25.27 13.82 9.38 26.74 19.15 25.15 15.88 65.00 86.17 40.79 19.50 1.50 19.50 8.50 56.65 52.73 29.54 4K 19.46 38.61 36.22 41.68 41.97 35.84 13.37 9.86 29.34 20.19 25.06 16.44 71.00 88.06 41.31 21.75 2.50 35.50 16.00 56.48 52.43 33.96 6K 19.33 40.29 37.95 49.68 46.48 40.29 15.31 11.10 30.43 21.35 25.14 16.64 71.50 88.93 42.04 23.50 3.50 47.00 19.50 56.55 54.11 36.22 8K 20.15 40.02 41.94 53.57 48.15 42.24 16.01 14.76 31.64 22.10 25.20 16.50 73.00 89.26 42.37 26.25 3.50 62.50 20.50 56.86 56.63 38.25 SnapKV 2K 17.38 31.37 31.48 29.65 37.77 30.05 11.54 9.66 27.03 19.93 24.97 15.97 59.00 88.13 40.48 16.25 3.50 32.50 9.00 56.32 55.91 30.85 4K 19.85 39.22 39.86 47.33 46.70 37.98 16.64 16.88 29.79 21.21 25.01 16.74 65.50 89.35 40.95 18.25 2.50 62.50 22.50 55.74 56.88 36.73 6K 20.83 39.65 44.48 51.84 49.30 40.18 20.28 25.32 31.27 22.73 25.09 16.81 69.00 89.95 41.47 18.75 4.00 85.00 20.50 55.69 57.82 39.52 8K 20.49 40.80 48.16 55.44 48.78 41.65 24.79 30.40 31.81 23.46 25.17 16.44 70.00 90.17 41.99 22.00 5.00 94.00 21.50 55.77 57.29 41.20 KeyDiff + 10% Recent Tokens 2K 19.92 37.58 45.99 46.41 44.24 34.81 15.20 29.22 28.23 22.49 24.99 14.71 59.00 84.08 41.38 24.38 3.00 70.50 15.50 57.49 51.91 36.72 4K 21.97 40.70 49.76 52.53 47.79 41.92 20.71 32.83 31.12 23.37 25.07 15.77 67.50 87.47 41.28 25.00 2.50 91.00 20.00 57.14 56.25 40.56 6K 23.76 40.56 50.79 54.23 50.54 42.15 25.20 34.12 31.95 23.10 25.21 15.91 71.50 88.17 41.67 26.50 3.50 96.00 20.50 56.80 55.18 41.78 8K 23.65 40.58 49.96 55.71 51.52 44.10 25.95 34.41 32.80 23.77 25.25 15.75 73.50 89.67 41.83 28.75 3.00 96.00 20.50 57.00 56.12 42.37 KeyDiff + 20% Recent Tokens 2K 19.27 34.86 45.26 44.94 42.81 34.15 14.27 27.31 28.10 22.13 25.08 14.96 61.50 85.07 41.62 24.17 2.00 71.50 16.00 57.56 53.58 36.48 4K 22.56 41.28 48.37 52.00 47.50 42.64 19.64 32.28 31.17 22.83 25.08 15.68 68.50 88.17 41.38 24.25 2.00 90.50 19.50 57.04 55.05 40.35 6K 22.88 40.74 50.37 55.02 49.90 42.34 25.02 34.95 31.97 23.28 25.24 16.32 71.50 89.61 41.43 27.00 3.00 95.50 20.50 56.82 55.77 41.86 8K 23.82 40.58 50.02 55.71 51.25 43.64 24.52 34.33 33.02 23.82 25.20 16.08 73.50 89.67 41.99 29.25 3.00 97.00 20.50 57.05 56.48 42.40 KeyDiff 2K 18.29 36.65 45.44 47.47 46.09 35.41 13.79 28.89 28.16 21.45 25.01 13.56 60.00 85.24 37.00 24.88 1.00 60.50 12.00 54.13 42.01 35.09 4K 22.34 40.60 49.15 52.56 50.14 40.30 21.65 32.46 31.38 23.44 25.06 15.28 66.50 87.92 41.41 27.50 2.50 88.50 19.50 55.55 52.24 40.28 6K 22.29 40.68 50.14 54.51 51.74 42.19 24.83 34.64 32.39 23.53 25.19 15.88 71.00 90.02 42.00 28.75 3.00 95.00 21.50 55.86 54.39 41.88 8K 22.41 40.77 50.10 55.62 49.83 43.58 28.09 34.30 32.78 23.60 25.17 15.77 72.00 90.17 42.46 30.75 3.50 96.50 21.50 55.85 55.65 42.40
-Appendix C Attention Sinks and Approximate Collinearity
-(a) Keys
-(b) Keys
-(c) Queries
-(d) Queries
-Figure 11: Angular distributions of keys and queries measured with respect to to mean key and mean query
-(a) Layer 25
-(b) Layer 26
-Figure 12: Cosine Similarity of mean key and mean query for each head in several layers.
-To better understand why there exists a negative correlation between cosine similarity between keys and attention scores, we look to recent research that seeks to the importance of attention sinks in decoder-based LLMs.
-The authors in (Barbero et al., 2025) show that attention sinks emerge from training decoder-based LLMs since they can denoise the model and prevent rank collapse by limiting over mixing in attention heads.
-Moreover, attention patterns in decoder based models demonstrate that most attention logits are quite small (and almost always negative) for most keys and queries.
-This allows the attention sink to have high attention activation, preventing over mixing in addition to allowing the heads to specialize and selectively identify important tokens.
-At the same time, the results in (Godey et al., 2024) suggest that hidden states, keys and queries are all approximately collinear in the sense that $\textrm{CosSim}(x_{i},x_{j})\gg 0$.
-In geometric terms, this means that most key tokens and query tokens lie in the same direction in Euclidean space.
-Our own results, seen in Figure 11 demonstrate that this is the case for both keys and queries.
-Our experiments show that most keys and queries lie within a small angular distance from the mean key and query.
-More than this, we see that across all heads, the mean key and mean query have negative cosine similarity Figure 12.
-Moreover, as seen in Figure 13, we find that the norms of keys and queries are tightly clustered around a relatively fixed value.
-This means that variations in the norm of key and query tokens have less impact on the magnitude of attention scores than their direction.
-These three observations indicate that most keys and queries combine to create uniformly small attention logits, and that larger attention weights are constructed by projecting keys closer to the direction of the mean query.
-This appears to be the fundamental mechanism through which over mixing is prevented: if most attention activations are very small, each head can increase the activations of a small number of keys across most queries selectively projecting them to be more aligned with the distribution of queries.
-This hypothesis is further supported by the fact that sink tokens themselves often have small norm, which results in an approximate no-op in the attention head as in (Barbero et al., 2025), however in this case, the only way for a key corresponding to a sink token to have high attention scores is if it is as parallel as possible to the set of query tokens.
-(a) Keys
-(b) Keys
-(c) Queries
-(d) Queries
-Figure 13: Distribution of L2 norms for keys and queries in several layers.
-Figure 14: Average cosine similarity per head between high scoring keydiff key tokens and mean query.
-To verify the above hypothesis, we show Figure 14 that keys which have maximum angular difference from the mean key are aligned with the mean query, resulting in very large attention weights.
-This demonstrates how LLMs exploit the geometry of the hidden states and projections to limit over mixing, and selectively identify important tokens.
-To summarize, we have for all attention heads in decoder based transformer models:
-  • The majority of keys and queries are approximately collinear with their mean.
-  • Mean keys and mean queries have negative cosine similarity across all heads.
-  • Most keys and queries have a similar L2 norm.
-  • Decoder based attention heads can selectively increase attention weights for a fixed key by aligning it with the mean query.
-  • Key token importance can hence be measured by the angular distance between a key and the mean key.
-We can show mathematically with some reasonable assumptions based on the above observations that key tokens with persistently high attention scores must be geometrically aligned with queries.
-Theorem C.1.
-Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$.
-Without loss of generality suppose $||q||=1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$:
-$w=\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}.$
-Then
-$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1\leq\textrm{CosSim}({k^{*}},q)$
-Proof.
-To show this we have that
-$w$ $=\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}$
-$w\left(\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)\right)$ $=\exp({k^{*}}^{\top}q)$
-$w\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)$ $=(1-w)\exp({k^{*}}^{\top}q)$
-Note that $-M\leq k_{i}^{\top}q\leq M$ and hence
-$w\sum\limits_{i=1}^{n}\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
-$wn\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
-$\frac{wn\exp(-M)}{1-w}$ $\leq\exp({k^{*}}^{\top}q)$
-$\ \log(wn)-M-\log(1-w)$ $\leq M\textrm{CosSim}({k^{*}},q)$
-$\frac{\log(wn)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
-$\frac{\log(\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}n)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
-$\frac{\log(\frac{\exp({k^{*}}^{\top}q)}{(n+1)\exp(M)}n)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
-$\frac{-M\textrm{CosSim}(k^{*},q)+\log(\frac{n}{n+1})-2M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
-$\frac{\log(\frac{n}{n+1})-2M-\log(1-w)}{M}$ $\leq 2\textrm{CosSim}({k^{*}},q)$
-$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1$ $\leq\textrm{CosSim}({k^{*}},q)$
-The above proof demonstrates that as long as the norms of the keys are bounded, in order for an attention head to be able to freely allocate $w$ attention weight to a given key ${k^{*}}$, the cosine similarity between ${k^{*}}$ and $q$ must be high, therefore ${k^{*}}$ and $q$ must be approximately collinear.
-Using this result, we can also show that as long as the cosine similarity between ${k^{*}}$ and $q$ is high, while the cosine similarity between $\bar{k}$ and $q$ is low, $\textrm{CosSim}({k^{*}},\bar{k})$ is small.
-Note that, since empirical results demonstrate that most keys have high cosine similarity with their mean $\bar{k}$, a key with high importance, approximately collinear to $q$, will also have low cosine similarity to $\bar{k}$.
-Generally, this also suggests that key tokens with low cosine similarity to $\bar{k}$ have greater importance.
-In order to show this, we need the following auxiliary result.
-Lemma C.2.
-Suppose $\{x_{1},...,x_{n}\}$ is an orthonormal basis of $\mathbb{R}^{n}$ and $y\in\mathbb{R}^{n}$. Define $\alpha_{i}=\textrm{CosSim}(x_{i},y)$. Then $\sum\limits_{i=1}^{n}\alpha_{i}^{2}=1$.
-Proof.
-Note that $\alpha_{i}=\frac{y^{\top}x_{i}}{||y||||x_{i}||}=\frac{y^{\top}x_{i}}{||y||}$. If we expand $\frac{y}{||y||}$ in the basis $\{x_{1},...,x_{n}\}$ we see that
-$\frac{y}{||y||}$ $=\sum\limits_{i=1}^{n}\left({\frac{y}{||y||}}^{\top}x_{i}\right)x_{i}$
-$=\sum\limits_{i=1}^{n}\alpha_{i}x_{i}$
-But then, we know that since $\left|\left|\frac{y}{||y||}\right|\right|_{2}^{2}=1$, then $\left\langle\sum\limits_{i=1}^{n}\alpha_{i}x_{i},\sum\limits_{i=1}^{n}\alpha_{i}x_{i}\right\rangle=1$.
-But we have
-$\left\langle\sum\limits_{i=1}^{n}\alpha_{i}x_{i},\sum\limits_{i=1}^{n}\alpha_{i}x_{i}\right\rangle$ $=\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\langle\alpha_{i}x_{i},\alpha_{j}x_{j}\rangle$
-$=\sum\limits_{i=1}^{n}\langle\alpha_{i}x_{i},\alpha_{i}x_{i}\rangle$
-$=\sum\limits_{i=1}^{n}\alpha_{i}^{2}$
-proving the result. ∎
-Theorem C.3.
-Consider tokens ${k^{*}}$, $q$, $\bar{k}$ as above where $\bar{k}$ is the average of the keys tokens.
-Suppose $\textrm{CosSim}({k^{*}},q)=\beta_{q}>0$ and $\textrm{CosSim}(\bar{k},q)=\alpha_{q}<0$.
-Then $\textrm{CosSim}(\bar{k},{k^{*}})\leq 1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$.
-Proof.
-Consider the cosine similarity of $\bar{k}$ and ${k^{*}}$:
-$\textrm{CosSim}(\bar{k},{k^{*}})$ $=\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$
-expand $\bar{k}$ in an orthonormal basis which contains $q$, $\{q,r_{1},...,r_{n-1}\}$ such that
-$\bar{k}=||\bar{k}||\left(\alpha_{q}q+\sum\limits_{i=1}^{n-1}\alpha_{i}r_{i}\right)$
-where $\alpha_{i}=\textrm{CosSim}(\bar{k},r_{i})$.
-Additionally, define $\beta_{i}=\textrm{CosSim}({k^{*}},r_{i})$ and note that by the definition of an orthonormal basis and the cosine similarity operation, using the result from Lemma C.2 we have that $\alpha_{q}^{2}+\sum\limits_{i=1}^{n-1}\alpha_{i}^{2}=1$ and that $\beta_{q}^{2}+\sum\limits_{i=1}^{n-1}\beta_{i}^{2}=1$.
-Now we have that
-$\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$ $=\frac{{k^{*}}^{\top}\left(||\bar{k}||\alpha_{q}q+\sum\limits_{i=1}^{n-1}||\bar{k}||\alpha_{i}r_{i}\right)}{||{k^{*}}||||\bar{k}||}$
-$=\alpha_{q}\beta_{q}+\frac{1}{||{k^{*}}||}\sum\limits_{i=1}^{n-1}\alpha_{i}{k^{*}}^{\top}r_{i}$
-$=\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}\alpha_{i}\beta_{i}$
-$\leq\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}|\alpha_{i}||\beta_{i}|$
-Applying Young’s inequality we obtain
-$\leq\alpha_{q}\beta_{q}+\frac{1}{2}\sum\limits_{i=1}^{n-1}\alpha_{q}^{2}+\beta_{q}^{2}$
-$=\alpha_{q}\beta_{q}+\frac{1}{2}(1-\alpha_{q}^{2})+\frac{1}{2}(1-\beta_{q}^{2})$
-$=1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$
-Note that on the domain $\alpha_{q}\in[-1,0)$, $\beta_{q}\in(0,1]$ the function $1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$ is bounded above by $1$ and decreasing to $-\frac{1}{2}$ as $\alpha\rightarrow-1$.
-Hence, the smaller $\textrm{CosSim}(\bar{k},q)=\alpha_{q}$ is, the smaller $\textrm{CosSim}(\bar{k},{k^{*}})$ must be.
+Appendix G Ablation study
+Table 14: Anchor vector ablation study. Average of Full Longbench results for Llama 3.2-3B-Instruct. KeyDiff results match Table 10. (Higher is better)
+2K 4K 6K 8K
+KeyDiff 35.09 40.28 41.88 42.40
+$\operatorname{Pairwise}$ 35.24 40.61 41.87 42.45
+$\operatorname{Median}$ 35.43 40.67 41.89 42.26
+Selecting the Anchor Vector
+We have mainly evaluated KeyDiff using the method described in Equation 7.
+Scores to determine eviction are measured via cosine similarity with an anchor vector which can be computed in several ways.
+We run LongBench on Llama3.2-3B-Instruct with eviction policies using the following anchor choices: pairwise cosine similarity from Equation 7, denoted $\operatorname{Pairwise}$; KeyDiff, using the mean of all normalized keys as an anchor, and using the median of keys as an anchor, denoted $\operatorname{Median}$.
+Table 14 summarizes the average LongBench accuracy for the different methods to Llama 3.2-3B-Instruct.
+KeyDiff shows similar average scores to $\operatorname{Pairwise}$.
+Additionally, KeyDiff and $\operatorname{Median}$ show similar scores, demonstrating that KeyDiff is robust to the selection of the anchor design.
+Table 15: Distance metric ablation study. Average of Full Longbench results of Llama 3.2-3B-Instruct. KeyDiff results match Table 10. (Higher is better)
+2K 4K 6K 8K
+KeyDiff 35.09 40.28 41.88 42.40
+$\operatorname{DotProd}$ 30.14 38.01 41.09 42.23
+$\operatorname{Euclidean}$ 13.68 21.06 25.91 29.53
+Selecting the Similarity Metric
+We use cosine similarity as the scoring metric for eviction in KeyDiff based on our discussion in Section 2.2.
+This could be replaced with other metrics like the dot product or Euclidean distance.
+We evaluate KeyDiff variants using dot product and Euclidean distance as the similarity metric, denoted $\operatorname{DotProd}$ and $\operatorname{Euclidean}$ respectively, and report the results in Table 15.
+KeyDiff and $\operatorname{DotProd}$ show similar performance for 6K and 8K budgets.
+However, KeyDiff outperforms $\operatorname{DotProd}$ for smaller cache sizes.
+This implies that considering both the direction and the magnitude of the keys to compute similarity are important for identifying the tokens to evict.
+On the other hand, $\operatorname{Euclidean}$ shows a significant performance drop relative to KeyDiff.
+G.1 Needle in a Haystack results for Sink attention
+Figure 19: Needle in a Haystack results for Sink Attention [35]

### v4

@@
 To enforce strict memory bounds throughout the prompt prefill and token generation inference phases, we adopt a block-wise inference strategy: the input prompt is divided into smaller blocks which are processed sequentially by the model, similar to [18, 1, 37].
-After processing each block, we evict some of the cached KVs by according to an eviction policy that scores each KV, as illustrated in Figure 1.
+After processing each block, we evict some of the cached KVs by according to an eviction policy that scores each KV, as illustrated in Figure˜1.
 Unlike previous approaches that apply eviction after processing the entire prompt, this strategy satisfies memory constraints throughout the full inference process.
-However, we observe a degradation in accuracy when applying existing eviction methods in this setting (Table 1).
+However, we observe a degradation in accuracy when applying existing eviction methods in this setting (Table˜1).
 We hypothesize that the performance drop stems from a mismatch in design: existing eviction methods assume access to full-prompt attention, where key importance is computed over the entire input.
@@
 Our contributions are summarized as follows:
-  • Insight. We observe that lower pairwise cosine similarity among keys correlates with higher attention scores, suggesting its utility as a proxy for token importance. (Section 3.1)
-  • Method. We introduce KeyDiff, an eviction strategy that selects keys based on their similarity to other cached entries without relying on attention scores or future tokens. (Section 3.2)
+  • Insight. We observe that lower pairwise cosine similarity among keys correlates with higher attention scores, suggesting its utility as a proxy for token importance. (Section˜3.1)
+  • Method. We introduce KeyDiff, an eviction strategy that selects keys based on their similarity to other cached entries without relying on attention scores or future tokens. (Section˜3.2)
   • Theory.
@@
 We also show that KeyDiff solves an optimal subset selection problem that maximizes key diversity.
-(Section 3.3)
+(Section˜3.3)
   • Performance.
-KeyDiff achieves $\leq 1.5\%$ and $\leq 0.04\%$ accuracy drop on LongBench with 6K and 8K cache budgets, respectively, outperforming state-of-the-art eviction methods across Llama and Qwen models (Section 4.2), and near non-evicting baseline performance for Deepseek-R1-Distill-Llama-8B on the Math-500 reasoning benchmark.
-(Section 4.3)
-  • Efficiency. We observe up to $30\%$ end-to-end inference latency reduction using KeyDiff compared to existing KV cache eviction methods. (Section 4.5)
+KeyDiff achieves $\leq 1.5\%$ and $\leq 0.04\%$ accuracy drop on LongBench with 6K and 8K cache budgets, respectively, outperforming state-of-the-art eviction methods across Llama and Qwen models (Section˜4.2), and near non-evicting baseline performance for Deepseek-R1-Distill-Llama-8B on the Math-500 reasoning benchmark.
+(Section˜4.3)
+  • Efficiency. We observe up to $30\%$ end-to-end inference latency reduction using KeyDiff compared to existing KV cache eviction methods. (Section˜4.5)
 2 Background
@@
 $O^{\operatorname{attn}}=\operatorname{Softmax}\left(\Large\nicefrac{{QK^{\top}}}{{\sqrt{d}}}+M\right)V=AV$ (2)
-where $O^{\operatorname{attn}}\in\mathbb{R}^{T\times d}$, and the causal attention mask $M$ is an upper triangular matrix with nonzero values of $\infty$- ∞.
+where $O^{\operatorname{attn}}\in\mathbb{R}^{T\times d}$, and the causal attention mask $M$ is an upper triangular matrix with nonzero values of -$\infty$.
 2.2 KV Caching
@@
 This can be avoided by storing previously computed KVs in a KV cache $\mathcal{C}=(K,V)$ for later reuse and append the new KV corresponding $x_{T+1}$ to the cache.
-We can apply Equation 2 to an existing KV cache $\mathcal{C}$ as follows:
+We can apply Equation˜2 to an existing KV cache $\mathcal{C}$ as follows:
 $o^{\operatorname{attn}}_{T+1}=\operatorname{Softmax}\left(\Large\nicefrac{{q_{T+1}[K\|k_{T+1}]^{\top}}}{{\sqrt{d}}}+M\right)[V\|v_{T+1}],$ (3)
-where $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ are the key, query, and value states of $x_{T+1}$, and $[X\|x_{T+1}]$ represents the concatenation of $x_{T+1}$ to an existing tensor $X$ along the time dimension, $M$ is the causal attention mask accounting for both the KV cache and $x_{T+1}$.
+where $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ are the key, query, and value states of $x_{T+1}$, and $X\|x_{T+1}]$ represents the concatenation of $x_{T+1}$ to an existing tensor $X$ along the time dimension, $M$ is the causal attention mask accounting for both the KV cache and $x_{T+1}$.
 KV caching dramatically reduces the latency of $\operatorname{Attention}$ by only computing $k_{T+1}$, $q_{T+1}$, $v_{T+1}$ for each token $x_{T+1}$ and reusing the KVs in $\mathcal{C}$.
@@
 Additionally, attention-based eviction often requires explicitly materializing $A$, which can be resource intensive.
-We discuss the attention-based eviction policies further in Section 5.
+We discuss the attention-based eviction policies further in Section˜5.
 2.4 KV Caching in Resource-Constrained Environments
@@
 This can often exceed model’s allocated memory limit when deploying long context applications in resource constrained environments.
-As demonstrated in efficient LLM inference frameworks [1, 14, 18, 36], one solution is to apply $\pi_{N}$ more frequently by segmenting $X$ into non-overlapping blocks $X=[X_{0},X_{1},\ldots,X_{m-1}]$, where $X_{i}=[x_{Bi},\ldots,x_{B(i+1)-1}]$, $B$ is the block size, and $m=\lceil T/B\rceil$, and iteratively updating the cache by exploiting causality, applying Equation 4 in a block-wise fashion:
+As demonstrated in efficient LLM inference frameworks [1, 14, 18, 36], one solution is to apply $\pi_{N}$ more frequently by segmenting $X$ into non-overlapping blocks $X=[X_{0},X_{1},\ldots,X_{m-1}]$, where $X_{i}=[x_{Bi},\ldots,x_{B(i+1)-1}]$, $B$ is the block size, and $m=\lceil T/B\rceil$, and iteratively updating the cache by exploiting causality, applying Equation˜4 in a block-wise fashion:
 $\mathcal{C}_{i}$ $\leftarrow([K_{i-1}\|k_{Bi:B(i+1)-1}],[V_{i-1}\|v_{Bi:B(i+1)-1}])$ (6)
@@
 where $k_{Bi:B(i+1)-1}$ and $v_{Bi:B(i+1)-1}$ are the keys and values selected from $X_{i}$ respectively, and $\mathcal{C}_{i}=(K_{i},V_{i})$ is the KV cache after processing the first $i$ prompt blocks.
-As in Equation 4, we concatenate the $B$ new KVs to the current cache, apply $\pi_{N}$ and update the cache in Equation 6.
+As in Equation˜4, we concatenate the $B$ new KVs to the current cache, apply $\pi_{N}$ and update the cache in Equation˜6.
 We refer to this as block prompt processing.
@@
 3 Method
-We demonstrate a negative correlation between attention scores and the cosine similarity among keys (Section 3.1) and leverage this observation to develop KeyDiff (Section 3.2), followed by a theoretical justification of KeyDiff (Section 3.3) and preliminary evidence of its efficacy (Section C.2).
+We demonstrate a negative correlation between attention scores and the cosine similarity among keys (Section˜3.1) and leverage this observation to develop KeyDiff (Section˜3.2), followed by a theoretical justification of KeyDiff (Section˜3.3) and preliminary evidence of its efficacy (Section˜C.2).
 3.1 Correlation of Attention Scores and Key Dissimilarity
-To address the shortcomings of attention-based KV cache eviction in Section 2.4, we develop an alternative attention-free scoring metric that retains significant KVs across blocks while being resource efficient.
+To address the shortcomings of attention-based KV cache eviction in Section˜2.4, we develop an alternative attention-free scoring metric that retains significant KVs across blocks while being resource efficient.
 We recall the “attention sink" phenomenon: LLMs often assign high attention weight to the first few tokens, regardless of the input [35, 29]; these highly weighted tokens are called sink tokens.
@@
 We evaluate our hypothesis by inspecting the cosine similarities between keys computed inside an attention block.
-We visualize the pairwise cosine similarities between keys along with the attention weights in two particular heads and layers in Figure 2.
-We observe that keys with lower cosine similarity with other keys exhibit higher relative attention scores regardless of the choice of query, such as the 4th and 15th keys in Figure 2(a), or the 1st key in Figure 2(b).
-Pairwise cosine similarity of keys is solely a function of the keys in the cache, which are independent of input queries; the surprising aspect of Figure 2 is the negative correlation with attention weights.
+We visualize the pairwise cosine similarities between keys along with the attention weights in two particular heads and layers in Figure˜2.
+We observe that keys with lower cosine similarity with other keys exhibit higher relative attention scores regardless of the choice of query, such as the 4th and 15th keys in Figure˜2(a), or the 1st key in Figure˜2(b).
+Pairwise cosine similarity of keys is solely a function of the keys in the cache, which are independent of input queries; the surprising aspect of Figure˜2 is the negative correlation with attention weights.
 These distinctive keys essentially recover the attention sink phenomenon [35].
 3.2 KeyDiff
-Based on the observation in Section 3.1, we propose KeyDiff, which evicts tokens from the KV cache based on key similarity.
+Based on the observation in Section˜3.1, we propose KeyDiff, which evicts tokens from the KV cache based on key similarity.
 If the cache $\mathcal{C}$ has intermediate size $n$ and budget $N$ where $n>N$, $\pi^{\textsc{KeyDiff}{}}_{N}$ is defined as:
@@
 However, computing the pairwise cosine similarities runs in $\mathcal{O}(n^{2})$ time.
-Fortunately, we can compute the score of each token in Equation 7 in $\mathcal{O}(n)$ as follows:
+Fortunately, we can compute the score of each token in Equation˜7 in $\mathcal{O}(n)$ as follows:
 $S$ $=\texttt{topk}(-\operatorname{CosSim}(\mu(\hat{K}),\hat{k}_{i}),N)$ (8)
-where $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\hat{k}_{i}$ and $\hat{k}_{i}=\frac{k_{i}}{||k_{i}||}$.
-We refer to $\mu(\hat{K})$ as the anchor vector.
-We show this formulation retains the same KVs of Equation 7 under a mild condition.
-(see Section C.2).
-Our experimentation has shown that the anchor vector $\mu(\hat{K})$ can be replaced with $\mu(K)$ without losing accuracy (see Table 14).
-We evaluate the efficient KeyDiff described in Figure 3 using unnormalized keys $k$ in all subsequent sections.
-Figures 5, 8, 9 and 10 visualize the keys retained and evicted by sink attention [35], TOVA [24] and KeyDiff via PCA.
-KeyDiff retains more varied keys.
-A full complexity and FLOP analysis can be found in Appendices B and B.1
+[…]
 KeyDiff with Sliding Window
 In tasks such as reasoning and coding, where the most recent tokens are often important, we can augment KeyDiff and its efficient variant to use a percentage of the cache budget for a sliding window [6], which we call KeyDiff with sliding window.
-This extension introduces no complexity or memory overhead and we observe better results on certain tasks than vanilla KeyDiff (Tables 13 and E).
+This extension introduces no complexity or memory overhead and we observe better results on certain tasks than vanilla KeyDiff (Tables˜14 and E).
 3.3 Why KeyDiff Works: A Theoretical Perspective
 To solidify the theoretical foundation of KeyDiff and show that KeyDiff ultimately selects keys most aligned with queries, we prove the following two results.
-We first validate the relationship between cosine similarity and attention scores observed in Figure 2 by bounding the attention score of a new incoming key $k^{\star}$ in terms cosine similarity with a fixed query $q$:
+We first validate the relationship between cosine similarity and attention scores observed in Figure˜2 by bounding the attention score of a new incoming key $k^{\star}$ in terms cosine similarity with a fixed query $q$:
 Lemma 3.1.
-Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$.
-Without loss of generality suppose $||q||=1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$.
+Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $k_{i}||_{2}^{2}<M,\;\forall\;i$.
+Without loss of generality suppose q=1||q||=1 and assume $k^{*}$} is a key not in $\{k_{i}\}_{i=1}^{n}$ with ${k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$.
 Then, for $n\to\infty$,
@@
 Figure 4: PCA embedding of keys and queries from Llama 3.2 3B
-By combining Lemmas 3.1 and 3.2, we establish a relationship between the attention weight $w$ and the KeyDiff score $\mathrm{CosSim}(\bar{k},k^{*})$.
-As $\mathrm{CosSim}(\bar{k},q)$ decreases and $\mathrm{CosSim}(k^{*},q)$ increases (along with the attention weight $w$), then $\mathrm{CosSim}(\bar{k},k^{*})$ tends to $-1$: this means KeyDiff selects distinct keys most aligned with $q$.
-We visualize this in Figure 4 with a PCA embedding of keys and queries from a single head of Llama 3.2 3B, highlighting the relationship between top scoring keys via KeyDiff, the anchor vector and queries.
-Similar trends are found from the other layers and heads as shown in Figure 11.
-The proofs of Lemmas 3.1 and 3.2 are in Section C.3, along with empirical motivation for the chosen assumptions.
+By combining Lemmas˜3.1 and 3.2, we establish a relationship between the attention weight $w$ and the KeyDiff score $\mathrm{CosSim}(\bar{k},k^{*})$.
+As $\mathrm{CosSim}(\bar{k},q)$ decreases and $\mathrm{CosSim}(k^{*},q)$ increases (along with the attention weight $w$), then $\mathrm{CosSim}(\bar{k},k^{*})$ tends to 1-1: this means KeyDiff selects distinct keys most aligned with $q$.
+We visualize this in Figure˜4 with a PCA embedding of keys and queries from a single head of Llama 3.2 3B, highlighting the relationship between top scoring keys via KeyDiff, the anchor vector and queries.
+Similar trends are found from the other layers and heads as shown in Figure˜11.
+The proofs of Lemmas˜3.1 and 3.2 are in Section˜C.3, along with empirical motivation for the chosen assumptions.
 (a) Sink Attention
@@
 Retained tokens are blue, while evicted tokens are orange.
-Keys are taken from layer $5$ and head $3$ of Llama3.2-3B-Instruct, and generated using the NarrativeQA dataset.
+Keys are taken from layer 55 and head $3$ of Llama3.2-3B-Instruct, and generated using the NarrativeQA dataset.
 (d) PCA visualization of the retained keys for each KV cache eviction method.
@@
 Our findings can be summarized as follows:
-  • Needle-In-a-Haystack. KeyDiff outperforms competing eviction policies on the Needle-In-A-Haystack benchmark (Section 4.1).
+  • Needle-In-a-Haystack. KeyDiff outperforms competing eviction policies on the Needle-In-A-Haystack benchmark (Section˜4.1).
   • LongBench.
-KeyDiff outperforms competing eviction policies with block size $B=128$ on LongBench, achieving an 1.5% accuracy drop with a 6K cache budget ($\sim$33compression rate) and $\leq$ .04% with a 8k cache budget ($\sim$23compression rate) with Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Section 4.2).
+KeyDiff outperforms competing eviction policies with block size $B=128$ on LongBench, achieving an 1.5% accuracy drop with a 6K cache budget ($\sim$33% compression rate) and $\leq$ .04% with a 8k cache budget ($\sim$23% compression rate) with Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Section˜4.2).
   • Reasoning.
-KeyDiff performs competitively on the Math-500 reasoning benchmark with other eviction methods using the DeepSeek-R1-Distill-Qwen-7B and Llama-8B, and shows near eviction-free baseline performance when augmented with a sliding window (Section 4.3) for DeepSeek-R1-Distill-Llama-8B.
+KeyDiff performs competitively on the Math-500 reasoning benchmark with other eviction methods using the DeepSeek-R1-Distill-Qwen-7B and Llama-8B, and shows near eviction-free baseline performance when augmented with a sliding window (Section˜4.3) for DeepSeek-R1-Distill-Llama-8B.
   • Ablation Study.
 We perform an ablation study on the main parameters of KeyDiff and show that utilizing negative cosine similarity as the eviction criteria and the mean of cached keys as the anchor vector performs best.
-(Section 4.4).
-  • Efficiency. We compare the end-to-end inference latency of KeyDiff, [20] and [24] and observe a 30% latency improvement with KeyDiff (Section 4.5).
+(Section˜4.4).
+  • Efficiency. We compare the end-to-end inference latency of KeyDiff, [20] and [24] and observe a 30% latency improvement with KeyDiff (Section˜4.5).
 Experimental Setup
-We apply several cache eviction methods to several decoder-only transformer-based language models, including Llama $3.1$-$8$B-Instruct [10], Llama $3.2$-$3$B-Instruct [10], and Qwen $2.5$-$3$B/7B-Instruct [38].
+We apply several cache eviction methods to several decoder-only transformer-based language models, including Llama 3.13.1-88B-Instruct [10], Llama 3.23.2-33B-Instruct [10], and Qwen 2.52.5-33B/7B-Instruct [38].
 We evaluate these models using H2O [43], TOVA [24], SnapKV [20], and StreamingLLM [35], (or “sink attention") cache eviction policies, along with the eviction-free model as a baseline.
-We simulate a resource constrained environment by processing prompts and generating responses using Equation 6, with a block size of $B=128$ for prompt processing and $B=1$ for token generation using greedy decoding for all experiments.
+We simulate a resource constrained environment by processing prompts and generating responses using Equation˜6, with a block size of $B=128$ for prompt processing and $B=1$ for token generation using greedy decoding for all experiments.
 We denote the cache budgets of 2048, 4096, 6144 and 8192 as 2K, 4K, 6K and 8K, respectively.
@@
 (c) KeyDiff
-Figure 6: Accuracy across document length and needle depth for needle in a haystack test. Cache size is $6$K with $B=128$.
+Figure 6: Accuracy across document length and needle depth for needle in a haystack test. Cache size is 66K with $B=128$.
 4.1 Needle In a Haystack
@@
 This test embeds specific information (“needle") at different points within a body of unrelated text (“haystack"); finding and retaining the needle is challenging for eviction policies, which can’t know what information must be retained during block prompt processing.
-The results are shown in Figure 6 and Figure 19, where we show the recall accuracy of Llama3.23.23.23.2-$3$B-Instruct across different document lengths (x-axis) and needle depths (y-axis) with a cache size of $6$K.
+The results are shown in Figure˜6 and Figure˜19, where we show the recall accuracy of Llama3.23.2-33B-Instruct across different document lengths (x-axis) and needle depths (y-axis) with a cache size of 66K.
 KeyDiff performs similarly to TOVA, SnapKV and sink attention for shorter documents and outperforms all three methods as the document length increases.
@@
 We omit Chinese dataset results and other model results due to space limit.
-The full evaluation results are in Table 10.
+The full evaluation results are in Table˜11.
 †: A subset of samples (183/200) were evaluated due to OOM errors.
@@
 LongBench is useful for evaluating cache eviction methods in a resource constrained environments with a fixed memory budget: 51% of prompts are longer than the largest KV cache size of 8K.
-For cache budgets of 6k and 8k tokens, prompts in LongBench are compressed by 33% and 23% respectively on average, (see Section F.3 for more detail.)
-Table 1 summarizes the evaluation results of Llama 3.1-8B-Instruct and Llama 3.2-3B-Instruct on the English subset of LongBench with 2K, 4K, 6K, and 8K cache budgets using various eviction policies with block prompt processing enabled with $B=128$.
-As shown in Table 1, KeyDiff outperforms other eviction strategies across most tasks, even demonstrating better performance with smaller cache budgets.
+For cache budgets of 6k and 8k tokens, prompts in LongBench are compressed by 33% and 23% respectively on average, (see Section˜F.3 for more detail.)
+Table˜1 summarizes the evaluation results of Llama 3.1-8B-Instruct and Llama 3.2-3B-Instruct on the English subset of LongBench with 2K, 4K, 6K, and 8K cache budgets using various eviction policies with block prompt processing enabled with $B=128$.
+As shown in Table˜1, KeyDiff outperforms other eviction strategies across most tasks, even demonstrating better performance with smaller cache budgets.
 KeyDiff shows significant a improvement on the PassageRetrieval-en (PR-en) dataset, which tests whether long-term dependencies within a long prompt can be correctly recognized [4], while achieving near full-context model performance even with the smallest budget.
-Adding a sliding window to KeyDiff improves coding task performance (Table 13).
-We observed similar trends in the full LongBench task suite as shown in Table 10 and in the additional results in Appendix F.
+Adding a sliding window to KeyDiff improves coding task performance (Table˜14).
+We observed similar trends in the full LongBench task suite as shown in Table˜11 and in the additional results in Appendix˜F.
 KeyDiff exhibits similar or better performance compared to competing methods.
@@
 Additional Results
-We present the full evaluation results in Table 10 and more complete comparisons on LongBench in Appendix F, such as: standard prompt processing with a single large block (i.e.
-$B=\infty$) in Table 9; eviction method performance with Qwen 2.5-3B/7B-Instruct in Table 11; performance behavior with block sizes $B=[64,256]$ in Table 12; and performance on KeyDiff combined with a sliding window as described in Section 3.2.
-We also compare against the $L_{2}$-norm minimizing eviction method of [9] in Table 10.
+We present the full evaluation results in Table˜11 and more complete comparisons on LongBench in Appendix˜F, such as: standard prompt processing with a single large block (i.e.
+$B=\infty$) in Table˜10; eviction method performance with Qwen 2.5-3B/7B-Instruct in Table˜12; performance behavior with block sizes $B=[64,256]$ in Table˜13; and performance on KeyDiff combined with a sliding window as described in Section˜3.2.
+We also compare against the $L_{2}$-norm minimizing eviction method of [9] in Table˜11.
 4.3 Math-500 Reasoning Benchmark
@@
 Surprisingly, we found that Llama equipped with KeyDiff and a moderate KV cache budget performs comparably to, or slightly better than, the eviction-free baseline, while also outperforming SnapKV.
-We kindly refer the reader to Appendix E for additional details on the reasoning task evaluation.
+We kindly refer the reader to Appendix˜E for additional details on the reasoning task evaluation.
 4.4 Ablation Study
 We evaluate the design choices of KeyDiff, including the similarity metrics and the choice of the anchor vector, and validate the efficacy of KeyDiff.
-We provide a full description of the test setup in Appendix G and summarize the findings here:
-  • KeyDiff anchor choice does not greatly impact benchmark accuracies (See Table 14).
-  • KeyDiff using cosine similarity as the distance metric outperforms other metrics (See Table 15)
+We provide a full description of the test setup in Appendix˜G and summarize the findings here:
+  • KeyDiff anchor choice does not greatly impact benchmark accuracies (See Table˜15).
+  • KeyDiff using cosine similarity as the distance metric outperforms other metrics (See Table˜16)
 4.5 Latency and Complexity
 Additionally, in order to demonstrate that KeyDiff decreases end-to-end inference latency, we measured time to first token for the Llama 3.2 3B instruct model using different block prompt processing sizes and cache strategies.
-These results are visualized in Figures 17 and 7.
+These results are visualized in Figures˜17 and 7.
 Since KeyDiff does not require attention weight materialization, FlashAttention [8] can be used, resulting in up to 30% lower latency than TOVA and SnapKV.
-We compare the complexity of KeyDiff with competitors in Appendix B and perform a complete FLOP count in Section B.1.
+We compare the complexity of KeyDiff with competitors in Appendix˜B and perform a complete FLOP count in Section˜B.1.
 Figure 7: Time-to-first-token (TTFT) for Llama 3.2-3B using Flash Attention with different eviction strategies with block prompt processing sizes $64$, $128$, and $256$.
@@
 Unlike the sparse attention and KV cache compression methods, eviction methods evict KVs from the cache to reduce the size of the KV cache.
-As discussed in Section 2.3, the majority of the token eviction methods employ their own rules to decide the importance of the tokens by manipulating the attention score $A$.
-For example, by appropriately choosing the aggregation functions $\phi(A)$ of Equation 4, we can obtain existing attention-based eviction methods as discussed in Section A.1.
+As discussed in Section˜2.3, the majority of the token eviction methods employ their own rules to decide the importance of the tokens by manipulating the attention score $A$.
+For example, by appropriately choosing the aggregation functions $\phi(A)$ of Equation˜4, we can obtain existing attention-based eviction methods as discussed in Section˜A.1.
 Attention-based eviction may be a better choice when the entire prompt is being processed at once, as the eviction can be done by assessing the importance of all tokens simultaneously.
@@
 Supplementary Material
-[…]
 Appendix A Extended Related Work
@@
 In this section, we provide a unified framework to understand prominent attention-based eviction methods.
-As mentioned in Equation 5, we can specify attention-based eviction methods under the unified framework with proper selection of the aggregation function $\phi(A)$ as follows:
+As mentioned in Equation˜5, we can specify attention-based eviction methods under the unified framework with proper selection of the aggregation function $\phi(A)$ as follows:
   • TOVA [24]: $\phi^{\mathrm{TOVA}}(A)=A_{-1,:}$,
@@
 Appendix B Runtime and Memory Complexity
-We analyze the runtime and memory complexity for the prominent KV cache eviction algorithms in Table 2.
+We analyze the runtime and memory complexity for the prominent KV cache eviction algorithms in Table˜2.
 For a given block size $B$ and cache budget $N$, KeyDiff requires $\mathcal{O}(N+B)$ runtime and memory.
@@
 H2O accumulates attention weights over all tokens, and computes the attention over the current block, so it will require $\mathcal{O}(NB+B^{2})$ memory overhead and runtime.
-We summarize these details in Table 2
+We summarize these details in Table˜2
 Table 2: Runtime and memory complexity of token eviction methods.
@@
 We count the FLOPs required for each operation as follows:
-  • $\|k_{i}\|=\sqrt{k_{i}\cdot k_{i}}$: since $k_{i}\in\mathbb{R}^{d}$: $d$ multiplications, $d-1$ additions, one square root.
-Repeating for each $i$, this contributes $nd$ multiplications, $n(d-1)$ additions, and $n$ square roots.
+  • $\|k_{i}\|=\sqrt{k_{i}\cdot k_{i}}$: since $k_{i}\in\mathbb{R}^{d}$: $d$ multiplications, d1d-1 additions, one square root.
+Repeating for each $i$, this contributes $nd$ multiplications, n(d1)n(d-1) additions, and $n$ square roots.
   • $\frac{k_{i}}{c}$: naively, $d$ divisions, but can be rewritten as one division and $d$ multiplications. Repeating for each $i$, this contributes $nd$ multiplications and $n$ divisions.
-  • $\frac{1}{n}\sum_{i=1}^{n}c_{i}$, for $c\in\mathbb{R}^{d}$: one division, $(n-1)d$ additions.
-Combining the above, we can compute the anchor vector using $2nd$ multiplications, $2nd-n-d$ additions, $n$ square roots and $n+1$ divisions.
-To compute the cosine similarity score, we have from above that $\mu(\hat{K})\cdot k_{i}$ requires $nd$ multiplications and $n(d-1)$ additions.
-Also from above, we have that computing $\|\mu(\hat{K})\|$ requires $d$ multiplications, $d-1$ additions and one square root.
-We reuse the computation of $\|k_{i}\|$ from the previous step and compute $\|k_{i}\|\|\mu(\hat{K})\|$ in $n$ multiplications and $\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)$ with more $n$ additions (assuming boolean comparison equals addition in cost).
-We can then divide through to compute $\frac{\mu(\hat{K})\cdot k_{i}}{\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)}$ with $n$ divisions.
-Therefore, computing the cosine similarity between the anchor and each key requires $nd+d+n$ multiplications, $nd+d-1$ additions, $n+1$ divisions and one square root.
+  • $\frac{1}{n}\sum_{i=1}^{n}c_{i}$, for $c\in\mathbb{R}^{d}$: one division, (n1)d(n-1)d additions.
+Combining the above, we can compute the anchor vector using $2nd$ multiplications, 2ndnd2nd-n-d additions, $n$ square roots and $n+1$ divisions.
+[…]
 Adding everything up, KeyDiff requires:
@@
   2. 2.
-    $13nd-n-$ 1 additions,
+    3ndn13nd-n-1 additions,
   3. 3.
-    $22n+$ 2 divisions,
+    $2n+2$ divisions,
   4. 4.
@@
 can be computed in one cycle), multiplications cost three FLOPs, division is roughly 47 FLOPs, we arrive at a final FLOP count of:
-$943(3nd+d+n)+(3nd-n-1)+47*(2n+2)+(n+1)=(12d+97)n+3d+$ 94 . (A.1)
+3(3nd+d+n)+(3ndn1)+47(2n+2)+(n+1)=(12d+97)n+3d+94.3(3nd+d+n)+(3nd-n-1)+47*(2n+2)+(n+1)=(12d+97)n+3d+94. (A.1)
 This is linear in the number of keys $n$ with a small constant, relative to the quadratic complexity of the attention operator.
@@
 C.1 Additional PCA Visualizations
-In order to demonstrate the phenomena in Figure 11 persists across all heads and layers, we have included several more visualizations as seen in Figure 8, Figure 9, and Figure 10.
+In order to demonstrate the phenomena in Figure˜11 persists across all heads and layers, we have included several more visualizations as seen in Figure˜8, Figure˜9, and Figure˜10.
 (a) Sink Attention
@@
 (d) Retained keys only
-Figure 9: Keys taken from layer $20$ and head 00 of Llama3.2-3B-Instruct
+Figure 9: Keys taken from layer 2020 and head 0 of Llama3.2-3B-Instruct
 (a) Sink Attention
@@
 (d) Retained keys only
-Figure 10: Keys taken from layer $8$ and head $1$ of Llama3.2-3B-Instruct
+Figure 10: Keys taken from layer 88 and head $1$ of Llama3.2-3B-Instruct
 (a) Layer 0 Head 1
@@
 C.2 Derivation of KeyDiff from an Optimization Perspective
+To leverage the observation in Section˜3.1, we minimize the sum of pairwise cosine similarities of each key retained in the cache.
+This can be formulated as a constrained optimization problem with the keys $K\in\mathbb{R}^{n\times d}$ whose element is $k_{i}$, and budget $N$ smaller than $n$:
 […]
+[…]
+$|S|$ $=N$
 This is a combinatorial optimization problem, which is difficult to solve efficiently, particularly during inference.
-However, we can relax Section C.2 to produce a more tractable approximation to the original problem.
-First, we rewrite Section C.2 by normalizing keys $k_{i}$ such that $\hat{k}_{i}=\frac{k_{i}}{\|k_{i}\|}$ resulting in:
+However, we can relax Equation˜A.2 to produce a more tractable approximation to the original problem.
+First, we rewrite Equation˜A.2 by normalizing keys $k_{i}$ such that $\hat{k}_{i}=\frac{k_{i}}{\|k_{i}\|}$ resulting in:
 $\sum\limits_{i\in S}\sum\limits_{j\in S}\hat{k}_{i}\cdot\hat{k}_{j}$ $=\sum\limits_{i\in S}\hat{k}_{i}\cdot\left(\sum\limits_{j\in S}\hat{k}_{j}\right)=\sum\limits_{i\in S}\langle\hat{k}_{i},N\mu(\hat{K}_{S})\rangle$
+where $\mu(\hat{K}_{S})=\frac{1}{N}\sum\limits_{j\in S}\hat{k}_{j}$ is the empirical mean of normalized keys in $S$.
+This objective requires recomputing $\mu(\hat{K}_{S})$ for each candidate subset $S$.
+The sub-sampled mean tends to converge to the mean over the entire set, the problem can be further relaxed by replacing $\mu(\hat{K}_{S})$ with $\mu(\hat{K})=\frac{1}{n}\sum_{i=1}^{n}\hat{k}_{i}$.
+Dropping $N$ from the objective (since it doesn’t affect the solution) yields:
+[…]
+[…]
+$|S|$ $=N$
 […]
@@
 To empirically verify KeyDiff’s ability to retain diverse keys, we apply PCA to the keys computed in an attention block of Llama 3.2-3B-Instruct after evaluating a long context prompt.
-We visualize the distribution of retained and evicted keys by applying sink attention [35], TOVA [24], and KeyDiff as eviction policies in Figure 11.
+We visualize the distribution of retained and evicted keys by applying sink attention [35], TOVA [24], and KeyDiff as eviction policies in Figure˜11.
 Visual observation reveals the tokens retained by sink attention and TOVA tend to tightly cluster together while KeyDiff’s retained tokens are more evenly distributed.
-This observation generalizes across heads and layers, as shown in Section C.1.
-We also visualize the log determinant of the Gram matrix of the key cache $\log(\det(KK^{T}))$ generated using different eviction policies in Figure 12.
+This observation generalizes across heads and layers, as shown in Section˜C.1.
+We also visualize the log determinant of the Gram matrix of the key cache $\log(\det(KK^{T}))$ generated using different eviction policies in Figure˜12.
 This quantity corresponds to the volume of space spanned by the keys in $\mathcal{C}$.
-The distribution of volumes for KeyDiff attains higher values, indicating that the retained keys are more distinctive than TOVA and sink, corroborating the results of Figure 11.
-Details of how these plots were generated are discussed in Section F.1.
+The distribution of volumes for KeyDiff attains higher values, indicating that the retained keys are more distinctive than TOVA and sink, corroborating the results of Figure˜11.
+Details of how these plots were generated are discussed in Section˜F.1.
 Figure 12: Distribution of $\log\left(\textrm{det}\left(KK^{T}\right)\right)$ from the Qasper dataset in LongBench.
@@
 In geometric terms, this means that most key tokens and query tokens lie in the same direction in Euclidean space.
-Our own results, seen in Figure 13 demonstrate that this is the case for both keys and queries.
+Our own results, seen in Figure˜13 demonstrate that this is the case for both keys and queries.
 Our experiments show that most keys and queries lie within a small angular distance from the mean key and query.
-More than this, we see that across all heads, the mean key and mean query have negative cosine similarity Figure 14.
-Moreover, as seen in Figure 15, we find that the norms of keys and queries are tightly clustered around a relatively fixed value.
+More than this, we see that across all heads, the mean key and mean query have negative cosine similarity Figure˜14.
+Moreover, as seen in Figure˜15, we find that the norms of keys and queries are tightly clustered around a relatively fixed value.
 This means that variations in the norm of key and query tokens have less impact on the magnitude of attention scores than their direction.
@@
 Figure 16: Cosine similarity between highest KeyDiff scoring key token with mean query.
-To verify the above hypothesis, we show Figure 16 that keys which have maximum angular difference from the mean key are aligned with the mean query, resulting in very large attention weights.
+To verify the above hypothesis, we show Figure˜16 that keys which have maximum angular difference from the mean key are aligned with the mean query, resulting in very large attention weights.
 This demonstrates how LLMs exploit the geometry of the hidden states and projections to limit over mixing, and selectively identify important tokens.
@@
 Theorem C.1.
-Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$.
-Without loss of generality suppose $||q||=1$, the scaling parameter is $1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$:
+Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $k_{i}||_{2}^{2}<M,\;\forall\;i$.
+Without loss of generality suppose q=1||q||=1, the scaling parameter is $1$ and assume $k^{*}$} is a key not in $\{k_{i}\}_{i=1}^{n}$ with ${k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$:
 $w=\frac{\exp({k^{*}}^{\top}q)}{\exp({k^{*}}^{\top}q)+\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)}.$
@@
 $w\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)$ $=(1-w)\exp({k^{*}}^{\top}q)$
-Note that $-M\leq k_{i}^{\top}q\leq M$ and hence
+Note that $M\leq k_{i}^{\top}q\leq M$ and hence
 $w\sum\limits_{i=1}^{n}\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
@@
 $\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1$ $\leq\textrm{CosSim}({k^{*}},q)$
-Taking the limit as $n\to\infty$ produces Lemma 3.1
+Taking the limit as $n\to\infty$ produces Lemma˜3.1
 The above proof demonstrates that as long as the norms of the keys are bounded, in order for an attention head to be able to freely allocate $w$ attention weight to a given key $k^{*}$}, the cosine similarity between $k^{*}$} and $q$ must be high, therefore $k^{*}$} and $q$ must be approximately collinear.
@@
 Proof.
-Note that $\alpha_{i}=\frac{y^{\top}x_{i}}{||y||||x_{i}||}=\frac{y^{\top}x_{i}}{||y||}$. If we expand $\frac{y}{||y||}$ in the basis $\{x_{1},...,x_{n}\}$ we see that
-$\frac{y}{||y||}$ $=\sum\limits_{i=1}^{n}\left({\frac{y}{||y||}}^{\top}x_{i}\right)x_{i}$
+[…]
+[…]
 $=\sum\limits_{i=1}^{n}\alpha_{i}x_{i}$
-But then, we know that since $\left|\left|\frac{y}{||y||}\right|\right|_{2}^{2}=1$, then $\left\langle\sum\limits_{i=1}^{n}\alpha_{i}x_{i},\sum\limits_{i=1}^{n}\alpha_{i}x_{i}\right\rangle=1$.
-But we have
+[…]
 $\left\langle\sum\limits_{i=1}^{n}\alpha_{i}x_{i},\sum\limits_{i=1}^{n}\alpha_{i}x_{i}\right\rangle$ $=\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\langle\alpha_{i}x_{i},\alpha_{j}x_{j}\rangle$
@@
 Consider the cosine similarity of $\bar{k}$ and $k^{*}$}:
-$\textrm{CosSim}(\bar{k},{k^{*}})$ $=\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$
+[…]
 expand $\bar{k}$ in an orthonormal basis which contains $q$, $\{q,r_{1},...,r_{n-1}\}$ such that
-$\bar{k}=||\bar{k}||\left(\alpha_{q}q+\sum\limits_{i=1}^{n-1}\alpha_{i}r_{i}\right)$
+[…]
 where $\alpha_{i}=\textrm{CosSim}(\bar{k},r_{i})$.
-Additionally, define $\beta_{i}=\textrm{CosSim}({k^{*}},r_{i})$ and note that by the definition of an orthonormal basis and the cosine similarity operation, using the result from Lemma C.2 we have that $\alpha_{q}^{2}+\sum\limits_{i=1}^{n-1}\alpha_{i}^{2}=1$ and that $\beta_{q}^{2}+\sum\limits_{i=1}^{n-1}\beta_{i}^{2}=1$.
+Additionally, define $\beta_{i}=\textrm{CosSim}({k^{*}},r_{i})$ and note that by the definition of an orthonormal basis and the cosine similarity operation, using the result from Lemma˜C.2 we have that $\alpha_{q}^{2}+\sum\limits_{i=1}^{n-1}\alpha_{i}^{2}=1$ and that $\beta_{q}^{2}+\sum\limits_{i=1}^{n-1}\beta_{i}^{2}=1$.
 Now we have that
-$\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$ $=\frac{{k^{*}}^{\top}\left(||\bar{k}||\alpha_{q}q+\sum\limits_{i=1}^{n-1}||\bar{k}||\alpha_{i}r_{i}\right)}{||{k^{*}}||||\bar{k}||}$
-$=\alpha_{q}\beta_{q}+\frac{1}{||{k^{*}}||}\sum\limits_{i=1}^{n-1}\alpha_{i}{k^{*}}^{\top}r_{i}$
+[…]
+[…]
 $=\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}\alpha_{i}\beta_{i}$
@@
 $=1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$
-Note that on the domain $\alpha_{q}\in[-1,0)$, $\beta_{q}\in(0,1]$ the function $1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$ is bounded above by $1$ and decreasing to $-\frac{1}{2}$ as $\alpha\rightarrow-1$.
+Note that on the domain $\alpha_{q}\in[-1,0)$, $\beta_{q}\in(0,1]$ the function $1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$ is bounded above by $1$ and decreasing to $\frac{1}{2}$ as $\alpha\rightarrow-1$.
 Hence, the smaller $\textrm{CosSim}(\bar{k},q)=\alpha_{q}$ is, the smaller $\textrm{CosSim}(\bar{k},{k^{*}})$ must be.
+C.4 Correlation Analysis
+We report the Spearman rank correlation between key cosine similarity and attention scores for each layer of the Llama-3.2-3B-Instruct model.
+Correlations are averaged over randomly sampled LongBench-Musique prompts.
+From Table˜3, we observe a consistently high correlation ($\rho\approx 0.94$ on average) across all layers, indicating that geometrically distinctive keys (i.e., those with low pairwise cosine similarity) are strongly aligned with tokens receiving higher attention scores.
+This empirical evidence supports our theoretical claim in Section˜3.3 that key diversity serves as a reliable proxy for token importance.
+Table 3: Spearman correlation ($\rho$) between negative key cosine similarity and attention scores for each transformer layer of Llama-3.2-3B-Instruct.
+Layer 1 2 3 4 5 6 7
+$\rho$ 0.8997 0.9561 0.9332 0.9496 0.9517 0.9484 0.9587
+Layer 8 9 10 11 12 13 14
+$\rho$ 0.9578 0.9534 0.9570 0.9628 0.9554 0.9658 0.9578
+Layer 15 16 17 18 19 20 21
+$\rho$ 0.9477 0.9538 0.9280 0.9373 0.9328 0.9340 0.9311
+Layer 22 23 24 Mean $\pm$ Std
+$\rho$ 0.9140 0.9060 0.8950 $0.94\pm 0.02$
 Appendix D TTFT Analysis
-We have measured end-to-end inference latency (measured as time to first token (TTFT)) for Llama 3.2-3B via the standard Huggingface API when the model is using eager attention and FlashAttention as in Figure 17 and Figure 7.
+We have measured end-to-end inference latency (measured as time to first token (TTFT)) for Llama 3.2-3B via the standard Huggingface API when the model is using eager attention and FlashAttention as in Figure˜17 and Figure˜7.
 Tests are performed on NVIDIA A100 80GB GPUs.
@@
 The reported scores are the average of accuracies over the random samples.
-Table 3 summarizes the Math-500 evaluation results for Llama-8B.
+Table˜4 summarizes the Math-500 evaluation results for Llama-8B.
 As shown in the table, token eviction methods generally perform well even with KV cache budgets that are strictly smaller than the maximum generation length.
@@
 To further analyze the effectiveness of token eviction, we measure accuracy in cases where the context length of the baseline method (i.e., prompt length + generation length) exceeds the available KV cache budget.
-As shown in Table 4, for samples where eviction is actively triggered, KeyDiff continues to outperform the token eviction baseline (SnapKV), and often achieves accuracy close to or slightly better than the non-evicting baseline.
-We also conducted a similar evaluation on DeepSeek-R1-Distill-Qwen-7B and observed a slight performance degradation for token eviction methods compared to full KV cache baselines (See Tables 5 and 6.) However, KeyDiff still demonstrates comparable performance to SnapKV overall.
+As shown in Table˜5, for samples where eviction is actively triggered, KeyDiff continues to outperform the token eviction baseline (SnapKV), and often achieves accuracy close to or slightly better than the non-evicting baseline.
+We also conducted a similar evaluation on DeepSeek-R1-Distill-Qwen-7B and observed a slight performance degradation for token eviction methods compared to full KV cache baselines (See Tables˜6 and 7.) However, KeyDiff still demonstrates comparable performance to SnapKV overall.
 This discrepancy may stem from architectural differences that Llama uses a lower GQA [2] ratio than Qwen, which results in more information compression in the KV cache.
@@
 Table 5: Math 500 results on DeepSeek-R1-Llama-8B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
-Max Gen. Length $=$ 4K
-Num Tokens $>$ 1K (497/500 samples)
+Max Gen. Length == 4K
+Num Tokens >> 1K (497/500 samples)
 Budget Flex Exact
@@
 SnapKV 1024 0.687 0.526
-Num Tokens $>$ 2K (240/500 samples)
+Num Tokens >> 2K (240/500 samples)
 Full N/A 0.604 0.449
@@
 SnapKV 2048 0.610 0.458
-Max Gen. Length $=$ 8K
-Num Tokens $>$ 2K (353/500 samples)
+Max Gen. Length == 8K
+Num Tokens >> 2K (353/500 samples)
 Budget Flex Exact
@@
 SnapKV 2048 0.736 0.560
-Num Tokens $>$ 4K (195/500 samples)
+Num Tokens >> 4K (195/500 samples)
 Full N/A 0.650 0.455
@@
 SnapKV 4096 0.637 0.453
-Max Gen. Length $=$ 32K
-Num Tokens $>$ 2K (353/500 samples)
+Max Gen. Length == 32K
+Num Tokens >> 2K (353/500 samples)
 Budget Flex Exact
@@
 SnapKV 2048 0.736 0.560
-Num Tokens $>$ 4K (195/500 samples)
+Num Tokens >> 4K (195/500 samples)
 Full N/A 0.650 0.455
@@
 SnapKV 4096 0.637 0.453
-Num Tokens $>$ 8K (162/500 samples)
+Num Tokens >> 8K (162/500 samples)
 Full N/A 0.793 0.550
@@
 2048 0.749 0.566 2637
-Full 8K N/A 0.877 0.658 3287
+Full N/A 0.877 0.658 3287
 KeyDiff 8K 2048 0.811 0.613 3348
@@
 Table 7: Math 500 results on DeepSeek-R1-Qwen-7B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
-Max Gen. Length $=$ 4K
-Num Tokens $>$ 1K (497/500 samples)
+Max Gen. Length == 4K
+Num Tokens >> 1K (497/500 samples)
 Budget Flex Exact
@@
 SnapKV 1024 0.690 0.532
-Num Tokens $>$ 2K (328/500 samples)
+Num Tokens >> 2K (328/500 samples)
 Full N/A 0.650 0.478
@@
 SnapKV 2048 0.629 0.459
-Max Gen. Length $=$ 8K
-Num Tokens $>$ 2K (317/500 samples)
+Max Gen. Length == 8K
+Num Tokens >> 2K (317/500 samples)
 Budget Flex Exact
@@
 SnapKV 2048 0.718 0.533
-Num Tokens $>$ 4K (179/500 samples)
+Num Tokens >> 4K (179/500 samples)
 Full N/A 0.626 0.441
@@
 SnapKV 4096 0.615 0.422
-Max Gen. Length $=$ 32K
-Num Tokens $>$ 2K (360/500 samples)
+Max Gen. Length == 32K
+Num Tokens >> 2K (360/500 samples)
 Budget Flex Exact
@@
 SnapKV 2048 0.718 0.533
-Num Tokens $>$ 4K (141/500 samples)
+Num Tokens >> 4K (141/500 samples)
 Full N/A 0.787 0.524
@@
 SnapKV 4096 0.615 0.422
-Num Tokens $>$ 8K (53/500 samples)
+Num Tokens >> 8K (53/500 samples)
 Full N/A 0.550 0.275
@@
 F.1 Empirical Motivation for KeyDiff Setup
-To generate Figure 11, we used the first sample from the test split of the narrativeqa task in LongBench to prefill the KV cache of Llama3.2-3B-Instruct with a block size of $B=128$.
+To generate Figure˜11, we used the first sample from the test split of the narrativeqa task in LongBench to prefill the KV cache of Llama3.2-3B-Instruct with a block size of $B=128$.
 The KV cache had a maximum size of $4096$ while the sample was much longer, requiring KV eviction.
 We applied PCA to the key cache and repeated the process for sink attention, TOVA and KeyDiff.
-To construct Figure 12, we sample 100 prompts from the Qasper dataset in LongBench [4], compute the log determinant of $KK^{T}$ of the keys in the KV caches of each head and layer of Llama 3.2-3B-Instruct using a cache budget of $N=2048$ and a block size of $B=128$, and plot the distribution in Figure 12.
+To construct Figure˜12, we sample 100 prompts from the Qasper dataset in LongBench [4], compute the log determinant of $KK^{T}$ of the keys in the KV caches of each head and layer of Llama 3.2-3B-Instruct using a cache budget of $N=2048$ and a block size of $B=128$, and plot the distribution in Figure˜12.
 We show this key distribution for sink attention, TOVA and KeyDiff.
@@
 Prompt lengths
-We measure the number of tokens in the samples using LLama tokenizer [31]. As shown in Figure 18, LongBench exhibits variability in sample length from the datasets.
+We measure the number of tokens in the samples using LLama tokenizer [31]. As shown in Figure˜18, LongBench exhibits variability in sample length from the datasets.
 Figure 18: Histograms of sample lengths measured by number of tokens
@@
 Instead, we provide the average compression ratio, which is defined as:
-Average Compression $\text{Average Compression Ratio}=\frac{1}{I}\sum_{i=1}^{I}\frac{N}{L_{i}},$
+$\text{Average Compression Ratio}=\frac{1}{I}\sum_{i=1}^{I}\frac{N}{L_{i}},$
 where $N$ is the KV cache budget, and $L_{i}$ is the length of the $i$-th prompt (sample). We replace the summand with 1 whenever $N\geq L_{i}$, as compression doesn’t occur in that setting.
-As summarized in Table 8, 2K cache budgets have a 0.31 average compression ratio, which indicates 69% of input prompts are compressed.
+As summarized in Table˜9, 2K cache budgets have a 0.31 average compression ratio, which indicates 69% of input prompts are compressed.
 Our largest cache budget, 8K, exhibits a 0.77 average compression ratio.
@@
 Table 10: Resource unrestricted LongBench Results (Higher is better).
-All methods processes the input prompt in parallel (i.e., block size =absent=$\infty$= ∞) and make an eviction decision with all token information in the input.
+All methods processes the input prompt in parallel (i.e., block size $\infty$) and make an eviction decision with all token information in the input.
 The token eviction is made at every step of generation if the budget exceeds.
@@
 $B=64$ 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 H2O 2K 1.30 18.23 16.96 14.25 12.26 16.84 0.72 6.88 0.78 1.29 16.24 2.97 35.00 18.07 9.78 13.50 0.50 1.50 0.25 39.98 12.68 11.43 4K 2.30 31.92 31.59 32.44 25.02 22.58 4.89 9.29 5.36 5.57 23.02 4.80 50.50 33.05 18.76 13.75 1.00 2.50 3.75 50.11 21.57 18.75 6K 3.12 38.87 37.63 44.58 34.38 35.35 12.13 12.82 10.43 9.38 24.31 6.63 63.00 45.46 26.08 14.25 0.00 8.50 17.50 53.84 30.66 25.19 8K 9.11 40.09 45.04 52.58 39.24 38.25 15.88 17.82 17.97 12.96 24.70 9.26 70.50 60.57 32.71 16.00 0.50 21.50 19.50 54.72 39.26 30.39 TOVA 2K 17.24 30.03 31.04 32.07 36.58 28.97 12.17 10.50 26.35 19.78 25.07 15.20 60.50 87.45 41.07 15.50 1.00 10.50 6.00 55.30 52.36 29.27 4K 19.59 39.27 42.16 44.54 44.58 37.63 18.62 17.44 28.82 21.46 25.18 16.49 62.50 89.48 41.89 17.50 3.50 24.00 15.00 55.14 56.58 34.35 6K 21.53 40.32 46.16 52.81 49.44 40.35 18.73 25.74 30.47 22.30 25.23 16.23 66.00 90.00 42.48 21.00 3.00 47.00 18.50 55.15 58.03 37.64 8K 21.32 40.87 50.20 54.84 49.35 42.11 24.52 30.71 31.60 23.05 25.20 16.57 69.00 90.50 41.80 23.25 5.50 74.50 19.50 55.45 58.34 40.39 Sink 2K 15.68 29.91 26.61 27.42 33.16 25.43 13.36 9.37 26.70 19.25 25.01 16.04 64.50 86.33 41.04 19.00 1.50 19.00 8.50 56.48 52.91 29.39 4K 19.35 37.77 36.91 41.61 41.46 35.26 12.88 10.30 29.59 20.27 25.01 16.09 69.00 88.06 41.88 21.25 2.50 36.00 16.50 55.85 52.51 33.81 6K 19.36 40.01 38.16 49.57 46.39 39.01 14.20 10.41 30.54 21.64 25.07 16.93 71.00 88.50 42.09 23.50 3.50 47.50 19.50 56.02 53.97 36.04 8K 20.49 39.89 42.21 53.08 47.72 41.44 16.37 15.08 31.80 22.04 25.07 16.89 72.00 89.26 42.20 25.25 3.50 61.50 20.50 56.21 56.42 38.04 SnapKV 2K 18.07 31.21 30.60 31.32 37.31 30.69 11.71 9.98 26.98 19.87 25.13 16.05 61.00 87.85 40.36 16.25 3.00 32.00 8.00 56.78 55.77 30.95 4K 19.30 39.01 40.81 44.86 47.83 37.74 16.75 16.94 29.90 21.37 25.25 16.70 65.00 88.88 40.99 16.75 3.50 59.50 18.00 55.15 57.20 36.26 6K 20.85 40.32 45.58 52.80 48.03 41.63 18.47 24.98 30.68 22.32 25.12 16.71 68.50 90.00 41.49 18.25 4.50 85.00 18.50 55.46 57.30 39.36 8K 20.64 41.10 47.89 54.70 48.49 41.79 21.58 31.46 32.03 23.32 25.22 16.65 71.00 90.00 41.44 21.00 4.00 95.00 19.50 55.44 57.54 40.94 KeyDiff 2K 17.40 38.12 45.25 47.09 45.28 34.23 13.97 27.96 28.34 21.09 24.94 13.45 56.00 83.29 38.53 24.25 1.00 63.50 12.50 54.41 40.00 34.79 4K 22.38 42.00 50.84 53.16 47.34 40.56 21.43 32.82 30.96 23.32 25.08 15.35 67.50 87.09 42.53 27.50 2.00 89.00 18.50 54.39 53.09 40.33 6K 22.25 41.55 50.32 54.91 51.61 42.02 24.62 34.60 32.40 23.63 25.26 16.18 71.00 88.42 41.90 29.25 3.50 95.00 19.50 55.69 55.27 41.85 8K 21.57 41.24 50.12 55.33 49.98 43.78 27.45 34.30 32.43 23.67 25.21 15.99 71.50 90.84 42.32 30.00 3.00 96.50 19.50 55.54 56.56 42.23 $B=256$ 23.76 40.23 50.09 55.84 50.69 42.29 26.84 36.24 33.09 24.30 25.21 16.41 72.50 90.11 42.58 34.00 3.00 96.50 20.50 56.22 56.52 42.71 H2O 2K 1.87 19.19 23.35 16.06 20.95 17.91 2.33 7.31 0.83 1.73 16.29 3.28 42.50 25.82 9.83 14.50 2.75 0.00 1.00 38.59 14.57 13.36 4K 5.58 31.49 33.28 32.36 26.85 28.34 8.61 10.83 5.46 6.27 23.13 4.98 54.00 41.73 19.14 15.00 1.00 8.00 6.50 51.39 24.24 20.87 6K 7.49 37.99 41.54 44.71 39.53 36.18 15.46 13.87 10.77 10.46 24.60 6.94 61.50 58.64 27.20 15.50 1.00 15.00 19.00 54.30 32.45 27.34 8K 9.92 39.71 43.84 52.15 39.14 39.23 19.28 18.04 17.78 13.60 24.98 8.99 71.00 67.64 32.83 17.25 2.50 30.00 21.00 55.25 39.33 31.59 TOVA 2K 18.46 30.80 33.74 32.24 39.73 32.18 14.10 10.17 26.32 20.17 25.18 15.67 62.00 89.36 40.60 16.25 2.00 16.00 4.50 55.98 53.42 30.42 4K 20.36 38.18 42.53 46.18 46.83 36.60 17.81 17.09 28.99 20.54 25.23 16.34 63.50 89.13 41.55 19.12 3.50 29.50 14.50 55.55 55.91 34.71 6K 20.71 40.46 45.82 52.95 51.33 40.92 21.47 24.80 30.71 22.37 25.48 16.40 67.00 88.50 41.91 20.25 3.50 49.00 20.00 55.85 56.74 37.91 8K 20.84 40.79 48.02 54.82 50.12 40.71 25.17 30.85 31.47 22.98 25.49 16.41 71.00 89.00 41.99 22.25 2.00 76.50 21.00 55.93 57.59 40.23 Sink 2K 15.48 30.74 26.93 27.76 33.86 25.63 13.30 9.56 26.70 19.41 25.15 15.56 66.00 86.44 41.17 18.50 1.00 19.00 8.50 56.47 52.37 29.50 4K 18.91 38.38 37.28 41.76 41.81 35.17 13.07 9.96 29.23 20.61 25.01 16.38 70.50 88.06 42.07 21.75 1.50 36.00 17.00 56.05 51.98 33.93 6K 19.37 40.30 38.14 49.77 46.14 39.60 14.18 11.12 30.48 21.54 25.20 16.28 71.00 88.80 41.79 23.50 3.50 47.00 20.00 55.65 53.53 36.04 8K 20.35 40.19 42.96 53.25 47.82 41.10 17.87 14.83 31.32 22.14 25.17 16.56 72.50 89.26 42.47 26.25 5.00 62.50 20.00 56.00 56.03 38.27 SnapKV 2K 17.04 31.55 31.75 31.87 37.25 34.03 12.17 9.80 27.17 20.16 25.26 16.27 61.00 87.63 40.95 17.50 4.00 35.00 10.00 55.93 54.39 31.46 4K 19.67 39.34 40.95 45.81 44.27 38.78 16.17 16.61 29.99 21.15 25.49 16.71 66.50 89.15 40.76 19.75 3.50 66.00 19.00 55.44 56.45 36.74 6K 22.98 40.13 44.80 52.39 50.61 39.28 20.31 24.84 31.28 22.17 25.42 16.55 70.00 89.79 41.91 19.75 3.50 86.50 20.00 55.93 57.49 39.79 8K 20.23 40.75 47.40 54.77 49.74 41.47 24.74 31.19 31.95 22.96 25.45 16.27 71.50 90.17 42.12 23.25 4.00 93.50 21.00 55.93 57.36 41.23 KeyDiff 2K 18.99 37.20 46.57 46.61 45.14 33.12 15.15 29.54 28.30 21.92 25.41 14.44 58.50 86.30 37.77 23.75 1.00 67.00 13.50 54.07 41.59 35.52 4K 21.00 41.48 48.69 52.94 47.07 40.26 22.69 33.07 31.11 23.35 25.22 15.57 65.50 88.72 41.79 26.50 2.00 91.00 19.00 55.79 51.50 40.20 6K 21.61 40.78 49.68 55.08 50.64 41.38 24.57 34.62 32.02 23.36 25.56 15.86 71.50 89.42 42.37 28.75 3.00 94.00 21.00 55.96 55.96 41.77 8K 22.24 40.83 50.23 55.45 50.50 43.28 28.37 33.88 32.67 23.85 25.48 15.83 72.50 90.34 42.31 30.50 3.00 95.50 21.00 56.01 55.58 42.35
-[…]
+Table 14: KeyDiff + Recent tokens Llama-3.1-8B/3.2-3B-Instruct LongBench Results with $B=128$ (Higher is better).
+We highlight the methods showing the best performance within a given budget with boldface.
+$\text{X}\%$ indicates $\text{X}\%$ of KV cache budget is reserved to keep the recent tokens, while the remaining cache budget is managed by KeyDiff algorithm.
+†: A subset of samples were evaluated due to OOM errors (183/200 samples are evaluated).
 Single Doc.
@@
 Appendix G Ablation study
-Table 14: Anchor vector ablation study. Average of Full Longbench results for Llama 3.2-3B-Instruct. KeyDiff results match Table 10. (Higher is better)
+Table 15: Anchor vector ablation study. Average of Full Longbench results for Llama 3.2-3B-Instruct. KeyDiff results match Table˜11. (Higher is better)
 2K 4K 6K 8K
@@
 Selecting the Anchor Vector
-We have mainly evaluated KeyDiff using the method described in Equation 7.
+We have mainly evaluated KeyDiff using the method described in Equation˜7.
 Scores to determine eviction are measured via cosine similarity with an anchor vector which can be computed in several ways.
-We run LongBench on Llama3.2-3B-Instruct with eviction policies using the following anchor choices: pairwise cosine similarity from Equation 7, denoted $\operatorname{Pairwise}$; KeyDiff, using the mean of all normalized keys as an anchor, and using the median of keys as an anchor, denoted $\operatorname{Median}$.
-Table 14 summarizes the average LongBench accuracy for the different methods to Llama 3.2-3B-Instruct.
+We run LongBench on Llama3.2-3B-Instruct with eviction policies using the following anchor choices: pairwise cosine similarity from Equation˜7, denoted $\operatorname{Pairwise}$; KeyDiff, using the mean of all normalized keys as an anchor, and using the median of keys as an anchor, denoted $\operatorname{Median}$.
+Table˜15 summarizes the average LongBench accuracy for the different methods to Llama 3.2-3B-Instruct.
 KeyDiff shows similar average scores to $\operatorname{Pairwise}$.
 Additionally, KeyDiff and $\operatorname{Median}$ show similar scores, demonstrating that KeyDiff is robust to the selection of the anchor design.
-Table 15: Distance metric ablation study. Average of Full Longbench results of Llama 3.2-3B-Instruct. KeyDiff results match Table 10. (Higher is better)
+Table 16: Distance metric ablation study. Average of Full Longbench results of Llama 3.2-3B-Instruct. KeyDiff results match Table˜11. (Higher is better)
 2K 4K 6K 8K
@@
 Selecting the Similarity Metric
-We use cosine similarity as the scoring metric for eviction in KeyDiff based on our discussion in Section 2.2.
+We use cosine similarity as the scoring metric for eviction in KeyDiff based on our discussion in Section˜2.2.
 This could be replaced with other metrics like the dot product or Euclidean distance.
-We evaluate KeyDiff variants using dot product and Euclidean distance as the similarity metric, denoted $\operatorname{DotProd}$ and $\operatorname{Euclidean}$ respectively, and report the results in Table 15.
+We evaluate KeyDiff variants using dot product and Euclidean distance as the similarity metric, denoted $\operatorname{DotProd}$ and $\operatorname{Euclidean}$ respectively, and report the results in Table˜16.
 KeyDiff and $\operatorname{DotProd}$ show similar performance for 6K and 8K budgets.
@@
 Figure 19: Needle in a Haystack results for Sink Attention [35]
+Appendix H Additional Experiments and Analyses
+To further substantiate the empirical findings presented in Sections˜3 and 4, we provide an extended set of experiments and complementary analyses.
+These additional studies examine the correlation between key geometry and attention, evaluate retrieval-critical and reasoning scenarios, and measure the efficiency of KeyDiff under constrained hardware settings.
+Together, they offer a broader view of the method’s robustness, efficiency, and general applicability across diverse inference conditions.
+H.1 Retrieval-Critical Evaluation: Phonebook Lookup
+We evaluate KeyDiff on a retrieval-critical setting using the Phonebook Lookup task, where the model retrieves a phone number corresponding to a queried name from a long list of entries.
+Accuracy is averaged across five random phonebooks of varying lengths.
+As shown in Table˜17, KeyDiff maintains high retrieval accuracy for shorter contexts and degrades more gracefully than attention-based baselines as context length increases.
+Table 17: Accuracy (%) on the Phonebook Lookup task using Llama-3.2-3B-Instruct with a 6k cache budget.
+Method 100 478 856 1233 1611 1989 2367 2744 3122 3500
+Dense 1.0 1.0 1.0 0.8 1.0 0.8 0.6 0.6 0.8 0.6
+KeyDiff 1.0 1.0 1.0 0.6 0.4 0.2 0.0 0.0 0.2 0.2
+TOVA 1.0 1.0 1.0 0.4 0.2 0.0 0.0 0.0 0.0 0.0
+H.2 RULER Benchmark Validation
+To assess the generalizability of KeyDiff across diverse architectures, we reference the results reported by the community-maintained KVPress RULER benchmark.
+KeyDiff consistently achieves competitive or superior leaderboard scores compared to methods such as TOVA, SnapKV, QFilter, and Knorm on both Llama-3.2-3B and Qwen-3-8B backbones, demonstrating its robustness and transferability.
+H.3 Needle-in-a-Haystack Recall Saturation
+We further analyze the recall behavior of KeyDiff under varying context lengths in the Needle-in-a-Haystack benchmark.
+Table 18 reports the recall difference between KeyDiff and the full-cache baseline as a function of depth and context length.
+As shown in Table˜18, KeyDiff achieves near-parity recall with the full-cache baseline up to 30k tokens, confirming its stability under long-context compression.
+Table 18: Recall difference (KeyDiff-Full) on NIAH benchmark for Llama-3.2-3B.
+Depth 1000 4222 7444 10667 13889 17111 20333 23556 26778 30000
+0.0 0.02 -0.06 -0.06 0.20 0.16 -0.16 0.18 0.20 0.06 0.04
+11.0 0.10 -0.04 -0.04 0.30 0.28 -0.18 0.26 0.28 0.10 -0.08
+22.0 0.04 -0.12 0.00 0.32 0.30 -0.18 0.16 0.28 0.14 -0.04
+33.0 0.02 0.06 -0.20 0.34 0.08 -0.16 0.26 0.26 0.20 -0.16
+44.0 -0.06 -0.08 -0.16 0.30 0.26 -0.12 0.24 0.22 -0.10 0.04
+56.0 -0.06 0.00 -0.24 0.26 0.26 -0.20 0.18 0.14 0.04 -0.02
+67.0 0.02 0.02 -0.14 0.36 0.28 -0.26 0.24 0.16 0.04 -0.08
+78.0 -0.10 0.08 -0.26 0.26 0.30 -0.26 0.20 0.30 0.02 -0.04
+89.0 0.02 0.08 -0.06 0.30 0.26 -0.32 0.30 0.22 0.06 -0.12
+100.0 0.00 -0.06 -0.26 0.00 -0.06 -0.02 0.00 -0.04 0.14 -0.04
+H.4 On-Device Latency Evaluation
+We further evaluate the runtime characteristics of KeyDiff on a mobile-class device by measuring the latency required to compute key-eviction scores on an Android platform.
+All methods are tested on a recent Samsung smartphone under identical FP16 inference precision, and the results are normalized by the latency of KeyDiff with a cache budget of 512 KVs.
+As shown in Table˜19, KeyDiff performs on par with competing methods for small cache sizes and achieves substantially lower scoring latency as the cache size increases, demonstrating both scalability and minimal overhead on edge hardware.
+Minor runtime fluctuations can be attributed to kernel-level optimizations and proprietary hardware characteristics.
+Table 19: Relative latency of key scoring on an Android device (normalized by KeyDiff at 512 KVs). Lower is better.
+Method 512 1024 2048 4096 8192
+SnapKV 1.92 2.46 4.50 8.94 11.42
+TOVA 1.02 0.94 1.25 1.95 3.06
+KeyDiff 1.00 1.03 1.01 1.07 1.37
+H2O 1.10 0.99 1.52 2.32 4.23
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+    • Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+    • Answer: [Yes]
+    • Justification: The claims made in the abstract are corroborated by both our theoretical justification Section˜3.3, experiments Section˜4 and auxiliary results and discussion.
+  2. 2.
+    Limitations
+    • Question: Does the paper discuss the limitations of the work performed by the authors?
+    • Answer: [Yes]
+    • Justification: In Section˜6 we discuss the fact that KeyDiff is only tested on decoder based models featuring GQA and would like to extend it in the future to models with different designs such as MLA.
+  3. 3.
+    Theory Assumptions and Proofs
+    • Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+    • Answer: [Yes]
+    • Justification: Proofs of all results are discussed briefly in their statement in Section˜3.3 and rigorously proved in Section˜C.3.
+  4. 4.
+    Experimental Result Reproducibility
+    • Question: If the contribution is a dataset or model, what steps did you take to make your results reproducible or verifiable?
+Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), release of a model checkpoint, or other means that are appropriate to your research.
+    • Answer: [N/A]
+    • Justification: The paper releases neither a dataset or a model, though the cache management algorithm can be easily implemented using the description from the paper.
+  5. 5.
+    Open access to data and code
+    • Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+    • Answer: [No]
+    • Justification: Code is not provided as it is proprietary. However the algorithm may be readily implemented and results replicated using the description in the paper.
+  6. 6.
+    Experimental Setting/Details
+    • Question: If you ran experiments, did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)?
+The full details can be provided with the code, but the important details should be in the main paper, and information about how hyperparameters were selected should appear either in the paper or supplementary materials.
+    • Answer: [Yes]
+    • Justification: All hyperparameters selected for the benchmarks run to evaluate KeyDiff are specified in Sections˜4, F.2 and E.
+Hyperparameters to recreate figures are given in figure descriptions.
+  7. 7.
+    Experiment Statistical Significance
+    • Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+    • Answer: [Yes]
+    • Justification: The only experiment using stochasticity Figure˜7 and Figure˜17 reports error bars and statistical significance.
+The Math500 results as well discuss the use of sampling and stochasticity in generating responses.
+Aside from this, the remaining results are deterministic.
+  8. 8.
+    Experiments Compute Resources
+    • Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+    • Answer: [Yes]
+    • Justification: In all relevant cases, particularly experiments on TTFT Figure˜17 and Figure˜7, we mention the GPUs and compute resources used.
+  9. 9.
+    Code Of Ethics
+    • Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+    • Answer: [Yes]
+    • Justification: The research proposes a new cache management algorithm based on the geometry of the attention mechanism in decoder-based LLMs.
+It does not involve human subjects or obviously ethically sensitive applications, and it conforms to the NeurIPS Code of Ethics.
+  10. 10.
+    Broader Impacts
+    • Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+    • Answer: [N/A]
+    • Justification: This research is foundational in nature and is not tied to particular applications or deployments. As such we do not see any direct path to negative applications of our work.
+  11. 11.
+    Safeguards
+    • Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+    • Answer: [N/A]
+    • Justification: The paper proposes a new cache eviction algorithm.
+Neither the models used, data or cache eviction algorithm appear to pose a high risk for abuse necessitating release safeguards beyond standard practices.
+  12. 12.
+    Licenses for existing assets
+    • Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+    • Answer: [Yes]
+    • Justification: Creators for datasets/benchmarks LongBench and Math500 are properly credited with citations.
+Open source models from the Llama and Qwen model classes are used to test KeyDiff are properly cited as are their deepseek distill r1 variants.
+  13. 13.
+    Assets
+    • Question: If you are releasing new assets, did you document them and provide these details alongside the assets?
+Researchers should communicate the details of the dataset or the model as part of their submissions via structured templates.
+This includes details about training, license, limitations, etc.
+    • Answer: [N/A]
+    • Justification: The research does not release new assets.
+  14. 14.
+    Crowdsourcing and Research with Human Subjects
+    • Question: If you used crowdsourcing or conducted research with human subjects, did you include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+Including this information in the supplemental material is fine, but if the main contribution of your paper involves human subjects, then we strongly encourage you to include as much detail as possible in the main paper.
+According to the NeurIPS Code of Ethics, you must pay workers involved in data collection, curation, or other labor at least the minimum wage in your country.
+    • Answer: [N/A]
+    • Justification: The research does not involve crowdsourcing experiments or research with human subjects.
+  15. 15.
+    IRB Approvals
+    • Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+    • Answer: [N/A]
+    • Justification: The research does not involve human subjects, therefore IRB approval is not applicable.
+  16. 16.
+    Declaration of LLM usage
+    • Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+    • Answer: [N/A]
+    • Justification: The research does not include any input from an LLM. In particular the core methodology and research is original to the authors.